### v1

# Differential Privacy of Cross-Attention with Provable Guarantee

Differential Privacy of Cross-Attention with Provable Guarantee
Jiuxiang Gu jigu@adobe.com. Adobe Research.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song zsong@adobe.com. Adobe Research.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Cross-attention has become a fundamental module nowadays in many important artificial intelligence applications, e.g., retrieval-augmented generation (RAG), system prompt, guided stable diffusion, and many so on. Ensuring cross-attention privacy is crucial and urgently needed because its key and value matrices may contain sensitive information about companies and their users, many of which profit solely from their system prompts or RAG data. In this work, we design a novel differential privacy (DP) data structure to address the privacy security of cross-attention with a theoretical guarantee. In detail, let $n$ be the input token length of system prompt/RAG data, $d$ be the feature dimension, $0<\alpha\leq 1$ be the relative error parameter, $R$ be the maximum value of the query and key matrices, $R_{w}$ be the maximum value of the value matrix, and $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods. Then, our data structure requires $\widetilde{O}(ndr^{2})$ memory consumption with $\widetilde{O}(nr^{2})$ initialization time complexity and $\widetilde{O}(\alpha^{-1}r^{2})$ query time complexity for a single token query. In addition, our data structure can guarantee that the user query is $(\epsilon,\delta)$-DP with $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}R^{2s}R_{w}r^{2})$ additive error and $n^{-1}(\alpha+\epsilon_{s})$ relative error between our output and the true answer. Furthermore, our result is robust to adaptive queries in which users can intentionally attack the cross-attention system. To our knowledge, this is the first work to provide DP for cross-attention. We believe it can inspire more privacy algorithm design in large generative models (LGMs).
1 Introduction
The development of Artificial Intelligence (AI) has four stages: (1) prediction AI, e.g., ResNet [51] in image classification; (2) generation AI, e.g., ChatGPT [1] in language generation; (3) autonomous agent AI, Voyager [90] autonomously plays Minecraft game [28]; (4) Artificial Generalization Intelligence (AGI). Humans have made rapid progress in generative AI, and we are excitingly heading to the third stage, the era of AI agent [64]. One prevalent application of AI agents is customized large generative models (LGMs) agents [69], e.g., AgentGPT [31], SuperAGI [34], MetaGPT [50, 47], GPT Researcher [33] and many so on. In particular, recently, Apple Inc. introduced Apple Intelligence [5], signaling the integration of LGMs into physical devices. This innovation allows devices to use personal information for real-life assistance, such as entering passport numbers when booking flights or informing users of their latest meetings. With increased AI capabilities, privacy concerns become significant, as the more personal information devices handle, the greater the potential privacy risks.
One fundamental technique used in LGMs is cross-attention [86], which is an essential module in retrieval-augmented generation (RAG) [62], system prompt, guided stable diffusion, and many so on. In RAG, to be more professional, the LGMs can answer user input queries by using a domain-specific database under cross-attention, which may contain specific privacy data and knowledge so that the LGMs gain additional power. For system prompts, based on cross-attention, some customized long prompts, e.g., user information or concrete rules, are concatenated before user input to follow human instructions better, which are commonly used in ChatGPT [32], Claude3 [4] and other commercial LGMs.
Consequently, protecting the privacy of domain-specific data in RAG or system prompts is crucial as they contain sensitive information about users and companies. These data and prompts are the core assets of many start-ups. However, these data and prompts can be easily recovered [55], jailbroken [52], and released [54] by user adversarial attack [95], e.g., there are 1700 tokens in ChatGPT system prompts [72]. These findings highlight the critical importance of robust privacy protections in LGMs, making privacy not just essential but an urgent issue that demands immediate attention.
To fundamentally preserve cross-attention privacy, we can borrow the powerful tools from differential privacy (DP) [17], which provides measurable privacy and combines with statistical machine learning seamlessly [73]. Thus, in this work, we would like to ask and answer the following question,
How can we use differential privacy to protect the security of cross-attention in LGMs?
Our work demonstrates that the cross-attention computation is equivalent to computing the weighted distance problem.
Definition 1.1 (Cross-attention).
Let $n$ and $m$ be the token length of the system prompt/RAG data and input query, respectively. Let $d$ be the feature dimension. Let $\phi:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be any function that measures vector similarity. Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is defined as
$\mathrm{Attn}(Q,K,V):=D^{-1}AV,$
where $A\in\mathbb{R}^{m\times n}$ and $D\in\mathbb{R}^{m\times m}$ is a diagonal matrix, i.e.,
$A_{i,j}:=\phi(Q_{i},K_{j}),\quad\text{and}\quad D:=\mathrm{diag}(A{\bf 1}_{n}).$
Our main theorem, presented below, provides a robust solution of cross-attention, ensuring privacy and accuracy guarantees.
Theorem 1.2 (Main result; Informal version of Theorem 5.1).
Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma 6.7). Let $l=\widetilde{O}(r)$. Let $\phi$ be the softmax distance function (Definition 4.1). Then, our Alg. 4 requires $O(lndr)$ memory to make cross-attention DP with $O(lnr)$ initialization time and $\widetilde{O}(\alpha^{-1}lr)$ query time, such that with probability $1-p_{f}$, the output is $(\epsilon,\delta)$-DP and robust to adaptive query with relative error $n^{-1}(\alpha+\epsilon_{s})$ and additive error $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}lR^{2s}R_{w}r)$.
Our main technique in Theorem 1.2 makes cross-attention differentially private by using the polynomial kernel approximation method and transforming it into a weighted distance problem. Then, the above problem can be solved by summing over weighted distances (depending on the value embedding) between the query embedding and the key embedding. We build a data structure for weighted softmax queries in Section 4.2, and we extend this data structure to handle adaptive queries using the $\epsilon_{0}$-net/metric entropy argument in Section 4.3.
Our contributions can be summarized as follows:
  • We demonstrate that cross-attention computations are equivalent to the weighted distance problem (Section 3.4).
  • We design a novel algorithm (Alg. 3) that can privately answer weighted softmax queries with high probability and a concrete accuracy bound.
  • Our algorithm (Alg. 4) can handle multiple queries and is robust against adaptive query attacks (Theorem 4.4), meaning that potential attackers cannot intentionally extract information of system prompts/RAG data.
To our knowledge, this is the first work to utilize DP to protect prompts in LGMs with theoretically provable guarantees. While some papers have explored protecting user/system prompts with DP [24, 66], they are primarily empirical and lack theoretical guarantees. Additionally, many others are working on protecting private datasets by applying DP to the fine-tuning stage of LGMs training [7, 77, 59, 96, 63, 81], which diverges from our work. The strength of DP lies in its strong, unambiguous, and concrete definition of privacy, enabling algorithm designs with provable privacy and accuracy analysis. Therefore, we believe that the theoretical aspects of DP applications in LGMs remain a highly impactful direction, and we aim to pave the way for further exploration in this area.
Roadmap.
In Section 2, we provide a review of related literature. In Section 3, we present the preliminary of differential privacy (DP) and cross-attention. In Section 4, we outline the main results of our algorithms. In Section 5, we present the main result of our cross-attention theorem (Theorem 5.1). In Section 6, we offer an overview of our techniques. In Section 7, we discuss DP-related topics and potential extensions. In Section 8, we conclude our paper.
2 Related Work
2.1 Differential Privacy Guarantee Analysis
Ever since [17] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [20, 61, 98, 73, 94]. DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [22, 3, 57, 49, 35, 8, 10, 23, 14, 46, 67, 68, 53, 58, 26, 27, 56, 15, 11, 16, 25, 42, 37, 60, 36, 21, 84, 19, 91, 83, 44]. Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [18]. For example, truncated Laplace mechanism [30] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions.
2.2 Differential Privacy in Data Structure and Attention
Differential privacy (DP) is a flourishing and powerful technique that has enormous applications in the topic of private machine learning. In the era of Large Generative Models (LGMs), there are three primary approaches to ensuring privacy: (1) during the pre-training stage: to protect training data [2, 73], (2) during the adaptation stage: to protect target data [7, 77, 59, 96, 63, 81], (3) during the inference stage: to protect user/system prompts [24] and RAG data [62].
To protect training data, DP-SGD [2] uses DP optimizer to ensure data privacy, severing as the traditional baseline method. Recently, numerous works have aimed to improve this method by integrating DP in both the pre-training and fine-tuning stages of LGMs [96, 63, 29, 7, 81, 65, 77, 99, 59]. To protect user/system prompts, [24] provides a survey on both DP and non-DP methods. In the use of LGMs, prompting methods almost become a standard way for inference [80]. Given the billions of prompt interactions daily, ensuring privacy is essential [66].
2.3 Cross-Attention in System Prompt, RAG, Stable Diffusion and More
Cross-attention [86], first introduced in language translation, is a widely used technique in many advanced AI systems. For example, Stable Diffusion [76] and SORA [70] employ cross-attention as a core module for a text-to-image conditional generation. This technique is also utilized by other multimodal models [40], including Imagen [79] and Diffusion Transformer [74].
In the realm of text-to-image editing, [48] analyzes and controls the cross-attention module to enable editing without requiring additional training. Furthermore, [97] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas.
In addition, Retrieval Augmented Generation (RAG) [62, 9, 45], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module.
Cross-attention also has other applications. [71] demonstrates that the prompt-tuning [39] task can be formulated as cross-attention, while [12] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation. Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [89], such as spatial reasoning [87], mathematical reasoning [38], in-context learning ability [82], compositional ability [92], few-shot adaptation ability [78, 93], and so on.
3 Preliminary
In this section, we give the preliminary of differential privacy (DP) and cross-attention. In Section 3.1, we describe the notations we use. In Section 3.2, we give several definitions related to DP. In Section 3.3, we give the DP facts we use in the paper. In Section 3.4, we show that the cross-attention module is equivalent to the weighted distance problem.
3.1 Notations
We use $\Pr[]$ to denote the probability. We use $\operatorname*{{\mathbb{E}}}[]$ to denote the expectation. We use $\operatorname*{\mathrm{Var}}[]$ to denote the variance. For two vectors $x\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$, we use $\langle x,y\rangle$ to denote the inner product between $x,y$, i.e., $\langle x,y\rangle=\sum_{i=1}^{d}x_{i}y_{i}$. We use $X\subset\mathbb{R}^{d}$ and $|X|=n$ to mean the same thing as $X\in\mathbb{R}^{n\times d}$. Also, we denote $x_{i}^{\top}$ as the $i$-th row of $X$. We use $x_{i,j}$ to denote the $j$-th coordinate of $x_{i}\in\mathbb{R}^{n}$. We use ${\bf 1}_{n}$ to denote a length-$n$ vector where all the entries are ones. We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e., $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$.
3.2 Differential Privacy Definitions
In this section, we give several definitions related to differential privacy (DP). We refer the reader to [18] for more background and details on DP.
Definition 3.1 (Neighboring dataset).
We define the two neighboring datasets as $X,X^{\prime}\in\mathbb{R}^{n}$ such that $\|X-X^{\prime}\|_{1}\leq 1$, i.e., they differ on a single data point.
Definition 3.2 (Sensitivity).
The sensitivity of a function $f:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is defined by:
$\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n},\|X-X^{\prime}\|_{1}=1}\|f(X)-f(X^{\prime})\|_{1}.$
Definition 3.3 ($(\epsilon,\delta)$-DP).
For $\epsilon>0,\delta\geq 0$, a randomized algorithm $\mathcal{A}$ is $(\epsilon,\delta)$-DP, if for all $\mathcal{S}\subseteq\mathrm{Range}({\mathcal{A}})$ and for all $X,X^{\prime}$ such that $\|X-X^{\prime}\|_{1}\leq 1$:
$\Pr[{\mathcal{A}}(X)\in\mathcal{S}]\leq\exp(\epsilon)\Pr[{\mathcal{A}}(X^{\prime})\in{\mathcal{S}}]+\delta.$
When $\delta=0$, the algorithm is said to have pure differential privacy.
For the Laplace mechanism, we have the following definitions.
Definition 3.4 (Laplace distribution).
We use $\mathrm{Lap}(b)$ to denote the pdf:
$p(z)=\frac{1}{2b}\exp(-\frac{|z|}{b}).$
Fact 3.5.
For $z\sim\mathrm{Lap}(b)$, $\operatorname*{{\mathbb{E}}}[z]=0$, and $\operatorname*{\mathrm{Var}}[z]=2b^{2}$. Furthermore, if $b=\Delta/\epsilon$, we have $\operatorname*{\mathrm{Var}}[z]=2\Delta^{2}/\epsilon^{2}$.
Definition 3.6 (Truncated Laplace distribution).
We use $\mathrm{TLap}(\Delta,\epsilon,\delta)$ to denote the Truncated Laplace distribution with pdf proportional to $\exp(-\epsilon|z|/\Delta)$ on the region $[-B,B]$, where $B=\frac{\Delta}{\epsilon}\cdot\log(1+\frac{\exp(\epsilon)-1}{2\delta})$.
Fact 3.7 (Theorem 3 in [30]).
Let $z$ denote a $\mathrm{TLap}(\Delta,\epsilon,\delta)$ random variable. Then we have $\operatorname*{{\mathbb{E}}}[z]=0$, and
$\operatorname*{\mathrm{Var}}[z]=\frac{2\Delta^{2}}{\epsilon^{2}}(1-\delta\cdot\frac{\log^{2}(1+\frac{e^{\epsilon}-1}{2\delta})+2\log(1+\frac{e^{\epsilon}-1}{2\delta})}{e^{\epsilon}-1}).$
Furthermore, if $\delta=0$, we have $\operatorname*{\mathrm{Var}}[z]=2\Delta^{2}/\epsilon^{2}$, meaning truncated Laplacian mechanism will be reduced to the standard Laplacian mechanism.
Lemma 3.8 (Laplace mechanism, [18, 30], see Lemma 2.2 in [3]).
Given a numeric function $f$ that takes a dataset $X$ as the input, and has sensitivity $\Delta$, the mechanism that outputs $f(X)+z$ where $z\sim\mathrm{Lap}(\Delta/\epsilon)$ is $(\epsilon,0)$-DP. In addition, if $\epsilon,\delta\in(0,0.5)$, $f(X)+z$, where $z\sim\mathrm{TLap}(\Delta,\epsilon,\delta)$ is $(\epsilon,\delta)$-DP. Moreover, the truncated Laplace mechanism is always accuracy up to error $B$.
3.3 DP Facts
In this section, we present several facts about differential privacy (DP). We first state the post-processing property, which means, in an algorithm, if one step is DP, all the following steps are DP.
Fact 3.9 (Post-processing, see Fact 2.1 in [35]).
Let ${\mathcal{A}}_{1}$ be an $(\epsilon,\delta)$-DP algorithm and ${\mathcal{A}}_{2}$ be a (randomized) post-processing algorithm. Then the algorithm ${\mathcal{A}}(X)={\mathcal{A}}_{2}({\mathcal{A}}_{1}(X))$ is still an $(\epsilon,\delta)$-DP algorithm.
If we have many DP algorithms, we need a composition rule. The most straightforward composition is the basic/sequential composition rule.
Fact 3.10 (Basic composition, see Fact 2.3 in [35]).
Let ${\mathcal{A}}_{1}$ be an $(\epsilon_{1},\delta_{1})$-DP algorithm and ${\mathcal{A}}_{2}$ be an $(\epsilon_{2},\delta_{2})$-DP algorithm. Then ${\mathcal{A}}(X)=({\mathcal{A}}_{1}(X),{\mathcal{A}}_{2}({\mathcal{A}}_{1}(X),X))$ is an $(\epsilon_{1}+\epsilon_{2},\delta_{1}+\delta_{2})$-DP algorithm.
We can do much better if we know that the inputs are disjoint.
Fact 3.11 (Parallel composition, see Fact 2.4 in [35]).
Let ${\mathcal{A}}_{1}$ be an $(\epsilon_{1},\delta_{1})$-DP algorithm and ${\mathcal{A}}_{2}$ be an $(\epsilon_{2},\delta_{2})$-DP algorithm. Assume ${\mathcal{A}}_{1}$ and ${\mathcal{A}}_{2}$ depend on disjoint subsets of input coordinates. Then the algorithm ${\mathcal{A}}(X)=({\mathcal{A}}_{1}(X),{\mathcal{A}}_{2}({\mathcal{A}}_{1}(X),X))$ is a $(\max\{\epsilon_{1},\epsilon_{2}\},$ $\max\{\delta_{1},\delta_{2}\})$-DP algorithm.
In addition, we have the advanced composition, which improves the dependence of the number of DP algorithms to square root but compromises the term $\delta^{\prime}$.
Theorem 3.12 (Advanced composition, see Theorem 3.20 in [18]).
For all $\epsilon,\delta,\delta^{\prime}\geq 0$, the class of $(\epsilon,\delta)$-differentially private mechanisms satisfies $(\epsilon^{\prime},k\delta+\delta^{\prime})$-differential privacy under $k$-fold adaptive composition for:
$\epsilon^{\prime}=k\epsilon(e^{\epsilon}-1)+\epsilon\sqrt{2k\log(1/\delta^{\prime})}.$
Comparison of truncated Laplace, Gaussian, and Laplace mechanisms
In this paper, we use the Chebyshev inequality to bound the error, and from [30], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
The variance of Gaussian mechanism in Theorem 3.22 in [18]:
$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}\log(1.25/\delta)}{\epsilon^{2}}.$
The variance of Laplace mechanism in Fact 3.5:
$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}}{\epsilon^{2}}.$
The variance of truncated Laplace mechanism in Fact 3.7, for $c\in(0,1]$:
$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}c}{\epsilon^{2}}.$
Thus, since it has the minimum variance, we choose the truncated Laplace mechanism to design our algorithms among these popular mechanisms.
3.4 Cross-Attention
Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1. We can access $K$ and $V$ before cross-attention inference but not user input $Q$. For the attention mechanism $\mathrm{Attn}$ (Definition 1.1), matrix $AV$ can be made DP.
Now, we describe how we can use our techniques to solve this problem. Let $A_{i,j}=\phi(Q_{i},K_{j})$ for $i\in[m],j\in[n]$. Let $V_{j,k}\in\mathbb{R}$ be the $(j,k)$-th entry of $V$, for $j\in[n],k\in[d]$.
In the following discussion, we focus on $\phi$ being the exponential inner product/softmax case, so we have $\phi(Q_{i},K_{j})=\exp(\langle Q_{i},K_{j}\rangle/d)$.
By post-processing (Fact 3.9), we only need to make $AV$ DP, where the $(i,k)$-th entry of $AV$ for each $i\in[m],k\in[d]$ can be computed by
$(AV)_{i,k}=$ $~{}\sum_{j=1}^{n}\phi(Q_{i},K_{j})V_{j,k}=~{}\sum_{j=1}^{n}V_{j,k}\exp(\langle Q_{i},K_{j}\rangle/d),$
which can be viewed as a weighted softmax problem, where $V$ is the weights, $Q$ is the query, and $K$ is the dataset. Thus, we only need to add noise to $K$. Furthermore, we find that we can only handle one column of $V$, $V_{*,k}\in\mathbb{R}^{n}$, in one data structure. Therefore, we need to initialize a total of $d$ different DPTreeSoftmaxAdaptive (Alg. 4) structures, each with weights $V_{*,k}$ for $k\in[d]$.
Algorithm 1 DPTree initialization function
1:datastructure DPTree $\triangleright$ Theorem 6.1
2:members
3:     $a:\mathbb{R}^{n}$
4:     $b:\mathbb{R}^{2n-1}$
5:     $c:\mathbb{R}^{2n-1}$
6:end members
7:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1)$) $\triangleright$ Lemma B.3, Lemma 6.3
8:     $a\leftarrow a$
9:     $b[n,2n-1]\leftarrow a$
10:     for $i=n\to 2n-1$ do
11:         $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
12:     end for
13:     for $i=(\log n)\to 1$ do
14:         for $j=1\to 2^{i-1}$ do
15:              $k\leftarrow 2^{i-1}+j-1$
16:              $b[k]\leftarrow b[2k]+b[2k+1]$
17:              $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
18:         end for
19:     end for
20:end procedure
21:end datastructure
4 Main Results: DPTree
In this section, we provide the main results for our data structures: DPTree (Algorithm 1 and 2), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 4). In Section 4.1, we give our basic building block algorithm DPTree. In Section 4.2, we present ourDPTreeSoftmax algorithm that can solve the weighted softmax problem. In Section 4.3, we present our DPTreeSoftmaxAdaptive algorithm that enables DPTreeSoftmax to handle adaptive query problem.
Algorithm 2 DPTree query
1:datastructure DPTree $\triangleright$ Theorem 6.1
2:procedure Query($x\in[n],y\in[n]$) $\triangleright$ Lemma B.4, B.5, B.6
3:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $c$) by using at most $2\log n$ nodes on the path. Let Value be the above summation.
4:     return Value
5:end procedure
6:procedure TrueQuery($x\in[n],y\in[n]$)
7:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $b$) by using at most $2\log n$ nodes on the path. Let Value be the above summation.
8:     return Value
9:end procedure
10:end datastructure
4.1 DPTree Initialization and Query
We design a basic data structure DPTree (Algorithm 1 and 2) that can answer summation queries by a summation segment tree with truncated Laplace noise. The algorithm first builds a binary summation tree in an array and then adds truncated Laplace noises to each node (Algorithm 1).
In query time, we first trace from bottom nodes to find their lowest common ancestor, then report the summation by using at most $O(\log n)$ nodes on the path (Algorithm 2).
Algorithm 3 Softmax query
1:datastrucutre DPTreeSoftmax $\triangleright$ Theorem 4.2
2:members
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$triangleright Alg. 6, Theorem G.7
4:     $P:[0,\Gamma_{R,s}]^{n\times r}$triangleright Def. G.3 for $\Gamma_{R,s}$, Eq. (6) for $s$, Eq. (7) for $r$, and Lem. G.5 for $P$
5:     $w:[-R_{w},R_{w}]^{n}$
6:     $P_{wx},~{}s_{w},~{}\epsilon_{s}:\mathbb{R}$
7:end members
8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma 6.7
9:     $\epsilon_{s},~{}w,~{}P,~{}P_{wx},~{}s_{w}\leftarrow\epsilon_{s},~{}w,~{}0^{n\times r},~{}0,~{}0$
10:     for $j=1\to n$ do
11:         Compute $P(x_{j})$
12:         Compute $w_{j}\|P(x_{j})\|_{2}^{2}$
13:         $P_{wx}\leftarrow P_{wx}+w_{j}\|P(x_{j})\|_{2}^{2}$
14:         $s_{w}\leftarrow s_{w}+w_{j}$
15:         $P_{j,:}\leftarrow P(x_{j})$
16:     end for
17:     for $i=1\to r$ do
18:         $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{r\log(1/\delta^{\prime})}$, $\delta/r$) $\triangleright$ Alg. 6
19:     end for
20:end procedure
21:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma 6.7
22:     Value $\leftarrow 0$
23:     Compute $P(y)$
24:     Compute $\|P(y)\|_{2}^{2}$
25:     for $i=1\to r$ do
26:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(P(y)_{i},\alpha)$triangleright Alg. 7
27:     end for
28:     Value $\leftarrow$ $0.5\cdot(P_{wx}+s_{w}\|P(y)\|_{2}^{2}~{}-\text{Value})$
29:     return Value
30:end procedure
31:end datastrucutre
Algorithm 4 Adaptive query data structure
1:datastructure DPTreeSoftmaxAdaptive $\triangleright$ Theorem 4.4
2:members
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$triangleright Alg. 3
4:end members
[…]
6:     $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
7:     for $i=1\to l$ do
8:         ${\mathcal{D}}_{i}$.Init$(X,n,w,\epsilon/l,\delta/l,\delta^{\prime}/l,c,\epsilon_{s})$
9:     end for
10:end procedure
11:procedure DistanceQuery($y\in[0,R]^{d},\alpha\in(0,1)$)
12:     $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
13:     $r\leftarrow 0^{l}$
14:     for $i=1\to l$ do
15:         $r_{i}\leftarrow{\mathcal{D}}_{i}.\textsc{DistanceQuery}(y,\alpha)$
16:     end for
17:     return Median of $r$
18:end procedure
19:end datastructure
4.2 Softmax Activation
In this section, we present the algorithm (Algorithm 3) that can answer the weighted softmax query (Definition 4.1) and can be further used to design DP cross-attention. First, we introduce the definition of weighted softmax query, which is also an abstraction for the problem described in Section 3.4.
Definition 4.1 (Softmax query).
Let $\phi$ be the distance measurement function defined in Definition 1.1. For the dataset $X\in[0,R]^{n\times d}$ where $x_{i}^{\top}$ is the $i$-th row of $X$ and query $y\in[0,R]^{d}$, we define the weighted exponential inner product/softmax query to be:
$\sum_{i\in[n]}w_{i}\phi(x_{i},y)=\sum_{i\in[n]}w_{i}\exp(\langle x_{i},y\rangle/d)=w^{\top}\exp(Xy/d).$
Based on the Definition 4.1, we design a novel algorithm that can make weighted softmax query DP. More details can be found in Section G.
Theorem 4.2 (Softmax query, informal version of Theorem G.8).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1))$. (Algorithm 3) It takes $O(nr)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 3) It takes $O(\alpha^{-1}r\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $0.99$.
Remark 4.3.
In Theorem 4.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section 6.5. Besides, note that the error bound in Theorem 4.2 does not depend on $\delta$ but depends on $\delta^{\prime}$. The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ can reduce the error by a small constant (Fact 3.7). In practice, we set $\delta$ as a small positive constant close to 00. Please refer to the Lemma B.6 for more details.
4.3 Adaptive Query Data Structure
We can now present the theorem for our data structure that can answer weighted softmax queries with a DP guarantee and concrete error bound, and it is robust to adaptive queries. The corresponding proofs can be found in Section F and Section G.
Theorem 4.4 (Adaptive query softmax data structure, informal version of Theorem G.11).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnr)$ spaces to solve the weighted softmax query problem for the dataset $X\subset[0,R]^{d}$ and supports the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$. (Algorithm 4) It takes $O(lnr)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 4) It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
    • it is robust to adaptive query.
Remark 4.5.
We describe the parallelization of our algorithms. In the second for loop of Init and the for loop of DistanceQuery in Algorithm 3, the $r$ DPTreeDistance data structures instantiated for each coordinate are independent of each other. In addition, the for loops in Algorithm 4 can also be parallelized since the $l=O(r\log(dR/(\epsilon_{s}p_{f})))$ copies are independent. After parallelization, we have the final time complexity of Init to be $O(nr)$ and DistanceQuery to be $O(\alpha^{-1}\log^{2}n)$ in Algorithm 4 with $O(lr)$ GPU process.
In Theorem 4.4, we show that our DPTreeSoftmaxAdaptive data structure defined above have $(\epsilon,\delta)$-DP guarantee with $\alpha$ relative error and $O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}(n)+w^{\top}\exp(Xy/d))$ additive error under $O(lnr)$ memory consumption, $O(lnr)$ initialization time and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query. We can apply our data structure in cross-attention, and we show detail in the following section.
5 Main Results: Cross-Attention
In this section, we show our main result for cross-attention. Theorem 5.1 states that we can make the cross-attention module DP.
Theorem 5.1 (Cross-attention, informal version of Theorem G.12).
Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma 6.7). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. Let $\phi$ be softmax function (Definition 4.1). There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
  • We initialize $d$ data structures using Init$(K,n,V_{*,k},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$ (Algorithm 4), for $k\in[d]$. It takes $O(lnr)$ time to initialize one data structure.
  • At query time, for user input $Q$, we process one token at a time by passing the $i$-th row of $Q$, denoted $Q_{i}\in\mathbb{R}^{d}$, to DistanceQuery$(Q_{i},\alpha\in(0,1))$ (Algorithm 4) for each $i\in[m]$. It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP,
    • $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
    • $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
    • it is robust to adaptive query.
In Theorem 5.1, we use our DPTreeSoftmaxAdaptive (see Section 6.6) and guarantee that for each query token of cross-attention, the output is $(\epsilon,\delta)$-DP with $n^{-1}(\alpha+\epsilon_{s})$ relative error and $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query. More specifically, the algorithm creates $d$ DPTreeSoftmaxAdaptive data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time.
Thus, our algorithm can theoretically protect system prompts/RAG data in cross-attention as discussed in Section 1. In Section 6, we provide a detailed technical overview, and in Section 7, we will present self-attention and DP-related discussion.
6 Technique Overview
This section provides the technique overview of our paper. In Section 6.1, we analyze our DPTree data structure. In Section 6.2, we show the sensitivity of summation problem. In Section 6.3, we explain the high-level idea behind the weighted $\ell_{p}^{p}$ distance query. In Section 6.4, we show how to answer one-dimensional weighted $\ell_{1}$ distance query. In Section 6.5, we show how to answer softmax distance query using previous algorithms. In Section 6.6, we show how to handle adaptive query. By combining the results from these sections, we prove the main results in Section 4.
6.1 Summation Segment Tree
First, in order to solve the weighted distance problem, we need to have a basic DP algorithm (Algorithm 1 and 2) that can answer simple summation queries. After analyzing its DP and error in Section B, we state the data structure theorem.
Theorem 6.1 (DPTree data structure, informal version of Theorem B.1).
There is a data structure (see DPTree in Algorithm 1, 2) that uses $O(n)$ spaces to support the following operations:
  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
  • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
    • it holds with probability $0.99$.
During the design of the data structure, we found an interesting property based on the parallel composition rule of DP Fact 3.11. We will now state the lemma with its proof provided in Section B.
Lemma 6.2 (Weighted sum of disjoint interval queries, informal version of Lemma B.8).
If the following conditions hold that:
  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
  • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
Then, we have $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ (see Alg. 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$.
From Lemma 6.2, we can see that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries, as long as the sum of squared weights is finite.
6.2 Sensitivity for Range Summation Problem
Our DP summation tree data structure DPTree (Algorithm 1 and 2) requires sensitivity parameter $\Delta$. In this section, we show that for the summation problem, we have the sensitivity $\Delta=2R$ if the input $X\in[-R,R]^{n}$.
Lemma 6.3 (Sensitivity of summation).
Let $X\in[-R,R]^{n}$. We have the sensitivity $\Delta=2R$ for DPTree.Init in Algorithm 1.
Proof.
Let’s say 2 neighboring dataset $X$ and $X^{\prime}$ differ in $x_{i}$ and $x^{\prime}_{i}$ for some $i$ in the array $X$. Then for a summation problem, i.e. $f(X):=\sum_{i=1}^{n}x_{i}$, we have
$\Delta=$ $~{}\max_{X,X^{\prime}}|f(X)-f(X^{\prime})|=\max_{X,X^{\prime}}|x_{i}-x^{\prime}_{i}|=2R.$
where the first step follows from Definition 3.2, the second step follows from $X,X^{\prime}$ differ in $x_{i},x^{\prime}_{i}$, and the last step follows from each coordinate of the dataset is bounded in $[-R,R]$. ∎
6.3 Weighted $\ell_{p}^{p}$ Distance Problem
Figure 1: The visualization of how to build $c_{j}$ of rounded dataset $X\in[0,1]^{10}$ and compute the weighted $\ell_{1}$ distance. The number above each $x_{i}$ is $w_{i}$. See Algorithm 6 for details. Suppose $y=0$. Then $\sum_{i=1}^{n}w_{i}|y-x_{i}|=0.1\cdot 2.2+0.3\cdot 3.1+0.3\cdot(-2)+0.3\cdot(-3)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5+0.9\cdot(-1)+0.9\cdot 1=4.4$. And $\sum_{j=0}^{n}|k-j|c_{j}/n=0.1\cdot 2.2+0.3\cdot(-1.9)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5=4.4$. See details in Lemma C.1.
In this section, we introduce the intuition behind the method for handling the weighted $\ell_{p}^{p}$ distance problem. The formal lemmas and proofs can be found in Section C.
Given a dataset and a query point in $d$ dimensions, we round each coordinate of the data points and the query point to the nearest multiple of a small interval. We then aggregate the weights of data points that have been rounded to the same position. Finally, we compute the sum of these aggregated weights multiplied by the distances between the query point and the data points over the rounded positions. This approach makes the computation more efficient while maintaining sufficient accuracy.
We provide an example of weighted $\ell_{1}$-distance of a one-dimensional dataset consisting of 10 data points, i.e., $X\in[0,1]^{10}$ and a query $y=0$ in Figure 1.
Lemma 6.4 (Weighted $\ell_{p}^{p}$-distance high dimension, informal version of Lemma C.2).
If the following conditions hold:
  • Let data $X\in[0,R]^{n\times d}$ and $x_{i}^{\top}\in[0,R]^{d}$ be the $i$-th row of $x$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]^{d}$.
  • We round each dimension of $X$ and $y$ to an integer multiple of $R/n$.
  • Let $x_{i,k},y_{k}$ denote the $k$-th coordinates of $x_{i},y$ for $k\in[d]$.
  • Let $c_{j,k}:=\sum_{j_{0}\in S_{j,k}}w_{j_{0}}$ where the set $S_{j,k}$ is the set of index $i$ such that the corresponding $x_{i,k}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
  • After rounding, we assume that $y_{k}$ is in the $l_{k}R/n$ position for $l_{k}\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
For the weighted problem, we have
$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}.$
where $O(R/n)$ is the rounding error for each data point.
Remark 6.5.
In Lemma 6.4, we first round the dataset. This rounding simplifies the calculation by reducing the number of possible positions to consider, from real values in $[0,R]^{d}$ to the total $O(nd)$ spots. However, it also introduces an error $O(R/n)$ for one data point. Then, for one spot in the rounded dataset, we sum over the weights of that spot and multiply the corresponding distance raised to the power of $p$. Additionally, since we are dealing with $\ell_{p}^{p}$ distance, the rounding error is also raised to the power of $p$.
6.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
Based on previous discussions in Section 6.1 and 6.3, we can now describe our one-dimensional weighted $\ell_{1}$ distance data structure, DPTreeDistance, presented in Algorithm 6 and 7, which generalizes the results from [8]. Drawing from the intuition in Section 6.3, the initialization process is as follows: first, we round each data point in the dataset to the nearest multiple of a small interval and build an array that aggregates the corresponding weights. This array is then fed into our DPTree data structure in Algorithm 1. At query time, we query the DPTree to obtain the aggregated weights within a small interval and multiply these weights by the distance to the query point. Furthermore, we also introduce a relative error parameter $\alpha$ to reduce the total number of queries to $O(\log(n)/\alpha)$ instead of querying all $n$ positions. We also analyze the DP and the error bound; see details in Section D.
Theorem 6.6 (DPTreeDistance data structure, informal version of Theorem D.6).
There is a data structure DPTreeDistance (Algorithm 6,7) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 6) It takes $O(n)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 7) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$,
    • it holds with probability $0.99$.
6.5 Softmax Activation
We then describe how we extend the previous results to softmax activation, i.e. exponential inner product function (Definition 4.1). From [6], we know that softmax activation can be approximated by polynomial kernel function $P(\cdot)$ with a certain error. The following lemma shows that we can transform weighted softmax queries into polynomial kernels. More specifically, we have one term that computes the weighted $\ell_{2}^{2}$ distance, which is the place where we add DP noises. Because of the decomposability of the $\ell_{p}^{p}$ distance, i.e. $\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p}$, we can easily extend the results of Section 6.4 to handle the $\ell_{2}^{2}$ distance query. After that, we compute the term for the weighted $\ell_{2}^{2}$ norms of approximation kernel exactly. Summing all these terms, with a certain error, we can answer the softmax query. Related proofs can be found in Section G.
Lemma 6.7 (Weighted softmax approximation, informal version of Lemma G.6).
Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma G.5. Then, we can approximate the exponential inner product using the polynomial kernel function:
$w^{\top}\exp(Xy/d)=$ $~{}-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
$+~{}O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
Moreover, the vectors $P(\cdot)$ can be computed in $O(r)$ time.
6.6 Adaptive Query
We introduce how we can modify our algorithm to solve the adaptive query problem using some tools in [75]. Our approach is based on proving that our algorithm can handle any query within the query space with a certain error. Since adaptive queries must lie within this space, our algorithm can effectively handle them. In Section 6.5, we demonstrate our algorithm’s capability to answer weighted softmax distance queries with constant probability. We then use the Chernoff bound to boost the constant probability of our algorithm to a high probability. Next, we apply the notion of an $\epsilon_{0}$-net to bound all query points within the net using the union bound. Finally, we bound all points in the query space by utilizing the Lipschitz property of the weighted softmax distance function and introducing an additive error. See the proofs in Sections F and G.
Lemma 6.8 (Adaptive softmax, informal version of Lemma G.10).
If the following conditions hold:
  • Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and let $|N|$ be the size of the net $N$.
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let the relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
  • We create $l=O(\log((R/\epsilon_{0})^{r}/p_{f}))$ independent copies of the data structure $\{\textsc{DPTreeSoftmax}_{j}\}_{j=1}^{l}$ (Algorithm 3) and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
  • Let $Z(y):=w^{\top}\exp(Xy/d)$.
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
$|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
7 Discussion
How do we extend to self-attention?
As self-attention is a more fundamental module in LGMs, we would like to extend our data structure to this setting. However, the challenge we faced was the dynamic update in tree nodes for each query for self-attention, which our current analysis does not support. How we can solve this challenge is crucial, and we leave it as our future direction.
Why not add noise in some other places?
Where and how to add DP noises is an important problem to ask during the DP algorithm design. In this paper, we consider the problem of $\sum_{i=1}^{n}w_{i}\exp(\langle x_{i},y\rangle/d)$ where $y,x_{i}\in[0,R]^{d}$ and $w\in[-R_{w},R_{w}]^{n}$ (Definition 4.1). Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorihtm 1 and 2). From Lemma 6.3 and the way we initialize DPTree in Algorithm 6, we can see that the sensitivity $\Delta$ of this problem is $2R_{w}$.
A simple method for adding noise involves adding $n$ noises to a length $n$ array, with each item $w_{i}\exp(\langle x_{i},y\rangle/d)$ for $i\in[n]$. However, this approach increases the error by a factor of $n$ by basic composition (Fact 3.10) and also makes the model dependent on the number of queries. Besides, it only supports a single query and requires rebuilding the tree for each new query, rendering it impractical.
In contrast, our current noise-adding technique (Lines 11 and 17 of Algorithm 1) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$. This method also supports multiple queries, eliminating the need to rebuild the tree each time.
8 Conclusion
To our knowledge, we are the first work to provide differential privacy for cross-attention. This paper presents the DPTree data structures, which provide a differential privacy guarantee for the cross-attention module in large generative models. This is achieved by transforming the cross-attention mechanism into a weighted distance problem. Furthermore, our algorithm is robust to adaptive queries, allowing users to interact with the model arbitrarily without extracting sensitive information from the system prompts or RAG data. Our results may inspire more privacy algorithm design in large generative models.
Acknowledgement
Research is partially supported by the National Science Foundation (NSF) Grants 2023239-DMS, CCF-2046710, and Air Force Grant FA9550-18-1-0166.
Appendix
Roadmap.
The appendix is organized as follows. In Section A, we give the preliminary of our paper. In Section B, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee. In Section C, we show how to solve weighted distance problem. In Section D, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{1}$ distance problem with DP and accuracy guarantee. In Section E, we present the analysis of our DPTreeHighDim (Algorithm 8) data structure, which can address the high-dimensional $\ell_{1}$ distance problem while ensuring differential privacy and accuracy guarantees. In Section F, we show how we can handle adaptive query. In Section G, we show how to extend our algorithm to softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 4).
Appendix A More Preliminary
In Section A.1, we give the probability tools we use in the paper. In Section A.2, we provide the facts we use.
A.1 Probability Tools
In this section, we give several probability lemmas.
Lemma A.1 (Markov’s inequality).
If $x$ is a nonnegative random variable and $t>0$, we have
$\Pr[x\geq t]\leq\frac{\operatorname*{{\mathbb{E}}}[x]}{t}.$
Lemma A.2 (Chernoff bound, [13]).
Let $x_{i}$ be a Bernoulli random variable with probability $p_{i}$ of being equal to $1$ and $1-p_{i}$ of being equal to 00, and all $x_{i}$ for $i\in[n]$ are independent. Let $x=\sum_{i=1}^{n}x_{i}$. Let $\mu=\operatorname*{{\mathbb{E}}}[x]=\sum_{i=1}^{n}p_{i}$. Then, for all $\delta>0$ we have
$\Pr[x\geq(1+\delta)\mu]\leq\exp(-\delta^{2}\mu/3),$
and for all $10<\delta<$ 1
$\Pr[x\leq(1-\delta)\mu]\leq\exp(-\delta^{2}\mu/2).$
Lemma A.3 (Chebyshev’s inequality).
Let $x$ (integrable) be a random variable with finite non-zero variance $\sigma^{2}$ (and thus finite expected value $\mu$). Then for any real number $k>0$,
$\Pr[|x-\mu|\geq k\sigma]\leq\frac{1}{k^{2}}.$
A.2 Facts
Fact A.4 (Upper bound of exponential, Fact C.9 in [41]).
For $a\in\mathbb{R}$, $b\in\mathbb{R}$, $a,b\leq R$, where $R\geq 0$, we have
$|\exp(a)-\exp(b)|\leq\exp(R)|a-b|.$
Appendix B DPTree Algorithm
In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 1 and 2). In Section B.1, we give the theorem (Theorem B.1) of our data structure that can answer summation problem. In Section B.2, we improve our data structure from constant probability to high probability by applying Chernoff bound. In Section B.3, we give the analysis. In Section B.4, we show some results of our data structure if the input queries are disjoint.
B.1 Single Data Structure
We give the theorem of our DPTree data structure that can answer the summation problem with DP, accuracy, runtime guarantee.
Theorem B.1 (DPTree data structure, formal version of Theorem 6.1).
There is a data structure (see DPTree in Algorithm 1, 2) that uses $O(n)$ spaces to support the following operations:
  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
  • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
    • it holds with probability $0.99$.
Proof.
The proofs follow from combining Lemma B.3 (running time of initialization), Lemma B.4 (running time of query), Lemma B.5 (DP of query), and Lemma B.6 (error of query) together. ∎
B.2 Boost the Constant Probability to High Probability
Algorithm 5 Boost constant probability
1:datastructure DPTreeHighProb $\triangleright$ Theorem B.2
2:members
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(\log(1/\delta_{\mathrm{fail}}))}:\textsc{DPTree}$triangleright Alg. 1
4:end members
5:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1),\delta_{\mathrm{fail}}\in(0,0.01)$)
6:     for $i=1\to O(\log(1/\delta_{\mathrm{fail}}))$ do
7:         ${\mathcal{D}}_{i}$.Init$(a,n,\Delta,\epsilon/O(\log(1/\delta_{\mathrm{fail}})),\delta/O(\log(1/\delta_{\mathrm{fail}})))$
8:     end for
9:end procedure
10:procedure Query($x\in[n],y\in[n]$)
11:     $r\leftarrow 0^{O(\log(1/\delta_{\mathrm{fail}}))}$
12:     for $i=1\to O(\log(1/\delta_{\mathrm{fail}}))$ do
13:         $r_{i}\leftarrow{\mathcal{D}}_{i}.\textsc{Query}(x,y)$
14:     end for
15:     return Median of $r$
16:end procedure
17:end datastructure
We can use Chernoff bound to boost the high probability by repeating the data structure multiple times.
Theorem B.2 (High-probability).
There is a data structure (see DPTreeHighProb in Algorithm 5) that uses $O(n\log(1/\delta_{\mathrm{fail}}))$ spaces to support the following operations
  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1),\delta_{\mathrm{fail}}\in(0,0.01))$. It takes $O(n\log(1/\delta_{\mathrm{fail}}))$ time to initialize the data structure.
  • Query$(x\in[n],y\in[n])$. It takes $O(\log(n)\cdot\log(1/\delta_{\mathrm{fail}}))$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$,
    • it holds with probability $1-\delta_{\mathrm{fail}}$ for failure probability $\delta_{\mathrm{fail}}\in(0,0.01)$.
Proof.
Note that our data structure (Theorem B.1) succeeds with probability $0.99$. The success of the algorithm (Theorem B.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma A.2). By repeating the data structure $O(\log(1/\delta_{\mathrm{fail}}))$ times and taking the median of the outputs, we boost the success probability. The details are following.
To boost the success probability, we assume the query is repeated $l$ times. Let $i\in[l]$, and let $z_{i}$ denote the indicator random variable for the success of the $i$-th instance of the data structure for a single query. Let $z=\sum_{i=1}^{l}z_{i}$ be the total success times. Since $p=\Pr[z_{i}=1]=0.99$, we can have $\mu=\operatorname*{{\mathbb{E}}}[z]=\sum_{i=1}^{l}p=lp$. Note that $p=0.99$. By setting $\delta=0.1$ and using Chernoff bound from Lemma A.2, we can show
$\Pr[z\leq l/2]\leq\Pr[z\leq(1-\delta)lp]\leq\exp(-\delta^{2}lp/2).$
Note that we want $z>l/2$ (since we want at least half to succeed so we could take the median),
$\Pr[z>l/2]\geq 1-\exp(-\delta^{2}lp/2).$
To ensure that failure probability is $\delta_{\mathrm{fail}}$, we have
$\exp(-\delta^{2}lp/2)=\delta_{\mathrm{fail}}.$
We can make this hold by choosing $l=O(\log(1/\delta_{\mathrm{fail}}))$.
By the DP basic composition rule (Fact 3.10), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem B.1. ∎
B.3 Algorithm of Data Structure
In this section, we analyze the accuracy, DP, and runtime of Algorithm 1 and 2.
We first analyze the runtime.
Lemma B.3 (Runtime of initialization, Algorithm 1).
For the initialization, we have the time complexity of Algorithm 1 is $O(n)$.
Proof.
All the computations are dominated by $O(n)$ time. ∎
Lemma B.4 (Runtime of query, Algorithm 2).
For each query, we have the time complexity of Algorithm 2 is $O(\log n)$.
Proof.
Due to the property of tree, we will use at most $2\log n$ nodes in the tree, thus the running time is $O(\log n)$. ∎
We now analyze the DP.
Lemma B.5 (Privacy of query, Algorithm 2).
The output of Query (see Algorithm 2) is $(\epsilon,\delta)$-DP.
Proof.
Suppose that our dataset is $X\in[-R,R]^{n}$. Note that we only add noise into the pre-processing stage. There is no noise in the query stage. Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma 6.3). Since we add noise to each node in the tree, and each leaf node count will contribute to $\log n$ nodes, it is equivalent to our output function being in $\log n$ dimension. We will then blow up the DP parameter by $\log n$ factor. Thus, using the basic composition rule (Fact 3.10), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP. ∎
We now analyze the accuracy.
Lemma B.6 (Accuracy of query, Algorithm 2).
Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 3.7, we have the error of Query(see Algorithm 2) output is upper bounded by:
$O(\epsilon^{-1}\Delta\log^{3/2}n).$
with probability $0.99$.
Proof.
[…]
$z=\sum_{i=1}^{O(\log n)}z_{i}$
where $z_{i}\sim\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$, and every $z_{i}$ are independent to each other.
We know $\mu=\operatorname*{{\mathbb{E}}}[z]=0$ since $\operatorname*{{\mathbb{E}}}[z_{i}]=0$. From Fact 3.7, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
Therefore, we can show
$\operatorname*{\mathrm{Var}}[z]=$ $~{}\operatorname*{\mathrm{Var}}[\sum_{i=1}^{O(\log n)}z_{i}]$
$=$ $~{}\sum_{i=1}^{O(\log n)}\operatorname*{\mathrm{Var}}[z_{i}]$
$=$ $~{}O(c\epsilon^{-2}\Delta^{2}\log^{3}n)$ (1)
where the first step follows from definition of $z$, the second step follows from every $z_{i}$ are independent to each other, and the last step follows from $\operatorname*{\mathrm{Var}}[z_{i}]=O(c\epsilon^{-2}\Delta^{2}\log^{2}n)$.
Note that we wish to bound $|z|=|\textsc{Query}(S)-\textsc{TrueQuery}(S)|$ as our error.
Using Lemma A.3, we can have
$\Pr[|z|\geq k\sigma]\leq\frac{1}{k^{2}}.$
We know that $\sigma=\sqrt{\operatorname*{\mathrm{Var}}[z]}=O(c^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$. Picking $k=10$, we have
$\Pr[|z|<10\sigma]\geq 0.99.$
Thus, we conclude that error is bounded by $O(c^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)=O(\epsilon^{-1}\Delta\log^{3/2}n)$ (since $c\in(0,2]$) with probability $0.99$. ∎
B.4 Disjoint Intervals
In this section, we show some interesting results for our DPTree data structure if the input queries are disjoint.
Lemma B.7 (Sum of disjoint interval queries).
If the following conditions hold that:
  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
Then, we have $\sum_{j=1}^{t}\textsc{Query}(S_{j})$ (see Algorithm 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
$O(t^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$.
Proof.
From Lemma B.5, we know that $\mathrm{DPTree}$.Query returns a $(\epsilon,\delta)$-DP output. Then, from Fact 3.11 and the disjoint intervals in Algorithm 7, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
Let $\textsc{TrueQuery}(S_{j})$ denote the true interval query answer returned by DPTree.TrueQuery in Algorithm 2 for interval $S_{j}$. Let $\textsc{Query}(S_{j})$ denote the noised interval query answer returned by DPTree.Query in Algorithm 2 for interval $S_{j}$. Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}z_{j}$. From the proof of Lemma B.6, we know $z_{j}$ is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$ and the variance is bounded by
$\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
Since the intervals $S_{j}$ are disjoint, they are independent to each other. Then, we have
$\operatorname*{\mathrm{Var}}[z]=$ $~{}\operatorname*{\mathrm{Var}}[\sum_{j=1}^{t}z_{j}]$
$=$ $~{}\sum_{j=1}^{t}\operatorname*{\mathrm{Var}}[z_{j}]$
$=$ $~{}O(t\epsilon^{-2}\Delta^{2}\log^{3}n)$
where the first step follows from definition of $z$, the second step follows from the intervals are disjoint, and the last step follows from $\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{2}n)$.
Note that we wish to bound $|z|$ as our error.
Using Lemma A.3, we can have error bounded by
$O(t^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$. ∎
Moreover, this can be generalized to weighted sum of queries.
Lemma B.8 (Weighted sum of disjoint interval queries, formal version of Lemma 6.2).
If the following conditions hold that:
  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
  • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
Then, we have $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ (see Algorithm 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$.
Proof.
The DP proof is the same as in the proof of Lemma B.7.
Let $\textsc{TrueQuery}(S_{j})$ and $\textsc{Query}(S_{j})$ be same in the proof of Lemma B.7 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}a_{j}z_{j}$. From the proof of Lemma B.7, we know the variance of $z_{j}$ is bounded by
$\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
Since the intervals $S_{j}$ are disjoint, they are independent to each other. Then, we have
$\operatorname*{\mathrm{Var}}[z]=$ $~{}\operatorname*{\mathrm{Var}}[\sum_{j=1}^{t}a_{j}z_{j}]$
$=$ $~{}\sum_{j=1}^{t}\operatorname*{\mathrm{Var}}[a_{j}z_{j}]$
$=$ $~{}\sum_{j=1}^{t}a_{j}^{2}\operatorname*{\mathrm{Var}}[z_{j}]$
$=$ $~{}\sum_{j=1}^{t}a_{j}^{2}\cdot O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
$=$ $~{}O(a\epsilon^{-2}\Delta^{2}\log^{3}n)$
where the first step follows from the definition of $z$, the second step follows from the intervals are disjoint, the third step follows from the $\operatorname*{\mathrm{Var}}[az]=a^{2}\operatorname*{\mathrm{Var}}[z]$ for a random variable $z$ and a constant $a$, the fourth step follows from the $\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{2}n)$, and the last step follows from $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
Note that we wish to bound $|z|$ as our error.
Using Lemma A.3, we can have error bounded by
$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
with probability $0.99$. ∎
Appendix C Weighted $\ell_{p}^{p}$ Distance
In this section, we introduce how to handle weighted $\ell_{p}^{p}$ distance problem in the high level idea. In Section C.1, we show how to solve one dimensional weighted problem. In Section C.2, we show how to solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset.
Suppose we have the original data $X\in[0,R]^{n}$ and weight $w\in\mathbb{R}^{n}$ and query $y\in[0,R]$. We want to compute the weighted $\ell_{1}$-distance, i.e.
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|.$
For data in $d$-dimension, due to the decomposability of $\ell_{p}^{p}$ distance, our problem will be: given $x_{i}\in[0,R]^{d}$ and $w_{i}\in\mathbb{R}$ for $i\in[n]$, and $y\in[0,R]^{d}$, we can compute
$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=$ $~{}\sum_{j=1}^{d}\sum_{i=1}^{n}w_{i}\cdot|y_{j}-x_{i,j}|^{p}$
where $x_{i,j},y_{j}$ means the $j$-th coordinates of $x_{i},y$ for $j\in[d]$.
Therefore, we can solve one dimension problem first, and then the high dimension case can be solved automatically.
C.1 One Dimensional Weighted Distance
Now we can give the lemma for weighted distance of dataset.
Lemma C.1 (Weighted distance one dimension).
If the following conditions hold:
  • Let data $X\in[0,R]^{n}$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]$.
  • We round $X$ and $y$ to an integer multiple of $R/n$.
  • Let $c_{j}=\sum_{j_{0}\in S_{j}}w_{j_{0}}$ where set $S_{j}$ is the set of index $i$ such that the corresponding $x_{i}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$.
  • After rounding, we assume $y$ is in the $kR/n$ position for $k\in\{0,1,2,\dots,n\}$.
For the weighted problem, we have
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|=\sum_{j=0}^{n}(|k-j|R/n+O(R/n))c_{j}.$
Moreover, we have
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(|k-j|R/n+O(R/n))^{p}c_{j}.$
where $O(R/n)$ is the rounding error for each data point.
Proof.
For each $i$, we have:
$w_{i}\cdot|y-x_{i}|=w_{i}\cdot(\frac{|k-j|R}{n}+O(\frac{R}{n})).$
where $O(R/n)$ is the rounding error introduced by each data point, since each data point will be at most $O(R/n)$ away from its true position.
We can construct $c_{j}$ by
$c_{j}=\sum_{j_{0}\in S_{j}}w_{j_{0}}$
set $S_{j}$ is the set of index $i$ such that the corresponding $x_{i}$ is rounded to $jR/n$. Moreover, $c_{j}$ can be negative.
Summing over all $i$ and grouping by $j$, we get:
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|=\sum_{j=0}^{n}(\frac{|k-j|R}{n}+O(\frac{R}{n}))c_{j}.$
The total rounding error will be $O(R)$ because we have $n$ data points, each with an error of at most $O(R/n)$.
Moreover, we have
$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(\frac{|k-j|R}{n}+O(\frac{R}{n}))^{p}c_{j}.$
C.2 High Dimensional Weighted Distance
Finally, we can solve the problem of weighted distance for $d$-dimensional dataset.
Lemma C.2 (Weighted $\ell_{p}^{p}$-distance high dimension, formal version of Lemma 6.4).
If the following conditions hold:
  • Let data $X\in[0,R]^{n\times d}$ and $x_{i}^{\top}\in[0,R]^{d}$ be the $i$-th row of $x$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]^{d}$.
  • We round each dimension of $X$ and $y$ to an integer multiple of $R/n$.
  • Let $x_{i,k},y_{k}$ denote the $k$-th coordinates of $x_{i},y$ for $k\in[d]$.
  • Let $c_{j,k}:=\sum_{j_{0}\in S_{j,k}}w_{j_{0}}$ where set $S_{j,k}$ is the set of index $i$ such that the corresponding $x_{i,k}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
  • After rounding, we assume $y_{k}$ is in the $l_{k}R/n$ position for $l_{k}\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
For the weighted problem, we have
$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}.$
where $O(R/n)$ is the rounding error for each data point.
Proof.
We can show
$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=$ $~{}\sum_{k=1}^{d}\sum_{i=1}^{n}w_{i}\cdot|y_{k}-x_{i,k}|^{p}$
$=$ $~{}\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}$
where the first step follows from decomposability of $\ell_{p}^{p}$-distance by dimension, the second step follows from Lemma C.1.
Appendix D One-Dimensional Weighted $\ell_{1}$ Distance Query
In this section, we generalize the algorithms in [8] to weighted distance. Here, we compute the problem of one-dimensional weighted $\ell_{1}$ distance query i.e. $\sum_{i\in[n]}w_{i}|y-x_{i}|$ for a given query $y\in[0,R]$, weights $w\in[-R_{w},R_{w}]^{n}$ and dataset $X\subset[0,R]$ and $n=|X|$. In Section D.1, we analyze the runtime of our algorithm. In Section D.2, we analyze the DP and accuracy of our algorithm. In Section D.3, we give the theorem for our DPTreeDistance data structure.
Algorithm 6 Pre-processing data structure
1:datastructure DPTreeDistance $\triangleright$ Theorem D.6
2:members
3:     ${\mathcal{D}}:$ DPTree $\triangleright$ Alg. 1
4:     $X:[0,R]^{n}$
5:     $w:[-R_{w},R_{w}]^{n}$
6:end members
7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma C.1
8:     $X,w,a\leftarrow X,w,0^{n+1}$
9:     for $i=1\to n$ do
10:         $j\leftarrow\textsc{Round}(x_{i},n)$triangleright $x_{i}\in X$ for $i\in[n]$
11:         $a_{j}\leftarrow a_{j}+w_{i}$
12:     end for
13:     ${\mathcal{D}}$.$\textsc{Init}(a,n+1,2R_{w},\epsilon,\delta)$triangleright Alg. 1, Lemma 6.3
14:end procedure
15:procedure Round($x\in[0,R]$, $n\in\mathbb{N}_{+}$)
16:     Let $j\in\{0,1,2,\dots n-1\}$ denote the integer such that $jR/n\leq x<(j+1)R/n$
17:     if $|x-(j+1)R/n|\leq|x-jR/n|$ then
18:         $j\leftarrow j+1$
19:     end if
20:     return $j$
21:end procedure
22:end datastructure
Algorithm 7 One dimensional weighted $\ell_{1}$ distance query
1:datastructure DPTreeDistance $\triangleright$ Theorem D.6
2:procedure DistanceQuery($y\in[0,R]$, $\alpha\in(0,1)$) $\triangleright$ Lemma D.2, Lemma D.4, Lemma D.5
3:     $y\leftarrow\textsc{Round}(y,n)\cdot(R/n)$triangleright Alg. 6
4:     Value $\leftarrow 0$
5:     for $j=0,1,...,O(\log(n)/\alpha)$ do
6:         $l_{j}\leftarrow\textsc{Round}(y+\frac{R}{(1+\alpha)^{j+1}},n)$
7:         $r_{j}\leftarrow\textsc{Round}(y+\frac{R}{(1+\alpha)^{j}},n)$triangleright Consider the points to the right of $y$
8:         Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$triangleright Alg. 2
9:     end for
10:     for $j=0,1,...,O(\log(n)/\alpha)$ do
11:         $l_{j}\leftarrow\textsc{Round}(y-\frac{R}{(1+\alpha)^{j}},n)$
12:         $r_{j}\leftarrow\textsc{Round}(y-\frac{R}{(1+\alpha)^{j+1}},n)$triangleright Consider the points to the left of $y$
13:         Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$triangleright Alg. 2
14:     end for
15:     Return Value
16:end procedure
17:end datastructure
D.1 Runtime Analysis
We first analyze the runtime.
Lemma D.1 (Runtime of initialization, Algorithm 6).
For the initialization, we have the time complexity of Init in Algorithm 6 is $O(n)$.
Proof.
In the initialization of Init, the computations need $O(n)$ time to compute the count and $O(\log n)$ time to build the tree. Thus, total time is $O(n)$. ∎
Lemma D.2 (Runtime of DistanceQuery, Algorithm 7).
For the $\ell_{1}$ distance query, we have the time complexity of DistanceQuery in Algorithm 7 is $O(\alpha^{-1}\log^{2}n)$.
Proof.
In DistanceQuery, the computations need $O(\log n)$ time to compute one value from DPTree.Query and this process need to be repeated $O(\alpha^{-1}\log n)$ times. ∎
Remark D.3.
In Line 8 and 13 of Algorithm 7, we use $R/(1+\alpha)^{j}$ to approximate the distance of each data point to the query in Lemma C.1, i.e. $|k-j|R/n$. This will introduce $\alpha$ relative error but also reduce the numbers of iteration from $O(n)$ to $O(\log(n)/\alpha)$.
D.2 Privacy and Accuracy Analysis
We show the DP.
Lemma D.4 (Privacy of DistanceQuery, Algorithm 7).
The output of DistanceQuery (Algorithm 7) is $(\epsilon,\delta)$-DP.
Proof.
From Lemma B.5, we know that $\mathrm{DPTree}$.Query returns a $(\epsilon,\delta)$-DP output. We observe that intervals in Algorithm 7 are disjoint. Then, following the same logic in the proof of Lemma B.8, we can conclude that the value returned is $(\epsilon,\delta)$-DP. ∎
We now analyze the accuracy of the algorithm.
Lemma D.5 (Accuracy of DistanceQuery, Algorithm 7).
If the following conditions are satisfied:
  • Let $X\in[0,R]^{n}$ be a dataset consisting of $n$ one-dimensional numbers, with weights $w\in[-R_{w},R_{w}]^{n}$.
  • Let $\alpha\in(0,1)$ represent the relative error parameter utilized in Algorithm 7.
  • Let $\widetilde{A}$ denote the output of the DistanceQuery in Algorithm 7.
  • Let $A_{*}:=\sum_{i\in[n]}w_{i}|y-x_{i}|$ represent the true distance query value for a specific query $y$.
Then with probability $0.99$, we have
$|\widetilde{A}-A_{*}|\leq\alpha A_{*}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n).$
Proof.
[…]
$\widehat{A}:=$ $~{}\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot\textsc{TrueQuery}(S_{j}).$
Since TrueQuery returns the sum of the corresponding weights, it aligns with the true answer $A_{*}:=\sum_{i\in[n]}w_{i}|y-x_{i}|$. Thus, we have
$|\widehat{A}-A_{*}|\leq\alpha\cdot A_{*},$
because for all $j$, the distances between $y$ and different points in $X$ vary only by a multiplicative factor of $(1+\alpha)$.
Next we show the additive error. Let $\textsc{Query}(S_{j})$ denote the noised interval query answer returned by DPTree.Query in Algorithm 2. Algorithm 7 outputs $\widetilde{A}=\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot\textsc{Query}(S_{j})$. We wish to bound
$|\widehat{A}-\widetilde{A}|\leq$ $~{}|\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot(\textsc{TrueQuery}(S_{j})-\textsc{Query}(S_{j}))|.$
Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$, which from Algorithm 1 we can see this is the sum of $O(\log n)$ independent truncated Laplace random variables.
From Lemma B.8, we only need to show that the series $\frac{1}{(1+\alpha)^{j}}$ for $j\in\{0,1,\dots,O(\log(n)/\alpha)\}$ square converges to $1/\alpha$, since $R$ is a constant.
We can show
$\sum_{j=0}^{O(\log(n)/\alpha)}\frac{1}{(1+\alpha)^{2j}}\leq$ $~{}\sum_{j=0}^{\infty}\frac{1}{(1+\alpha)^{2j}}$
$\leq$ $~{}\sum_{j=0}^{\infty}\frac{1}{(1+\alpha)^{j}}$
$=$ $~{}\frac{1}{1-\frac{1}{1+\alpha}}$
$=$ $~{}1+\frac{1}{\alpha}$
$=$ $~{}O(1/\alpha)$
where the first step follows from we extend the finite sum to infinite sum, the second step follows from $\frac{1}{(1+\alpha)^{2j}}\leq\frac{1}{(1+\alpha)^{j}}$, the third step follows from the closed form of geometric sum, the fourth step follows from simple algebra, and the last step follows from $\alpha\in(0,1)$.
Then from the proof of Lemma B.8, we can know that the variance is given by
$O(\frac{R^{2}R_{w}^{2}\log^{3}n}{\alpha\epsilon^{2}})$ (2)
since the sensitivity $\Delta=2R_{w}$ from Lemma 6.3.
Using Lemma A.3, we can have additive error bounded by
$O(\frac{R\cdot R_{w}\log^{3/2}n}{\epsilon\sqrt{\alpha}}).$
with probability $0.99$. ∎
D.3 One Dimension Single Data Structure
We therefore have the data structure that can solve weighted $\ell_{1}$-distance problem.
Theorem D.6 (DPTreeDistance data structure, formal version of Theorem 6.6).
There is a data structure DPTreeDistance (Algorithm 6,7) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 6) It takes $O(n)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 7) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$,
    • it holds with probability $0.99$.
Proof.
The proofs follow from combining Lemma D.1 (running time of initialization), Lemma D.2 (running time of query), Lemma D.4 (DP of query), and Lemma D.5 (error of query) together. ∎
Appendix E High-Dimensional Weighted $\ell_{1}$ Query
In this section, we show how we can solve the high dimensional weighted $\ell_{1}$ distance problem, generalizing results from [8]. In Section E.1, we give the analysis of Algorithm 8. In Section E.2, we give the theorem of our DPTreeHighDim data structure.
Algorithm 6,7 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{1}$ distance function. We construct $d$ separate one-dimensional distance query data structures, each corresponding to a coordinate projection of the dataset.
E.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
We now give the analysis of our Algorithm 8 for high dimensional weighted $\ell_{1}$-distance query.
Algorithm 8 High-dimensional weighted $\ell_{1}$ distance query
1:datastrucutre DPTreeHighDim $\triangleright$ Theorem E.3
2:members
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$triangleright Alg. 6
4:     $X:[0,R]^{n\times d}$
5:     $w:[-R_{w},R_{w}]^{n}$
6:end members
7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$) $\triangleright$ Lemma C.2
8:     $X\leftarrow X$
9:     $w\leftarrow w$
10:     for $i=1\to d$ do
11:         $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 6
12:     end for
13:end procedure
14:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma E.1, Lemma E.2
15:     Value $\leftarrow 0$
16:     for $i=1\to d$ do
17:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(y_{i},\alpha)$triangleright Alg. 7
18:     end for
19:     return Value
20:end procedure
21:end datastrucutre
Lemma E.1 (Privacy of DistanceQuery, Algorithm 8).
If the following conditions hold
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$.
  • Let $c\in(0,0.1)$ be a small constant and $A$ be the output of DistanceQuery in Algorithm 8, where each one-dimensional algorithm is configured to be $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP (see Line 11).
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ represent the true distance query value.
  • Let $\epsilon\leq O(\log(1/\delta^{\prime}))$.
Then, we have the output of DistanceQuery (Algorithm 8) is $(\epsilon,\delta+\delta^{\prime})$-DP.
Proof.
The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem 3.12. Our algorithm instantiate each one-dimensional data structure with $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP total $d$ times.
From advanced composition in Theorem 3.12, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
$O(c\epsilon\sqrt{2d\log(1/\delta^{\prime})}/\sqrt{d\log(1/\delta^{\prime})})=O(\epsilon)$
and the final failure probability parameter be:
$d\delta/d+\delta^{\prime}=\delta+\delta^{\prime}.$
Lemma E.2 (Accuracy of DistanceQuery, Algorithm 8).
If the following conditions hold
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$.
  • Let $c\in(0,0.1)$ be a small constant and $A$ be the output of DistanceQuery in Algorithm 8, where each one-dimensional algorithm is configured to be $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP (see Line 11).
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ represent the true distance query value.
With probability $0.99$, we have
$|A-A_{*}|\leq\alpha A_{*}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}d\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
Proof.
Let $A_{i}$ be the $i$-th dimension output returned by $\mathcal{D}_{i}$ in Algorithm 8. Let $A_{*,i}$ be the true distance query value in the $i$-th dimension. Observe that $A_{*}=\sum_{i=1}^{d}A_{*,i}$ and $A=\sum_{i=1}^{d}A_{i}$.
We follow the similar idea in the proof of Lemma D.5. Let $z_{j,i}$ be the random variables that represent $z_{j}$ (used in the proof of Lemma D.5) for the $i$-th coordinate. We can observe that the overall error across $d$ coordinates can be upper bounded by
$|\sum_{i=1}^{d}\sum_{j=0}^{O(\log(n)/\alpha)}\frac{Rz_{j,i}}{(1+\alpha)^{j}}|$
where each $z_{j,i}$ is the sum of $O(\log n)$ truncated Laplace random variables independent to others. With $\epsilon$ scaled down by $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$ and $\delta$ scaled down by $\delta/d$, the variance of each individual dimension is given by (see Eq. (2))
$O(\alpha^{-1}\epsilon^{-2}dR^{2}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
Thus, the total variance for $d$ instantiated data structures is then
$O(\alpha^{-1}\epsilon^{-2}d^{2}R^{2}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
Finally, from Lemma A.3, we have the additive error given by
$O(\alpha^{-1/2}\epsilon^{-1}dRR_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
E.2 High Dimension Single Data Structure
We have the data structure that can solve weighted $\ell_{1}$-distance problem in $d$-dimensional data.
Theorem E.3 (DPTreeHighDim data structure).
There is a data structure DPTreeHighDim (Algorithm 8) that uses $O(nd)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1))$. (Algorithm 8) It takes $O(nd)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 8) It takes $O(\alpha^{-1}d\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$,
    • $|z-\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}|\leq\alpha\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}d\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $0.99$.
Proof.
For the runtime analysis, since we loop data structure DPTreeDistance $d$ times, an additional $d$ factor will appear for both initialization and query time complexity. The DP is proved by Lemma E.1. The accuracy is proved by Lemma E.2. ∎
Appendix F Adaptive Query
In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from [75]. Our idea is that, if we can prove that our algorithm can solve any query in the query space with certain error. Then, since adaptive query must lie in this space, we can handle adaptive query. In Section F.1, we show how we can boost the constant probability of our algorithm to high probability. In Section F.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net. In Section F.3, we show how we can bound all points in the query space by introducing an additive error. In Section F.4, we examine the effects of different norms on our adaptive query proof.
First, from Theorem E.3, given query $y\in[0,R]^{d},\alpha\in(0,1)$ we have DistanceQuery$(y,\alpha)$ that can solve $d$-dimension weighted $\ell_{1}$-distance problem with constant probability $0.99$. Now we show how to improve it to solve adaptive query problem.
F.1 Boost the Constant Probability to High Probability
We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma A.2.
Lemma F.1 (Using Chernoff bound to boost the probability).
If the following conditions hold:
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
  • We create $l=O(\log(1/p_{f}))$ independent copies of data structure DPTreeHighDim and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ be the true answer.
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then for each fixed query point $y$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by
$\alpha A_{*}+B$
with probability $1-p_{f}$.
Proof.
By basic composition Fact 3.10, we prove the DP. Similar to the proof of Theorem B.2, we prove the error by Chernoff bound (Lemma A.2). ∎
F.2 From Each Fixed Query Point to All On-net Points
In this section, we build $\epsilon_{0}$-net and generalize from each fixed query point to all on-net points.
Definition F.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in [85]).
We define $N$ be $\ell_{p}$ $\epsilon_{0}$-net of :=${\mathcal{B}}:=\{q\in[0,R]^{d}\}$ such that, for every point $q$ in ${\mathcal{B}}$, there exists $y\in N$ satisfying $\|y-q\|_{p}\leq\epsilon_{0}$.
Fact F.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$. We have $|N|\leq(5R/\epsilon_{0})^{d}$.
Fact F.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in [88]).
Let $N$ be the $\ell_{2}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$. We have $|N|\leq(5R/\epsilon_{0})^{d}$.
Fact F.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in [43]).
Let $N$ be the $\ell_{1}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$. We have $|N|\leq(5R\sqrt{d}/\epsilon_{0})^{d}$.
Lemma F.6 (From for each query point to for all points in net).
If the following conditions hold:
  • Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$.
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
  • We create $l=O(\log(|N|/p_{f}))$ independent copies of data structure DPTreeHighDim and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ be the true answer.
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by
$\alpha A_{*}+B.$
Proof.
By basic composition Fact 3.10, we prove the DP. From Lemma F.1, we know for each $y\in N$, the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}/|N|$.
Then, by union bound, with probability $1-p_{f}$, the error of all $|N|$ query points in the net $y\in N$ is upper bounded by $\alpha A_{*}+B$. ∎
F.3 From Net Points to All Points
In this section, we show how to generalize points from net to all points in the query space. Since adaptive query must lie in this space, we complete the proof of adaptive query.
Lemma F.7 (Lipschitz of query function).
If the following conditions hold:
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let $Z(y):=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$.
  • Let $L=nR_{w}$.
Then, we have $Z(y)$ is $L$-Lipschitz (note that we have $\ell_{1}$ Lipschitz here).
Proof.
We can show
$|Z(y)-Z(\widetilde{y})|=$ $~{}|\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}-\sum_{i\in[n]}w_{i}\|\widetilde{y}-x_{i}\|_{1}|$
$\leq$ $~{}\sum_{i\in[n]}|w_{i}|\cdot|\|y-x_{i}\|_{1}-\|\widetilde{y}-x_{i}\|_{1}|$
$\leq$ $~{}\sum_{i\in[n]}|w_{i}|\cdot\|y-\widetilde{y}\|_{1}$
$=$ $~{}nR_{w}\cdot\|y-\widetilde{y}\|_{1}$
where the first step follows from definition of $Z(y)$, the second step follows from triangular inequality, the third step follows from reverse triangular inequality, the fourth step follows from $w\in[-R_{w},R_{w}]^{n}$. ∎
Lemma F.8 (From points in net to all points in query space).
If the following conditions hold:
  • Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$.
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
  • We create $l=O(\log((R/\epsilon_{0})^{d}/p_{f}))$ independent copies of data structure $\{\textsc{DPTreeHighDim}_{j}\}_{j=1}^{l}$ and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeHighDim}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
  • Let $Z(y):=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$, where $Z(y)$ is $L$-Lipschitz with $L=nR_{w}$.
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
$|f(y)-Z(q)|\leq\alpha Z(q)+B+2Ld\epsilon_{0}.$
Proof.
By basic composition Fact 3.10, we prove the DP.
We define an event $E$ such that:
$~{}\forall y\in N$
$~{}|f(y)-Z(y)|\leq\alpha Z(y)+B.$
From Lemma F.1, with $l=O(\log(|N|/p_{f}))$ we know
$\Pr[\mathrm{event}~{}E~{}\mathrm{holds}]\geq 1-p_{f}$
We can show
$l=$ $~{}O(\log(|N|/p_{f})$
$=$ $~{}O(\log((R/\epsilon_{0})^{d}/p_{f})$
where the first step follows from definition of $l$, the second step follows from Fact F.3.
We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition F.2), for each $q\notin N$, there exists $y\in N$ such that
$\|y-q\|_{\infty}\leq\epsilon_{0}$ (3)
We know
$|Z(y)-Z(q)|\leq$ $~{}L\cdot\|y-q\|_{1}$
$\leq$ $~{}L\cdot d\|y-q\|_{\infty}$
$\leq$ $~{}L\cdot d\epsilon_{0}$ (4)
where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (3).
Using the on-net query $y$ to answer the off-net query $q$, for any $q\notin N$, we have
$|f(y)-Z(q)|\leq$ $~{}|f(y)-Z(y)|+|Z(q)-Z(y)|$
$\leq$ $~{}|f(y)-Z(y)|+L\cdot d\cdot\epsilon_{0}$
$\leq$ $~{}\alpha Z(y)+B+L\cdot d\cdot\epsilon_{0}$
$\leq$ $~{}\alpha Z(q)+B+2L\cdot d\cdot\epsilon_{0}$ (5)
where the first step follows from triangular inequality, the second step follows from Eq. (F.3), the third step follows from Lemma F.6, and the last step follows from Eq. (F.3).
Thus, we complete the proof. ∎
Therefore, even adaptive queries can be answered accurately, since any adaptive query can be assumed in ${\mathcal{B}}$.
F.4 Effect of Different Norms on the Result
In the above proof, we have two different measure spaces, i.e. $\ell_{\infty}$ distance of $\epsilon_{0}$-net (Definition F.2) and $\ell_{1}$ Lipschitz (Lemma F.7).
One might ask, will the norm we choose in two spaces have an impact on the final result? We can show that the norm we choose currently is sufficient to use.
For different norms, the only differences in the proofs will be Lipschitz smoothness in Eq. (F.3) and the cardinality of $\epsilon_{0}$-net, i.e. $|N|$ in Fact F.3.
Lemma F.9.
If we use $\ell_{\infty}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
$l=O(d\log(nR/p_{f}))$
Proof.
If we use $\ell_{\infty}$ to bound the distance to net, Eq. (F.3) is:
$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
$\leq$ $~{}nR_{w}\cdot d\|y-q\|_{\infty}$
$\leq$ $~{}nR_{w}\cdot d\epsilon_{0}$
where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{\infty}$ $\epsilon_{0}$-net.
Then, Eq. (F.3) is
$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot d\cdot\epsilon_{0}$
For $\ell_{\infty}$ distance, we have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact F.3.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot d\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
$l=$ $~{}O(\log(|N|/p_{f})$
$=$ $~{}O(\log((R/\epsilon_{0})^{d}/p_{f})$
$=$ $~{}O(\log((nR)^{d}/p_{f}))$
$=$ $~{}O(d\log(nR/p_{f}))$
where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
Lemma F.10.
If we use $\ell_{2}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
$l=O(d\log(nR/p_{f}))$
Proof.
If we use $\ell_{2}$ to bound the distance to net, Eq. (F.3) changes to be:
$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
$\leq$ $~{}nR_{w}\cdot\sqrt{d}\cdot\|y-q\|_{2}$
$\leq$ $~{}nR_{w}\cdot\epsilon_{0}\sqrt{d}$
where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq\sqrt{d}\cdot\|x\|_{2}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{2}$ $\epsilon_{0}$-net.
Then, Eq. (F.3) changes to be
$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot\epsilon_{0}\sqrt{d}$
For $\ell_{2}$ distance, we also have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact F.4.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\sqrt{d}\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
$l=$ $~{}O(\log(|N|/p_{f})$
$=$ $~{}O(\log((R/\epsilon_{0})^{d}/p_{f})$
$=$ $~{}O(\log((nR)^{d}/p_{f}))$
$=$ $~{}O(d\log(nR/p_{f}))$
where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
Lemma F.11.
If we use $\ell_{1}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
$l=O(d\log(nRd/p_{f}))$
Proof.
If we use $\ell_{1}$ to bound the distance to net, Eq. (F.3) changes to be:
$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
$\leq$ $~{}nR_{w}\cdot\epsilon_{0}$
where the first step follows from Lemma F.7, and the last step follows from $\ell_{1}$ $\epsilon_{0}$-net.
Then, Eq. (F.3) changes to be
$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot\epsilon_{0}$
For $\ell_{1}$ distance, we have $|N|\leq(5R\sqrt{d}/\epsilon_{0})^{d}$.
We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
$l=$ $~{}O(\log(|N|/p_{f})$
$=$ $~{}O(\log((R\sqrt{d}/\epsilon_{0})^{d}/p_{f})$
$=$ $~{}O(\log((nR\sqrt{d})^{d}/p_{f}))$
$=$ $~{}O(d\log(nRd/p_{f}))$
where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
From the above analysis, we can show that $\ell_{\infty}$ or $\ell_{2}$ $\epsilon_{0}$-net is slightly better than $\ell_{1}$ $\epsilon_{0}$-net.
  • $\ell_{\infty}$ $\epsilon_{0}$-net, Lemma F.9: we have $l=O(d\log(nR/p_{f}))$.
  • $\ell_{2}$ $\epsilon_{0}$-net, Lemma F.10: we have $l=O(d\log(nR/p_{f}))$.
  • $\ell_{1}$ $\epsilon_{0}$-net, Lemma F.11: we have $l=O(d\log(nRd/p_{f}))$.
Thus, the norm we choose for $\epsilon_{0}$-net is sufficient good.
Appendix G Softmax Activation
In this section, we introduce how we extend previous $\ell_{1}$ distance results to the softmax activation function, which is the most widely used distance measure in attention mechanism based models. In order to have softmax activation, we need the distance measure function $\phi$ in Definition 1.1 to be the softmax function in Definition 4.1.
In Section G.1, we show how to extend to the softmax distance function in Lemma G.6. In Section G.2, we show how to adjust our algorithms. In Section G.3, we extend our algorithm to be robust to adaptive query. In Section G.4, we give the proof of our main result Theorem 5.1.
G.1 Exponential Inner Product
In this section, we show how we obtain the softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from [6].
Definition G.1 (Definition 3.1 in [6]).
Let $r\geq 1$ denote a positive integer. Let $\epsilon\in(0,0.1)$ denote an accuracy parameter. Given a matrix $A\in\mathbb{R}^{n\times n}_{\geq 0}$, we say $\widetilde{A}\in\mathbb{R}^{n\times n}_{\geq 0}$ is an $(\epsilon,r)$-approximation of $A$ if
  • $\widetilde{A}=U_{1}\cdot U_{2}^{\top}$ for some matrices $U_{1},U_{2}\in\mathbb{R}^{n\times r}$ (i.e., $\widetilde{A}$ has rank at most $r$), and
  • $|\widetilde{A}_{i,j}-A_{i,j}|\leq\epsilon\cdot A_{i,j}$ for all $(i,j)\in[n]^{2}$.
Lemma G.2 (Lemma 3.4 in [6]).
Suppose $Q,K\in\mathbb{R}^{n\times d}$, with $\|Q\|_{\infty}\leq R$, and $\|K\|_{\infty}\leq R$. Let $A:=\exp(QK^{\top}/d)\in\mathbb{R}^{n\times n}$. For accuracy parameter $\epsilon\in(0,0.1)$, there is a positive integer $s$ bounded above by
$s=O\Big{(}\max\Big{\{}\frac{\log(1/\epsilon)}{\log(\log(1/\epsilon)/R)},R^{2}\Big{\}}\Big{)},$ (6)
and a positive integer $r$ bounded above by
$r\leq{2s+2d\choose 2s}$ (7)
such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition G.1) of $A\in\mathbb{R}^{n\times n}$. Furthermore, the matrices $U_{1}$ and $U_{2}$ defining $\widetilde{A}$ can be computed in $O(n\cdot r)$ time.
Here we consider the vector version of Lemma G.2.
Definition G.3.
We define $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$.
Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in [6].
Remark G.4.
We use $\Gamma_{R,s}$ to denote the value range of our polynomial kernel methods function, i.e., $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$. The factorial term in $\Gamma_{R,s}$ comes from Taylor approximation coefficients. We take the maximum overall $s$ order approximation terms to get the upper bound of our value range.
Lemma G.5 (Polynomial approximation).
For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in [6] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Then, for any $x,y\in[0,R]^{d}$, we have
$|P(x)^{\top}P(y)-\exp(x^{\top}y/d)|\leq\epsilon_{s}\cdot\min\{\exp(x^{\top}y/d),P(x)^{\top}P(y)\}$
Furthermore, the vectors $P(x)$ and $P(y)$ can be computed in $O(r)$ time.
Proof.
Let $n=1$. The proof follows from directly applying Lemma G.2. ∎
Using the results from [6] above, we can extend our results to softmax activation.
Lemma G.6 (Weighted softmax approximation, formal version of Lemma 6.7).
Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma G.5. Then we can approximate exponential inner product using polynomial kernel function:
$w^{\top}\exp(Xy/d)=$ $~{}-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
$+~{}O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s})$
Moreover, the vectors $P(\cdot)$ can be computed in $O(r)$ time.
Proof.
From Lemma G.5, we can use polynomial kernel to approximate the softmax function:
$w^{\top}\exp(Xy/d)=$ $~{}\sum_{i\in[n]}w_{i}P(x_{i})^{\top}P(y)+O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma G.5.
Then, we can show
$2\sum_{i\in[n]}w_{i}P(x_{i})^{\top}P(y)=$ $~{}-\sum_{i\in[n]}w_{i}\|P(x_{i})-P(y)\|_{2}^{2}+\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
$=$ $~{}-\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
where the first step follows from $\|x-y\|_{2}^{2}=\|x\|_{2}^{2}+\|y\|_{2}^{2}-2\langle x,y\rangle$, and the second step follows $\|x\|_{2}^{2}=\sum_{j=1}^{d}|x_{j}|^{2}$ for $x\in\mathbb{R}^{d}$. ∎
G.2 Algorithm Modifications
Based on Lemma G.6, we can now extend our DP algorithms to handle softmax activation. First, we need to construct $P(y)$ and $P(x_{i})$ for $i\in[n]$, each costing $O(r)$ time. Then, for the second term in Lemma G.6, i.e. $\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$, we don’t need to add DP noises in it; instead, we calculate this term exactly, preprocess it, and store the results in the algorithm. For the first term, $-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}$, we can adjust our high dimensional DP distance query algorithm to solve it.
Due to the decomposability of $\ell_{p}^{p}$ norm, i.e.
$\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p},$
we can compute $\ell_{2}^{2}$ norm easily (see details in Lemma C.2). We then show how to extend our one dimensional $\ell_{1}$ distance algorithm (Algorithm 6 and 7) to $\ell_{2}^{2}$ distance with minor modifications.
Theorem G.7 (DPTreeDistance $\ell_{2}^{2}$ distance).
With $\alpha$ scaled down by a factor of $2$ and all Query instead multiplied by $R^{2}/(1+\alpha/2)^{2j}$ in Lines 8 and 13 of Algorithm 7, i.e., from
  • Lines 8 and 13: Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$
to
  • Lines 8 and 13: Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R^{2}}{(1+\alpha/2)^{2j}}$.
The data structure DPTreeDistance (Algorithm 6,7) uses $O(n)$ spaces to solve weighted $\ell_{2}^{2}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 6) It takes $O(n)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 7)
    It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}$,
    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}|\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}+O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$,
    • it holds with probability $0.99$.
Proof.
The proof is similar to that of Theorem D.6, except that now our additive error includes $R$ increased by a power of $2$, i.e., from $O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$ to $O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$. ∎
Now we can give our result that can answer softmax query.
Theorem G.8 (Softmax query, formal version of Theorem 4.2).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition G.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1))$. (Algorithm 3) It takes $O(nr)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 3) It takes $O(\alpha^{-1}r\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
      ${}~{}~{}~{}~{}~{}~{}~{}+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $0.99$.
Proof.
The DP proof is the same as the proof of Lemma E.1.
We then show the time complexity. From Lemma G.6, we know that constructing $P(\cdot)$ requires $O(r)$ time. In the first for loop of Init, the dominating time consumption is $O(nr)$. The second for loop also has a time complexity of $O(nr)$. Therefore, the total time complexity for Init is $O(nr)$. In the DistanceQuery function, constructing $P(y)$ takes $O(r)$ time. Within the for loop, it requires $O(\alpha^{-1}r\log^{2}n)$. Thus, the total time complexity for DistanceQuery is $O(\alpha^{-1}r\log^{2}n)$.
The space complexity is $O(nr)$, since storing the $n\times r$ matrix $P$ is the dominating factor.
The proof of the error follows from the triangle inequality by combining the errors in Lemma G.6 and Theorem G.7.
G.3 Adaptive Softmax
In this section, we show how to make Algorithm 3 robust to adaptive query. We follow the same idea from Section F. We notice that, in the softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$-distance in Section F. Therefore, we need to recalculate Lipschitz constant first.
Lemma G.9 (Lipschitz of weighted softmax).
If the following conditions hold:
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let $Z(y):=w^{\top}\exp(Xy/d)$.
  • Let $L=nd^{-1/2}RR_{w}\exp(R^{2})$.
Then, we have $Z(y)$ is $L$-Lipschitz (note that we have $\ell_{1}$ Lipschitz here).
Proof.
We can show
$|Z(y)-Z(\widetilde{y})|=$ $~{}|\sum_{i\in[n]}w_{i}\exp(x_{i}^{\top}y/d)-\sum_{i\in[n]}w_{i}\exp(x_{i}^{\top}\widetilde{y}/d)|$
$\leq$ $~{}\sum_{i\in[n]}|w_{i}|\cdot|\exp(x_{i}^{\top}y/d)-\exp(x_{i}^{\top}\widetilde{y}/d)|$
$\leq$ $~{}\sum_{i\in[n]}|w_{i}|\exp(R^{2})|x_{i}^{\top}y/d-x_{i}^{\top}\widetilde{y}/d|$
$\leq$ $~{}\sum_{i\in[n]}|w_{i}|\exp(R^{2})\|x_{i}\|_{2}\cdot\|y-\widetilde{y}\|_{2}/d$
$\leq$ $~{}nR_{w}\exp(R^{2})\sqrt{d}R\cdot\|y-\widetilde{y}\|_{2}/d$
$\leq$ $~{}nd^{-1/2}RR_{w}\exp(R^{2})\|y-\widetilde{y}\|_{1}$
where the first step follows from definition of $Z(y),Z(\widetilde{y})$, the second step follows from triangular inequality, the third step follows from Fact A.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$. ∎
Then we can show how to extend our algorithm to be robust to adaptive query.
Lemma G.10 (Adaptive softmax, formal version of Lemma 6.8).
If the following conditions hold:
  • Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$.
  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
  • We create $l=O(\log((R/\epsilon_{0})^{r}/p_{f}))$ independent copies of data structure $\{\textsc{DPTreeSoftmax}_{j}\}_{j=1}^{l}$ (Algorithm 3) and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
  • Let $Z(y):=w^{\top}\exp(Xy/d)$, where $Z(y)$ is $L$-Lipschitz with $L=nd^{-1/2}RR_{w}\exp(R^{2})$.
  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
$|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
Proof.
The proof follows from the same idea as the proof of Lemma F.8, except that we use Theorem G.8 and the Lipschitz in Lemma G.9. ∎
Theorem G.11 (Adaptive query softmax data structure, formal version of Theorem 4.4).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}$ be defined in Definition G.3. Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter, and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnr)$ spaces to solve weighted softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$. (Algorithm 4) It takes $O(lnd)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 4) It takes $O(\alpha^{-1}ld\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
      ${}~{}~{}~{}~{}~{}~{}~{}+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
    • it is robust to adaptive query.
Proof.
We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma G.10. We know the additive error introduced by adaptive query is $E_{a}:=O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0})$ and the relative error introduced by polynomial kernel approximation is $E_{p}:=w^{\top}\exp(Xy/d)\cdot\epsilon_{s}$. It can be shown that:
$E_{p}:=$ $~{}w^{\top}\exp(Xy/d)\cdot\epsilon_{s}$
$\leq$ $~{}\epsilon_{s}\|w\|_{2}\cdot\|\exp(Xy/d)\|_{2}$
$=$ $~{}O(nR_{w}\epsilon_{s}\exp(R^{2}))$
where the first step follows from definition of $E_{p}$, the second step follows from Cauchy–Schwarz inequality, and the last step follows from $w\in[-R_{w},R_{w}]^{n}$, $X\in[0,R]^{n\times d}$, and $y\in[0,R]^{d}$.
Picking $\epsilon_{0}=\Theta(\frac{\epsilon_{s}}{\sqrt{d}R})$, we can hide the error of adaptive query $E_{a}$ in $E_{p}$. Thus, we have
$l=$ $~{}O(\log((R/\epsilon_{0})^{r}/p_{f}))$
$=$ $~{}O(\log((\sqrt{d}R^{2}/\epsilon_{s})^{r}/p_{f}))$
$=$ $~{}O(r\log(dR/(\epsilon_{s}p_{f})))$
where the first step comes from the definition of $l$, the second step comes from picking $\epsilon_{0}=\Theta(\frac{\epsilon_{s}}{\sqrt{d}R})$, and the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
G.4 Proof of main result
In this section, we give the proof of our main result of Theorem 5.1.
Theorem G.12 (Cross-attention, formal version of Theorem 5.1).
Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma 6.7). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. Let $\phi$ be softmax function (Definition 4.1). There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
  • We initialize $d$ data structures using Init$(K,n,V_{*,k},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$ (Algorithm 4), for $k\in[d]$. It takes $O(lnr)$ time to initialize one data structure.
  • At query time, for user input $Q$, we process one token at a time by passing the $i$-th row of $Q$, denoted $Q_{i}\in\mathbb{R}^{d}$, to DistanceQuery$(Q_{i},\alpha\in(0,1))$ (Algorithm 4) for each $i\in[m]$. It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP,
    • $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
    • $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
    • it is robust to adaptive query.
Proof.
From Section 3.4, we know that we can make matrix $AV$ in cross-attention computation DP. Next, from Theorem 4.4, for $i\in[m],j\in[n],k\in[d]$, we have $(AV)_{i,k}$ is $(\epsilon,\delta+\delta^{\prime})$-DP and also robust to adaptive query. For $D$, we compute the exact true value. By the post-processing property of DP (Fact 3.9), our cross-attention is DP.
Let $(AV)_{i,k}$ be the true value and $\widetilde{(AV)}_{i,k}$ be the noisy value. Then, from Theorem 4.4, we have
$|(AV)_{i,k}-\widetilde{(AV)}_{i,k}|\leq(\alpha+\epsilon_{s})\cdot(AV)_{i,k}+O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n).$ (8)
For $D_{i,i}$, we can show
$D_{i,i}=(A\cdot{\bf 1}_{n})_{i}=\sum_{j=1}^{n}\exp(\langle Q_{i},K_{j}\rangle/d)\geq n$ (9)
because $\langle Q_{i},K_{j}\rangle\geq 0$ for bounded $Q,K$.
Finally, we can show the error of one entry is bounded by
$|(D^{-1}AV)_{i,k}-(D^{-1}\widetilde{AV})_{i,k}|=$ $~{}|D_{i,i}^{-1}((AV)_{i,k}-\widetilde{(AV)}_{i,k})|$
$=$ $~{}|D_{i,i}^{-1}|\cdot|((AV)_{i,k}-\widetilde{(AV)}_{i,k})|$
$\leq$ $~{}n^{-1}(\alpha+\epsilon_{s})\cdot(AV)_{i,k}$
$~{}+O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}(n))$
where the first step follows from definition, the second step follows from simple algebra, and the last step follows from Eq.(8) and (9). ∎

### v2

@@
 Differential Privacy of Cross-Attention with Provable Guarantee
-Jiuxiang Gu jigu@adobe.com.
-Adobe Research.
 Yingyu Liang yingyul@hku.hk.
@@
 University of Wisconsin-Madison.
-Zhao Song zsong@adobe.com.
-Adobe Research.
+Zhao Song magic.linuxkde@gmail.com.
+The Simons Institute for the Theory of Computing at the University of California, Berkeley.
 Yufa Zhou yufazhou@seas.upenn.edu.
 University of Pennsylvania.
-Cross-attention has become a fundamental module nowadays in many important artificial intelligence applications, e.g., retrieval-augmented generation (RAG), system prompt, guided stable diffusion, and many so on.
-Ensuring cross-attention privacy is crucial and urgently needed because its key and value matrices may contain sensitive information about companies and their users, many of which profit solely from their system prompts or RAG data.
+Cross-attention has become a fundamental module nowadays in many important artificial intelligence applications, e.g., retrieval-augmented generation (RAG), system prompt, guided stable diffusion, and many more.
+Ensuring cross-attention privacy is crucial and urgently needed because its key and value matrices may contain sensitive information about model providers and their users.
 In this work, we design a novel differential privacy (DP) data structure to address the privacy security of cross-attention with a theoretical guarantee.
@@
 Then, our data structure requires $\widetilde{O}(ndr^{2})$ memory consumption with $\widetilde{O}(nr^{2})$ initialization time complexity and $\widetilde{O}(\alpha^{-1}r^{2})$ query time complexity for a single token query.
-In addition, our data structure can guarantee that the user query is $(\epsilon,\delta)$-DP with $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}R^{2s}R_{w}r^{2})$ additive error and $n^{-1}(\alpha+\epsilon_{s})$ relative error between our output and the true answer.
+In addition, our data structure can guarantee that the process of answering user query satisfies $(\epsilon,\delta)$-DP with $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}R^{2s}R_{w}r^{2})$ additive error and $n^{-1}(\alpha+\epsilon_{s})$ relative error between our output and the true answer.
 Furthermore, our result is robust to adaptive queries in which users can intentionally attack the cross-attention system.
-To our knowledge, this is the first work to provide DP for cross-attention.
-We believe it can inspire more privacy algorithm design in large generative models (LGMs).
+To our knowledge, this is the first work to provide DP for cross-attention and is promising to inspire more privacy algorithm design in large generative models (LGMs).
 1 Introduction
@@
 One fundamental technique used in LGMs is cross-attention [97], which is an essential module in retrieval-augmented generation (RAG) [68], system prompt, guided stable diffusion, and many so on.
-In RAG, to be more professional, the LGMs can answer user input queries by using a domain-specific database under cross-attention, which may contain specific privacy data and knowledge so that the LGMs gain additional power.
+In RAG, to be more professional, the LGMs answer user input queries by using a domain-specific database under cross-attention, which may contain specific privacy data and knowledge so that the LGMs gain additional power.
 For system prompts, based on cross-attention, some customized long prompts, e.g., user information or concrete rules, are concatenated before user input to follow human instructions better, which are commonly used in ChatGPT [34], Claude3 [4] and other commercial LGMs.
@@
 These findings highlight the critical importance of robust privacy protections in LGMs, making privacy not just essential but an urgent issue that demands immediate attention.
-To fundamentally preserve cross-attention privacy, we can borrow the powerful tools from differential privacy (DP) [17], which provides measurable privacy and combines with statistical machine learning seamlessly [73].
+To fundamentally preserve cross-attention privacy, we borrow the powerful tools from differential privacy (DP) [19], which provides measurable privacy and combines with statistical machine learning seamlessly [84].
 Thus, in this work, we would like to ask and answer the following question,
 How can we use differential privacy to protect the security of cross-attention in LGMs?
-Our work demonstrates that the cross-attention computation is equivalent to computing the weighted distance problem.
-Definition 1.1 (Cross-attention).
-Let $n$ and $m$ be the token length of the system prompt/RAG data and input query, respectively.
+Our work demonstrates that the Softmax cross-attention computation is equivalent to computing the weighted distance problem.
+Definition 1.1 (Softmax cross-attention).
+Let $n$ and $m$ be the token length of the data and input query, respectively.
 Let $d$ be the feature dimension.
-Let $\phi:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be any function that measures vector similarity.
-Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is defined as
+Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Softmax Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is
 $\mathrm{Attn}(Q,K,V):=D^{-1}AV,$
-where $A\in\mathbb{R}^{m\times n}$ and $D\in\mathbb{R}^{m\times m}$ is a diagonal matrix, i.e.,
-$A_{i,j}:=\phi(Q_{i},K_{j}),\quad\text{and}\quad D:=\mathrm{diag}(A{\bf 1}_{n}).$
+where $A\in\mathbb{R}^{m\times n}$ satisfies $A_{i,j}:=\exp(\langle Q_{i},K_{j}\rangle/d)$ for any $i\in[m],j\in[n]$ ($Q_{i}$ and $K_{j}$ denote the $i$-th and $j$-th rows of $Q$ and $K$, respectively) and $D:=\mathrm{diag}(A{\bf 1}_{n})\in\mathbb{R}^{m\times m}$ is a diagonal matrix.
+Note that $\mathsf{Softmax}(QK^{\top})=D^{-1}A\in\mathbb{R}^{m\times n}$ in Definition 1.1, which is the standard function used in transformers, and usually, we call it as attention matrix.
 Our main theorem, presented below, provides a robust solution of cross-attention, ensuring privacy and accuracy guarantees.
 Theorem 1.2 (Main result; Informal version of Theorem 3.1).
-Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1.
+Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1.
 Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter.
-Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma 6.7).
+Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma C.7).
 Let $l=\widetilde{O}(r)$.
-Let $\phi$ be the softmax distance function (Definition 4.1).
-Then, our Alg.
-4 requires $O(lndr)$ memory to make cross-attention DP with $O(lnr)$ initialization time and $\widetilde{O}(\alpha^{-1}lr)$ query time, such that with probability $1-p_{f}$, the output is $(\epsilon,\delta)$-DP and robust to adaptive query with relative error $n^{-1}(\alpha+\epsilon_{s})$ and additive error $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}lR^{2s}R_{w}r)$.
-Our main technique in Theorem 1.2 makes cross-attention differentially private by using the polynomial kernel approximation method and transforming it into a weighted distance problem.
-Then, the above problem can be solved by summing over weighted distances (depending on the value embedding) between the query embedding and the key embedding.
+Then, our Algorithm 1 requires $O(lndr)$ memory with $O(lnr)$ initialization time and $\widetilde{O}(\alpha^{-1}lr)$ query time, such that with probability $1-p_{f}$, the output process of cross-attention satisfies $(\epsilon,\delta)$-DP and is robust to adaptive query with relative error $n^{-1}(\alpha+\epsilon_{s})$ and additive error $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}lR^{2s}R_{w}r)$.
+Our main technique in Theorem 1.2 ensures that cross-attention is differentially private by using the polynomial kernel approximation method and transforming it into a weighted distance problem.
+We then solve the problem by summing over weighted distances (depending on the value embedding) between the query embedding and the key embedding.
 We build a data structure for weighted Softmax queries in Section 4.3, and we extend this data structure to handle adaptive queries using the $\epsilon_{0}$-net/metric entropy argument in Section 4.4.
-Our contributions can be summarized as follows:
-  • We demonstrate that cross-attention computations are equivalent to the weighted distance problem (Section 3.4).
-  • We design a novel algorithm (Alg. 3) that can privately answer weighted softmax queries with high probability and a concrete accuracy bound.
-  • Our algorithm (Alg.
-4) can handle multiple queries and is robust against adaptive query attacks (Theorem 4.4), meaning that potential attackers cannot intentionally extract information of system prompts/RAG data.
+Furthermore, our error decreases as the input token length grows, diminishing both the relative and additive errors to zero.
+Our contributions are as follows:
+  • We demonstrate that cross-attention computations are equivalent to the weighted distance problem (Section 3).
+  • We design a novel algorithm (Algorithm 3) that privately answers weighted Softmax queries with high probability and a concrete accuracy bound.
+  • Our algorithm (Algorithm 1) handles multiple cross-attention queries and is robust against adaptive query attacks (Theorem 3.1), meaning that potential attackers cannot intentionally extract information of system prompts/RAG data.
 To our knowledge, this is the first work to utilize DP to protect prompts in LGMs with theoretically provable guarantees.
-While some papers have explored protecting user/system prompts with DP [24, 66], they are primarily empirical and lack theoretical guarantees.
-Additionally, many others are working on protecting private datasets by applying DP to the fine-tuning stage of LGMs training [7, 77, 59, 96, 63, 81], which diverges from our work.
+While some have explored protecting user/system prompts with DP [26, 77], they are primarily empirical and lack theoretical guarantees.
+Additionally, many others are working on protecting private datasets by applying DP to the fine-tuning stage of LGMs [9, 88, 63, 113, 74, 92], which diverges from our work.
 The strength of DP lies in its strong, unambiguous, and concrete definition of privacy, enabling algorithm designs with provable privacy and accuracy analysis.
 Therefore, we believe that the theoretical aspects of DP applications in LGMs remain a highly impactful direction, and we aim to pave the way for further exploration in this area.
-Roadmap.
-In Section 2, we provide a review of related literature.
-In Section 3, we present the preliminary of differential privacy (DP) and cross-attention.
-In Section 4, we outline the main results of our algorithms.
-In Section 5, we present the main result of our cross-attention theorem (Theorem 5.1).
-In Section 6, we offer an overview of our techniques.
-In Section 7, we discuss DP-related topics and potential extensions.
-In Section 8, we conclude our paper.
-2 Related Work
-2.1 Differential Privacy Guarantee Analysis
-Ever since [17] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [20, 61, 98, 73, 94].
-DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [22, 3, 57, 49, 35, 8, 10, 23, 14, 46, 67, 68, 53, 58, 26, 27, 56, 15, 11, 16, 25, 42, 37, 60, 36, 21, 84, 19, 91, 83, 44].
-Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [18].
-For example, truncated Laplace mechanism [30] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions.
-2.2 Differential Privacy in Data Structure and Attention
+1.1 Related Work
+Differential Privacy in Data Structure and Attention.
 Differential privacy (DP) is a flourishing and powerful technique that has enormous applications in the topic of private machine learning.
@@
 Recently, numerous works have aimed to improve this method by integrating DP in both the pre-training and fine-tuning stages of LGMs [113, 74, 31, 9, 92, 76, 88, 116, 63].
-To protect user/system prompts, [24] provides a survey on both DP and non-DP methods.
+To protect user/system prompts, EW [26] provides a survey on both DP and non-DP methods.
 In the use of LGMs, prompting methods almost become a standard way for inference [91].
 Given the billions of prompt interactions daily, ensuring privacy is essential [77].
-2.3 Cross-Attention in System Prompt, RAG, Stable Diffusion and More
-Cross-attention [86], first introduced in language translation, is a widely used technique in many advanced AI systems.
-For example, Stable Diffusion [76] and SORA [70] employ cross-attention as a core module for a text-to-image conditional generation.
-This technique is also utilized by other multimodal models [40], including Imagen [79] and Diffusion Transformer [74].
-In the realm of text-to-image editing, [48] analyzes and controls the cross-attention module to enable editing without requiring additional training.
-Furthermore, [97] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas.
-In addition, Retrieval Augmented Generation (RAG) [62, 9, 45], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module.
-Cross-attention also has other applications.
-[71] demonstrates that the prompt-tuning [39] task can be formulated as cross-attention, while [12] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation.
-Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [89], such as spatial reasoning [87], mathematical reasoning [38], in-context learning ability [82], compositional ability [92], few-shot adaptation ability [78, 93], and so on.
+We refer readers to Appendix A for more related works.
+Roadmap.
+In Section 2, we present the preliminary of differential privacy (DP) and cross-attention.
+In Section 3, we present the main result of our cross-attention theorem (Theorem 3.1).
+In Section 4, we outline the main results of our algorithms.
+In Section 5, we discuss DP-related topics and potential extensions.
+In Section 6, we conclude our paper.
 2 Preliminary
-In this section, we give the preliminary of differential privacy (DP) and cross-attention.
-In Section 3.1, we describe the notations we use.
-In Section 3.2, we give several definitions related to DP.
-In Section 3.3, we give the DP facts we use in the paper.
-In Section 3.4, we show that the cross-attention module is equivalent to the weighted distance problem.
+In this section, we give the preliminary of differential privacy (DP) and cross-attention. In Section 2.1, we describe the notations. In Section 2.2, we give definitions related to DP.
 2.1 Notations
@@
 2.2 Differential Privacy Definitions
-In this section, we give several definitions related to differential privacy (DP). We refer the reader to [18] for more background and details on DP.
+In this section, we give several definitions related to differential privacy (DP). We refer the reader to DR [20] for more background and details on DP.
 Definition 2.1 (Neighboring dataset).
@@
 Definition 2.2 (Sensitivity).
-The sensitivity of a function $f:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is defined by:
-$\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n},\|X-X^{\prime}\|_{1}=1}\|f(X)-f(X^{\prime})\|_{1}.$
+The sensitivity of a function $f:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is defined by: $\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n},\|X-X^{\prime}\|_{1}=1}\|f(X)-f(X^{\prime})\|_{1}.$
 Definition 2.3 ($(\epsilon,\delta)$-DP).
@@
 When $\delta=0$, the algorithm is said to have pure differential privacy.
-For the Laplace mechanism, we have the following definitions.
-Definition 3.4 (Laplace distribution).
-We use $\mathrm{Lap}(b)$ to denote the pdf:
-$p(z)=\frac{1}{2b}\exp(-\frac{|z|}{b}).$
-Fact 3.5.
-For $z\sim\mathrm{Lap}(b)$, $\operatorname*{{\mathbb{E}}}[z]=0$, and $\operatorname*{\mathrm{Var}}[z]=2b^{2}$.
-Furthermore, if $b=\Delta/\epsilon$, we have $\operatorname*{\mathrm{Var}}[z]=2\Delta^{2}/\epsilon^{2}$.
+We mainly use the truncated Laplace mechanism, which has the following definitions.
 Definition 2.4 (Truncated Laplace distribution).
 We use $\mathrm{TLap}(\Delta,\epsilon,\delta)$ to denote the Truncated Laplace distribution with pdf proportional to $\exp(-\epsilon|z|/\Delta)$ on the region $[-B,B]$, where $B=\frac{\Delta}{\epsilon}\cdot\log(1+\frac{\exp(\epsilon)-1}{2\delta})$.
-Fact 3.7 (Theorem 3 in [30]).
+Fact 2.5 (Theorem 3 in GDGK [32]).
 Let $z$ denote a $\mathrm{TLap}(\Delta,\epsilon,\delta)$ random variable. Then we have $\operatorname*{{\mathbb{E}}}[z]=0$, and
@@
 Furthermore, if $\delta=0$, we have $\operatorname*{\mathrm{Var}}[z]=2\Delta^{2}/\epsilon^{2}$, meaning truncated Laplacian mechanism will be reduced to the standard Laplacian mechanism.
-Lemma 3.8 (Laplace mechanism, [18, 30], see Lemma 2.2 in [3]).
+Lemma 2.6 (Laplace mechanism, [20, 32], see Lemma 2.2 in AIMN [3]).
 Given a numeric function $f$ that takes a dataset $X$ as the input, and has sensitivity $\Delta$, the mechanism that outputs $f(X)+z$ where $z\sim\mathrm{Lap}(\Delta/\epsilon)$ is $(\epsilon,0)$-DP.
@@
 Moreover, the truncated Laplace mechanism is always accuracy up to error $B$.
-3.3 DP Facts
-In this section, we present several facts about differential privacy (DP). We first state the post-processing property, which means, in an algorithm, if one step is DP, all the following steps are DP.
-Fact 3.9 (Post-processing, see Fact 2.1 in [35]).
-Let ${\mathcal{A}}_{1}$ be an $(\epsilon,\delta)$-DP algorithm and ${\mathcal{A}}_{2}$ be a (randomized) post-processing algorithm.
-Then the algorithm ${\mathcal{A}}(X)={\mathcal{A}}_{2}({\mathcal{A}}_{1}(X))$ is still an $(\epsilon,\delta)$-DP algorithm.
-If we have many DP algorithms, we need a composition rule. The most straightforward composition is the basic/sequential composition rule.
-Fact 3.10 (Basic composition, see Fact 2.3 in [35]).
-Let ${\mathcal{A}}_{1}$ be an $(\epsilon_{1},\delta_{1})$-DP algorithm and ${\mathcal{A}}_{2}$ be an $(\epsilon_{2},\delta_{2})$-DP algorithm.
-Then ${\mathcal{A}}(X)=({\mathcal{A}}_{1}(X),{\mathcal{A}}_{2}({\mathcal{A}}_{1}(X),X))$ is an $(\epsilon_{1}+\epsilon_{2},\delta_{1}+\delta_{2})$-DP algorithm.
-We can do much better if we know that the inputs are disjoint.
-Fact 3.11 (Parallel composition, see Fact 2.4 in [35]).
-Let ${\mathcal{A}}_{1}$ be an $(\epsilon_{1},\delta_{1})$-DP algorithm and ${\mathcal{A}}_{2}$ be an $(\epsilon_{2},\delta_{2})$-DP algorithm.
-Assume ${\mathcal{A}}_{1}$ and ${\mathcal{A}}_{2}$ depend on disjoint subsets of input coordinates.
-Then the algorithm ${\mathcal{A}}(X)=({\mathcal{A}}_{1}(X),{\mathcal{A}}_{2}({\mathcal{A}}_{1}(X),X))$ is a $(\max\{\epsilon_{1},\epsilon_{2}\},$ $\max\{\delta_{1},\delta_{2}\})$-DP algorithm.
-In addition, we have the advanced composition, which improves the dependence of the number of DP algorithms to square root but compromises the term $\delta^{\prime}$.
-Theorem 3.12 (Advanced composition, see Theorem 3.20 in [18]).
-For all $\epsilon,\delta,\delta^{\prime}\geq 0$, the class of $(\epsilon,\delta)$-differentially private mechanisms satisfies $(\epsilon^{\prime},k\delta+\delta^{\prime})$-differential privacy under $k$-fold adaptive composition for:
-$\epsilon^{\prime}=k\epsilon(e^{\epsilon}-1)+\epsilon\sqrt{2k\log(1/\delta^{\prime})}.$
-Comparison of truncated Laplace, Gaussian, and Laplace mechanisms
-In this paper, we use the Chebyshev inequality to bound the error, and from [30], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
-The variance of Gaussian mechanism in Theorem 3.22 in [18]:
-$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}\log(1.25/\delta)}{\epsilon^{2}}.$
-The variance of Laplace mechanism in Fact 3.5:
-$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}}{\epsilon^{2}}.$
-The variance of truncated Laplace mechanism in Fact 3.7, for $c\in(0,1]$:
-$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}c}{\epsilon^{2}}.$
-Thus, since it has the minimum variance, we choose the truncated Laplace mechanism to design our algorithms among these popular mechanisms.
-3.4 Cross-Attention
-Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1.
-We can access $K$ and $V$ before cross-attention inference but not user input $Q$.
-For the attention mechanism $\mathrm{Attn}$ (Definition 1.1), matrix $AV$ can be made DP.
-Now, we describe how we can use our techniques to solve this problem.
-Let $A_{i,j}=\phi(Q_{i},K_{j})$ for $i\in[m],j\in[n]$.
-Let $V_{j,k}\in\mathbb{R}$ be the $(j,k)$-th entry of $V$, for $j\in[n],k\in[d]$.
-In the following discussion, we focus on $\phi$ being the exponential inner product/softmax case, so we have $\phi(Q_{i},K_{j})=\exp(\langle Q_{i},K_{j}\rangle/d)$.
-By post-processing (Fact 3.9), we only need to make $AV$ DP, where the $(i,k)$-th entry of $AV$ for each $i\in[m],k\in[d]$ can be computed by
-$(AV)_{i,k}=$ $~{}\sum_{j=1}^{n}\phi(Q_{i},K_{j})V_{j,k}=~{}\sum_{j=1}^{n}V_{j,k}\exp(\langle Q_{i},K_{j}\rangle/d),$
-which can be viewed as a weighted softmax problem, where $V$ is the weights, $Q$ is the query, and $K$ is the dataset.
-Thus, we only need to add noise to $K$.
-Furthermore, we find that we can only handle one column of $V$, $V_{*,k}\in\mathbb{R}^{n}$, in one data structure.
-Therefore, we need to initialize a total of $d$ different DPTreeSoftmaxAdaptive (Alg.
-4) structures, each with weights $V_{*,k}$ for $k\in[d]$.
-Algorithm 1 DPTree initialization function
-1:datastructure DPTree $\triangleright$ Theorem 6.1
-2:members
-3:     $a:\mathbb{R}^{n}$
-4:     $b:\mathbb{R}^{2n-1}$
-5:     $c:\mathbb{R}^{2n-1}$
-6:end members
-7:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1)$) $\triangleright$ Lemma B.3, Lemma 6.3
-8:     $a\leftarrow a$
-9:     $b[n,2n-1]\leftarrow a$
-10:     for $i=n\to 2n-1$ do
-11:         $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
-12:     end for
-13:     for $i=(\log n)\to 1$ do
-14:         for $j=1\to 2^{i-1}$ do
-15:              $k\leftarrow 2^{i-1}+j-1$
-16:              $b[k]\leftarrow b[2k]+b[2k+1]$
-17:              $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
-18:         end for
-19:     end for
-20:end procedure
-21:end datastructure
-4 Main Results: DPTree
-In this section, we provide the main results for our data structures: DPTree (Algorithm 1 and 2), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 4).
-In Section 4.1, we give our basic building block algorithm DPTree.
-In Section 4.2, we present ourDPTreeSoftmax algorithm that can solve the weighted softmax problem.
-In Section 4.3, we present our DPTreeSoftmaxAdaptive algorithm that enables DPTreeSoftmax to handle adaptive query problem.
-Algorithm 2 DPTree query
-1:datastructure DPTree $\triangleright$ Theorem 6.1
-2:procedure Query($x\in[n],y\in[n]$) $\triangleright$ Lemma B.4, B.5, B.6
-3:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $c$) by using at most $2\log n$ nodes on the path.
-Let Value be the above summation.
-4:     return Value
-5:end procedure
-6:procedure TrueQuery($x\in[n],y\in[n]$)
-7:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $b$) by using at most $2\log n$ nodes on the path.
-Let Value be the above summation.
-8:     return Value
-9:end procedure
-10:end datastructure
-4.1 DPTree Initialization and Query
-We design a basic data structure DPTree (Algorithm 1 and 2) that can answer summation queries by a summation segment tree with truncated Laplace noise.
-The algorithm first builds a binary summation tree in an array and then adds truncated Laplace noises to each node (Algorithm 1).
-In query time, we first trace from bottom nodes to find their lowest common ancestor, then report the summation by using at most $O(\log n)$ nodes on the path (Algorithm 2).
-Algorithm 3 Softmax query
-1:datastrucutre DPTreeSoftmax $\triangleright$ Theorem 4.2
-2:members
-3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$triangleright Alg. 6, Theorem G.7
-4:     $P:[0,\Gamma_{R,s}]^{n\times r}$triangleright Def. G.3 for $\Gamma_{R,s}$, Eq. (6) for $s$, Eq. (7) for $r$, and Lem. G.5 for $P$
-5:     $w:[-R_{w},R_{w}]^{n}$
-6:     $P_{wx},~{}s_{w},~{}\epsilon_{s}:\mathbb{R}$
-7:end members
-8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma 6.7
-9:     $\epsilon_{s},~{}w,~{}P,~{}P_{wx},~{}s_{w}\leftarrow\epsilon_{s},~{}w,~{}0^{n\times r},~{}0,~{}0$
-10:     for $j=1\to n$ do
-11:         Compute $P(x_{j})$
-12:         Compute $w_{j}\|P(x_{j})\|_{2}^{2}$
-13:         $P_{wx}\leftarrow P_{wx}+w_{j}\|P(x_{j})\|_{2}^{2}$
-14:         $s_{w}\leftarrow s_{w}+w_{j}$
-15:         $P_{j,:}\leftarrow P(x_{j})$
-16:     end for
-17:     for $i=1\to r$ do
-18:         $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{r\log(1/\delta^{\prime})}$, $\delta/r$) $\triangleright$ Alg. 6
-19:     end for
-20:end procedure
-21:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma 6.7
-22:     Value $\leftarrow 0$
-23:     Compute $P(y)$
-24:     Compute $\|P(y)\|_{2}^{2}$
-25:     for $i=1\to r$ do
-26:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(P(y)_{i},\alpha)$triangleright Alg. 7
-27:     end for
-28:     Value $\leftarrow$ $0.5\cdot(P_{wx}+s_{w}\|P(y)\|_{2}^{2}~{}-\text{Value})$
-29:     return Value
-30:end procedure
-31:end datastrucutre
 Algorithm 1 Adaptive query data structure
@@
 2:members
-3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$triangleright Alg. 3
+3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$triangleright Algorithm 3
 4:end members
@@
 19:end datastructure
+3 Main Results: Cross-Attention
+In this section, we show our main result for cross-attention.
+Theorem 3.1 states that we can ensure the entire cross-attention module satisfies DP and is robust to adaptive queries.
+Our high-level idea is based on the similarity between weighted distance problem and cross-attention.
+For a typical weighted distance problem, we define the following: Let $w\in\mathbb{R}^{n}$ be the weights, $X\in\mathbb{R}^{n\times d}$ be the data matrix, where $x_{i}^{\top}$ is the $i$-th row of $X$ for $i\in[n]$, and let $y\in\mathbb{R}^{d}$ be the query.
+Suppose we need to answer $\ell_{1}$-distance query.
+We have
+$\sum_{i\in[n]}\underbrace{w_{i}}_{\mathrm{weight}}\|\underbrace{y}_{\mathrm{query}}-\underbrace{x_{i}}_{\mathrm{data}}\|_{1}.$
+[…]
+$(AV)_{i,k}=$ $~{}\sum_{j=1}^{n}\underbrace{V_{j,k}}_{\mathrm{weight}}\exp(\langle\underbrace{Q_{i}}_{\mathrm{query}},\underbrace{K_{j}}_{\mathrm{data}}\rangle/d),$ (1)
+which can be viewed as a weighted Softmax problem, where $V$ provides the weights, $Q$ is the query, and $K$ is the dataset.
+Thus, we choose to add noise to $K$ and $V$ based on the similarity between the weighted distance problem and cross-attention.
+Furthermore, we find that we can only handle one column of $V$, i.e., $V_{*,k}\in\mathbb{R}^{n}$, in a single data structure.
+Therefore, we need to initialize a total of $d$ different data structures, each with weights $V_{*,k}$ for $k\in[d]$.
+Here, we present our main result below.
+Theorem 3.1 (Softmax cross-attention, informal version of Theorem I.12).
+Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1.
+Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter.
+Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma C.7).
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
+Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$.
+There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnrd)$ spaces to ensure cross-attention satisfies DP and supports the following operations:
+  • We initialize $d$ data structures using Init$(K,n,V_{*,k},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$ (Algorithm 1), for $k\in[d]$.
+It takes $O(lnr)$ time to initialize one data structure.
+  • At query time, for user input $Q$, we process one token at a time by passing the $i$-th row of $Q$, denoted $Q_{i}\in\mathbb{R}^{d}$, to DistanceQuery$(Q_{i},\alpha\in(0,1))$ (Algorithm 1) for each $i\in[m]$.
+It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
+    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP,
+    • the process of output $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
+    • the process of output $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
+    • it is robust to adaptive query.
+In Theorem 3.1, we use our DPTreeSoftmaxAdaptive (Algorithm 1) and guarantee that, for each query token of cross-attention, the output process satisfies $(\epsilon,\delta)$-DP with $n^{-1}(\alpha+\epsilon_{s})$ relative error and $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query.
+More specifically, the algorithm creates $d$ DPTreeSoftmaxAdaptive data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time.
+Notably, our error is inversely proportional to $n$, meaning that as the input token length increases, both the relative and approximate errors approach zero.
+This is achieved by the normalizing matrix $D$ (Definition 1.1).
+We refer the reader to Section I for proof details.
+Thus, our algorithm theoretically protects system prompts/RAG data in cross-attention as discussed in Section 1.
+In Section 4, we provide a detailed technical overview, and in Section 5, we will present self-attention and DP-related discussion.
+Algorithm 2 DPTree initialization and query
+1:datastructure DPTree $\triangleright$ Theorem C.1
+2:members
+3:     $b:\mathbb{R}^{2n-1}$
+4:     $c:\mathbb{R}^{2n-1}$
+5:end members
+6:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{R},\epsilon\in(0,1),\delta\in(0,1)$) $\triangleright$ Lemma D.3, Lemma C.3
+7:     $b[n,2n-1]\leftarrow a$
+8:     for $i=n\to 2n-1$ do
+9:         $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
+10:     end for
+11:     for $i=(\log n)\to 1$ do
+12:         for $j=1\to 2^{i-1}$ do
+13:              $k\leftarrow 2^{i-1}+j-1$
+14:              $b[k]\leftarrow b[2k]+b[2k+1]$
+15:              $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
+16:         end for
+17:     end for
+18:end procedure
+19:procedure Query($x\in[n],y\in[n]$) $\triangleright$ Lemma D.4, D.5, D.6
+20:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $c$) by using at most $2\log n$ nodes on the path.
+Let Value be the above summation.
+21:     return Value
+22:end procedure
+23:procedure TrueQuery($x\in[n],y\in[n]$)
+24:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $b$) by using at most $2\log n$ nodes on the path, where the height of the tree is $\log n$, and we need left and right boundary points.
+Let Value be the above summation.
+25:     return Value
+26:end procedure
+27:end datastructure
+4 Key Data Structure: DPTree
+This section provides our key data structures: DPTree (Algorithm 2), DPTreeDistance (Algorithm 5 and 6), DPTreeHighDim (Algorithm 7), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 1).
+In Section 4.1, we provide our high-level proof insights.
+In Section 4.2, we give our basic building block algorithms DPTree, DPTreeDistance and DPTreeHighDim.
+In Section 4.3, we present our DPTreeSoftmax algorithm that solves the weighted Softmax problem.
+In Section 4.4, we present our DPTreeSoftmaxAdaptive algorithm that enables DPTreeSoftmax to handle adaptive query problem.
+4.1 Technique Overview
+Notice that Eq.
+(1) is not a typical distance measure like $\ell_{1}$ or $\ell_{2}$, but by using polynomial kernel method techniques, we transform it into a distance measure.
+AS [6] states that the exponential inner product can be approximated by polynomial kernel function $P(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{r}$, i.e., $P(x)^{\top}P(y)\approx\exp(x^{\top}y/d)$ for two vector $x,y\in\mathbb{R}^{d}$, with a relative error.
+Then, by the Law of Cosines, we transform the inner product of polynomial kernel functions into a distance measure, i.e.,
+$2P(x)^{\top}P(y)=$ $~{}-\|P(x)-P(y)\|_{2}^{2}+\|P(x)\|^{2}_{2}+\|P(y)\|^{2}_{2}.$ (2)
+After transforming Eq. (1) into a distance measure, we design the DPTree series data structures to provide cross-attention DP guarantee.
+In summary, we first design the data structure DPTree (Algorithm 2) that builds a binary segment tree with truncated Laplace noise added in the leaf nodes to ensure DP guarantee.
+Then, based on this data structure, we design DPTreeDistance (Algorithm 5 and 6) to answer one dimensional weighted distance queries $\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|$, which utilizes DPTree to store and return noised weights $w_{i}$ multiplied with the approximated distances between the query $y$ and data $x_{i}$.
+We further decompose high dimensional $\ell_{p}^{p}$-distance problem into one dimensional $\ell_{1}$-distance problems using
+$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=$ $~{}\sum_{k=1}^{d}\sum_{i=1}^{n}w_{i}\cdot|y_{k}-x_{i,k}|^{p}.$ (3)
+Based on this decomposition, we design DPTreeHighDim (Algorithm 7) which is capable of answering high dimension queries.
+Then, using Eq.
+(2) and DPTreeHighDim, we design DPTreeSoftmax (Algorithm 3) to answer Softmax queries.
+By building multiple copies of this data structure, we boost the success probability such that it can answer any query (including adaptive query) with an additive error, establishing the final data structure DPTreeSoftmaxAdaptive (Algorithm 1).
+See Section C for a more detailed outline of algorithms and proof techniques.
+4.2 DPTree, DPTreeDistance, and DPTreeHighDim
+We design a basic data structure DPTree (Algorithm 2) that answers summation queries by a summation segment tree with truncated Laplace noise (Definition 2.4).
+The algorithm first builds a binary summation tree in an array and then adds truncated Laplace noises to each node.
+In query time, we first trace from bottom nodes to find their lowest common ancestor, then report the summation by using at most $2\log n$ nodes on the path (Algorithm 2).
+Based on the parallel composition rule of DP (Fact B.7), we find that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries (Lemma D.8).
+See more details in Section D.
+We then design DPTreeDistance, a one-dimensional weighted $\ell_{1}$ distance data structure detailed in Algorithm 5 and 6.
+Initialization involves rounding each data point to the nearest multiple of a small interval and aggregating their weights into an array (illustrated in Figure 1), which is then input into our DPTree.
+At query time, we retrieve aggregated weights within small intervals and multiply them by their distances to the query point.
+We introduce a relative error parameter $\alpha$ to reduce the number of iterations to $O(\log(n)/\alpha)$, improving efficiency.
+Guided by Eq.(3), we design DPTreeHighDim (Algorithm 7), which extends DPTreeDistance to higher dimension by constructing independent data structures for each coordinate.
+See details in Section F and G.
+Algorithm 3 Softmax query
+1:datastrucutre DPTreeSoftmax $\triangleright$ Theorem 4.2
+2:members
+3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$triangleright Algorithm 5, Theorem I.7
+4:     $P:[0,\Gamma_{R,s}]^{n\times r}$triangleright Definition I.3 for $\Gamma_{R,s}$, Eq. (9) for $s$, Eq. (10) for $r$
+5:     $w:[-R_{w},R_{w}]^{n}$
+6:     $P_{wx},~{}s_{w},~{}\epsilon_{s}:\mathbb{R}$
+7:end members
+8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma C.7
+9:     $\epsilon_{s},~{}w,~{}P,~{}P_{wx},~{}s_{w}\leftarrow\epsilon_{s},~{}w,~{}0^{n\times r},~{}0,~{}0$
+10:     for $j=1\to n$ do
+11:         Compute $P(x_{j})$triangleright Polynomial kernel function $P(\cdot)$, Lemma I.5
+12:         Compute $w_{j}\|P(x_{j})\|_{2}^{2}$
+13:         $P_{wx}\leftarrow P_{wx}+w_{j}\|P(x_{j})\|_{2}^{2}$
+14:         $s_{w}\leftarrow s_{w}+w_{j}$
+15:         $P_{j,:}\leftarrow P(x_{j})$
+16:     end for
+17:     for $i=1\to r$ do
+18:         $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{r\log(1/\delta^{\prime})}$, $\delta/r$) $\triangleright$ Algorithm 5
+19:     end for
+20:end procedure
+21:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma C.7
+22:     Value $\leftarrow 0$
+23:     Compute $P(y)$
+24:     Compute $\|P(y)\|_{2}^{2}$
+25:     for $i=1\to r$ do
+26:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(P(y)_{i},\alpha)$triangleright Algorithm 6
+27:     end for
+28:     Value $\leftarrow$ $0.5\cdot(P_{wx}+s_{w}\|P(y)\|_{2}^{2}~{}-\text{Value})$
+29:     return Value
+30:end procedure
+31:end datastrucutre
 4.3 Softmax Activation
-In this section, we present the algorithm (Algorithm 3) that can answer the weighted softmax query (Definition 4.1) and can be further used to design DP cross-attention.
-First, we introduce the definition of weighted softmax query, which is also an abstraction for the problem described in Section 3.4.
-Definition 4.1 (Softmax query).
-Let $\phi$ be the distance measurement function defined in Definition 1.1.
+In this section, we present DPTreeSoftmax (Algorithm 3) that answers the weighted Softmax query (Definition 4.1) and is further used to design DP cross-attention.
+First, we introduce the definition of weighted Softmax query, an abstraction for the problem described in Eq.
+(1).
+Definition 4.1 (Weighted Softmax query (without normalization)).
 For the dataset $X\in[0,R]^{n\times d}$ where $x_{i}^{\top}$ is the $i$-th row of $X$ and query $y\in[0,R]^{d}$, we define the weighted exponential inner product/Softmax query to be:
-$\sum_{i\in[n]}w_{i}\phi(x_{i},y)=\sum_{i\in[n]}w_{i}\exp(\langle x_{i},y\rangle/d)=w^{\top}\exp(Xy/d).$
-Based on the Definition 4.1, we design a novel algorithm that can make weighted softmax query DP. More details can be found in Section G.
-Theorem 4.2 (Softmax query, informal version of Theorem G.8).
+$\sum_{i\in[n]}w_{i}\exp(\langle x_{i},y\rangle/d)=w^{\top}\exp(Xy/d).$
+Building on Definition 4.1, we develop a novel algorithm to answer differentially private weighted Softmax queries using the polynomial kernel method from AS [6].
+Specifically, in Eq.(2), there is a term that computes the weighted $\ell_{2}^{2}$ distance, which we calculate using DPTreeHighDim.
+We then compute the exact term for the weighted $\ell_{2}^{2}$ norms of the approximation kernel.
+By summing these terms with a controlled error, we extend DPTreeHighDim to answer the Softmax query efficiently.
+More details can be found in Section I.
+Theorem 4.2 (Softmax query, informal version of Theorem I.8).
 Let $R\geq 1$.
 Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
 Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
-There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
+Our data structure DPTreeSoftmax (Algorithm 3) uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support following operations:
   • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1))$.
@@
   • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 3) It takes $O(\alpha^{-1}r\log^{2}n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
-    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
-    • it holds with probability $0.99$.
+    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
+    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
+      $+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • it holds with probability at least $0.99$.
 Remark 4.3.
-In Theorem 4.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section 6.5.
+In Theorem 4.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section C.5.
 Besides, note that the error bound in Theorem 4.2 does not depend on $\delta$ but depends on $\delta^{\prime}$.
-The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ can reduce the error by a small constant (Fact 3.7).
+The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ reduces the error by a small constant (Fact 2.5).
 In practice, we set $\delta$ as a small positive constant close to 00.
-Please refer to the Lemma B.6 for more details.
+Please refer to the Lemma D.6 for more details.
 4.4 Adaptive Query Data Structure
-We can now present the theorem for our data structure that can answer weighted softmax queries with a DP guarantee and concrete error bound, and it is robust to adaptive queries.
-The corresponding proofs can be found in Section F and Section G.
-Theorem 4.4 (Adaptive query softmax data structure, informal version of Theorem G.11).
+We adapt our DPTreeSoftmax to DPTreeSoftmaxAdaptive (Algorithm 1) to solve the adaptive query problem.
+By proving it can handle any query within the query space with a certain error, we ensure it effectively processes adaptive queries.
+We first boost the constant probability to high probability using the Chernoff bound (Lemma B.2).
+Employing an $\epsilon_{0}$-net argument and the union bound, we bound all query points within the net.
+Finally, we use the Lipschitz property of the weighted Softmax distance function with an additive error to bound all points in the query space.
+The corresponding proofs can be found in Section H and Section I.
+Theorem 4.4 (Adaptive query Softmax data structure, informal version of Theorem I.11).
 Let $R\geq 1$.
 Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
 Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
@@
   • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$.
-(Algorithm 4) It takes $O(lnr)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 4) It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
-    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
-    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
+It takes $O(lnr)$ time to initialize the data structure.
+  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output a number $z$ such that
+    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
+    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
+      $+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • it holds with probability at least $1-p_{f}$ (where $p_{f}$ is used in $l$),
     • it is robust to adaptive query.
@@
 In the second for loop of Init and the for loop of DistanceQuery in Algorithm 3, the $r$ DPTreeDistance data structures instantiated for each coordinate are independent of each other.
-In addition, the for loops in Algorithm 4 can also be parallelized since the $l=O(r\log(dR/(\epsilon_{s}p_{f})))$ copies are independent.
+In addition, the for loops in Algorithm 1 are also parallelizable since the $l=O(r\log(dR/(\epsilon_{s}p_{f})))$ copies are independent.
 After parallelization, we have the final time complexity of Init to be $O(nr)$ and DistanceQuery to be $O(\alpha^{-1}\log^{2}n)$ in Algorithm 1 with $O(lr)$ GPU process.
-In Theorem 4.4, we show that our DPTreeSoftmaxAdaptive data structure defined above have $(\epsilon,\delta)$-DP guarantee with $\alpha$ relative error and $O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}(n)+w^{\top}\exp(Xy/d))$ additive error under $O(lnr)$ memory consumption, $O(lnr)$ initialization time and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query.
-We can apply our data structure in cross-attention, and we show detail in the following section.
-5 Main Results: Cross-Attention
-In this section, we show our main result for cross-attention. Theorem 5.1 states that we can make the cross-attention module DP.
-Theorem 5.1 (Cross-attention, informal version of Theorem G.12).
-Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1.
-Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter.
-Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma 6.7).
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
-Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$.
-Let $\phi$ be softmax function (Definition 4.1).
-There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
-  • We initialize $d$ data structures using Init$(K,n,V_{*,k},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$ (Algorithm 4), for $k\in[d]$.
-It takes $O(lnr)$ time to initialize one data structure.
-  • At query time, for user input $Q$, we process one token at a time by passing the $i$-th row of $Q$, denoted $Q_{i}\in\mathbb{R}^{d}$, to DistanceQuery$(Q_{i},\alpha\in(0,1))$ (Algorithm 4) for each $i\in[m]$.
-It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
-    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP,
-    • $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
-    • $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
-    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
-    • it is robust to adaptive query.
-In Theorem 5.1, we use our DPTreeSoftmaxAdaptive (see Section 6.6) and guarantee that for each query token of cross-attention, the output is $(\epsilon,\delta)$-DP with $n^{-1}(\alpha+\epsilon_{s})$ relative error and $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query.
-More specifically, the algorithm creates $d$ DPTreeSoftmaxAdaptive data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time.
-Thus, our algorithm can theoretically protect system prompts/RAG data in cross-attention as discussed in Section 1.
-In Section 6, we provide a detailed technical overview, and in Section 7, we will present self-attention and DP-related discussion.
-6 Technique Overview
-This section provides the technique overview of our paper.
-In Section 6.1, we analyze our DPTree data structure.
-In Section 6.2, we show the sensitivity of summation problem.
-In Section 6.3, we explain the high-level idea behind the weighted $\ell_{p}^{p}$ distance query.
-In Section 6.4, we show how to answer one-dimensional weighted $\ell_{1}$ distance query.
-In Section 6.5, we show how to answer softmax distance query using previous algorithms.
-In Section 6.6, we show how to handle adaptive query.
-By combining the results from these sections, we prove the main results in Section 4.
-6.1 Summation Segment Tree
-First, in order to solve the weighted distance problem, we need to have a basic DP algorithm (Algorithm 1 and 2) that can answer simple summation queries.
-After analyzing its DP and error in Section B, we state the data structure theorem.
-Theorem 6.1 (DPTree data structure, informal version of Theorem B.1).
-There is a data structure (see DPTree in Algorithm 1, 2) that uses $O(n)$ spaces to support the following operations:
-  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
-  • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
-    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
-    • it holds with probability $0.99$.
-During the design of the data structure, we found an interesting property based on the parallel composition rule of DP Fact 3.11. We will now state the lemma with its proof provided in Section B.
-Lemma 6.2 (Weighted sum of disjoint interval queries, informal version of Lemma B.8).
-If the following conditions hold that:
-  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
-  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
-  • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
-Then, we have $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ (see Alg. 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
-$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
-with probability $0.99$.
-From Lemma 6.2, we can see that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries, as long as the sum of squared weights is finite.
-6.2 Sensitivity for Range Summation Problem
-Our DP summation tree data structure DPTree (Algorithm 1 and 2) requires sensitivity parameter $\Delta$.
-In this section, we show that for the summation problem, we have the sensitivity $\Delta=2R$ if the input $X\in[-R,R]^{n}$.
-Lemma 6.3 (Sensitivity of summation).
-Let $X\in[-R,R]^{n}$. We have the sensitivity $\Delta=2R$ for DPTree.Init in Algorithm 1.
-Proof.
-Let’s say 2 neighboring dataset $X$ and $X^{\prime}$ differ in $x_{i}$ and $x^{\prime}_{i}$ for some $i$ in the array $X$. Then for a summation problem, i.e. $f(X):=\sum_{i=1}^{n}x_{i}$, we have
-$\Delta=$ $~{}\max_{X,X^{\prime}}|f(X)-f(X^{\prime})|=\max_{X,X^{\prime}}|x_{i}-x^{\prime}_{i}|=2R.$
-where the first step follows from Definition 3.2, the second step follows from $X,X^{\prime}$ differ in $x_{i},x^{\prime}_{i}$, and the last step follows from each coordinate of the dataset is bounded in $[-R,R]$.
-∎
-6.3 Weighted $\ell_{p}^{p}$ Distance Problem
-Figure 1: The visualization of how to build $c_{j}$ of rounded dataset $X\in[0,1]^{10}$ and compute the weighted $\ell_{1}$ distance.
-The number above each $x_{i}$ is $w_{i}$.
-See Algorithm 6 for details.
-Suppose $y=0$.
-Then $\sum_{i=1}^{n}w_{i}|y-x_{i}|=0.1\cdot 2.2+0.3\cdot 3.1+0.3\cdot(-2)+0.3\cdot(-3)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5+0.9\cdot(-1)+0.9\cdot 1=4.4$.
-And $\sum_{j=0}^{n}|k-j|c_{j}/n=0.1\cdot 2.2+0.3\cdot(-1.9)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5=4.4$.
-See details in Lemma C.1.
-In this section, we introduce the intuition behind the method for handling the weighted $\ell_{p}^{p}$ distance problem. The formal lemmas and proofs can be found in Section C.
-Given a dataset and a query point in $d$ dimensions, we round each coordinate of the data points and the query point to the nearest multiple of a small interval.
-We then aggregate the weights of data points that have been rounded to the same position.
-Finally, we compute the sum of these aggregated weights multiplied by the distances between the query point and the data points over the rounded positions.
-This approach makes the computation more efficient while maintaining sufficient accuracy.
-We provide an example of weighted $\ell_{1}$-distance of a one-dimensional dataset consisting of 10 data points, i.e., $X\in[0,1]^{10}$ and a query $y=0$ in Figure 1.
-Lemma 6.4 (Weighted $\ell_{p}^{p}$-distance high dimension, informal version of Lemma C.2).
-If the following conditions hold:
-  • Let data $X\in[0,R]^{n\times d}$ and $x_{i}^{\top}\in[0,R]^{d}$ be the $i$-th row of $x$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]^{d}$.
-  • We round each dimension of $X$ and $y$ to an integer multiple of $R/n$.
-  • Let $x_{i,k},y_{k}$ denote the $k$-th coordinates of $x_{i},y$ for $k\in[d]$.
-  • Let $c_{j,k}:=\sum_{j_{0}\in S_{j,k}}w_{j_{0}}$ where the set $S_{j,k}$ is the set of index $i$ such that the corresponding $x_{i,k}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
-  • After rounding, we assume that $y_{k}$ is in the $l_{k}R/n$ position for $l_{k}\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
-For the weighted problem, we have
-$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}.$
-where $O(R/n)$ is the rounding error for each data point.
-Remark 6.5.
-In Lemma 6.4, we first round the dataset.
-This rounding simplifies the calculation by reducing the number of possible positions to consider, from real values in $[0,R]^{d}$ to the total $O(nd)$ spots.
-However, it also introduces an error $O(R/n)$ for one data point.
-Then, for one spot in the rounded dataset, we sum over the weights of that spot and multiply the corresponding distance raised to the power of $p$.
-Additionally, since we are dealing with $\ell_{p}^{p}$ distance, the rounding error is also raised to the power of $p$.
-6.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
-Based on previous discussions in Section 6.1 and 6.3, we can now describe our one-dimensional weighted $\ell_{1}$ distance data structure, DPTreeDistance, presented in Algorithm 6 and 7, which generalizes the results from [8].
-Drawing from the intuition in Section 6.3, the initialization process is as follows: first, we round each data point in the dataset to the nearest multiple of a small interval and build an array that aggregates the corresponding weights.
-This array is then fed into our DPTree data structure in Algorithm 1.
-At query time, we query the DPTree to obtain the aggregated weights within a small interval and multiply these weights by the distance to the query point.
-Furthermore, we also introduce a relative error parameter $\alpha$ to reduce the total number of queries to $O(\log(n)/\alpha)$ instead of querying all $n$ positions.
-We also analyze the DP and the error bound; see details in Section D.
-Theorem 6.6 (DPTreeDistance data structure, informal version of Theorem D.6).
-There is a data structure DPTreeDistance (Algorithm 6,7) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
-  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 6) It takes $O(n)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 7) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
-    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$,
-    • it holds with probability $0.99$.
-6.5 Softmax Activation
-We then describe how we extend the previous results to softmax activation, i.e.
-exponential inner product function (Definition 4.1).
-From [6], we know that softmax activation can be approximated by polynomial kernel function $P(\cdot)$ with a certain error.
-The following lemma shows that we can transform weighted softmax queries into polynomial kernels.
-More specifically, we have one term that computes the weighted $\ell_{2}^{2}$ distance, which is the place where we add DP noises.
-Because of the decomposability of the $\ell_{p}^{p}$ distance, i.e.
-$\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p}$, we can easily extend the results of Section 6.4 to handle the $\ell_{2}^{2}$ distance query.
-After that, we compute the term for the weighted $\ell_{2}^{2}$ norms of approximation kernel exactly.
-Summing all these terms, with a certain error, we can answer the softmax query.
-Related proofs can be found in Section G.
-Lemma 6.7 (Weighted softmax approximation, informal version of Lemma G.6).
-Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
-Let $R\geq 1$.
-Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
-Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma G.5.
-Then, we can approximate the exponential inner product using the polynomial kernel function:
-$w^{\top}\exp(Xy/d)=$ $~{}-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
-$+~{}O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
-Moreover, the vectors $P(\cdot)$ can be computed in $O(r)$ time.
-6.6 Adaptive Query
-We introduce how we can modify our algorithm to solve the adaptive query problem using some tools in [75].
-Our approach is based on proving that our algorithm can handle any query within the query space with a certain error.
-Since adaptive queries must lie within this space, our algorithm can effectively handle them.
-In Section 6.5, we demonstrate our algorithm’s capability to answer weighted softmax distance queries with constant probability.
-We then use the Chernoff bound to boost the constant probability of our algorithm to a high probability.
-Next, we apply the notion of an $\epsilon_{0}$-net to bound all query points within the net using the union bound.
-Finally, we bound all points in the query space by utilizing the Lipschitz property of the weighted softmax distance function and introducing an additive error.
-See the proofs in Sections F and G.
-Lemma 6.8 (Adaptive softmax, informal version of Lemma G.10).
-If the following conditions hold:
-  • Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and let $|N|$ be the size of the net $N$.
-  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
-  • Let the relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
-  • We create $l=O(\log((R/\epsilon_{0})^{r}/p_{f}))$ independent copies of the data structure $\{\textsc{DPTreeSoftmax}_{j}\}_{j=1}^{l}$ (Algorithm 3) and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
-  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
-  • Let $Z(y):=w^{\top}\exp(Xy/d)$.
-  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
-Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
-$|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
 5 Discussion
@@
 How we can solve this challenge is crucial, and we leave it as our future direction.
-Why not add noise in some other places?
+Why not add noise to some other places?
 Where and how to add DP noises is an important problem to ask during the DP algorithm design.
 In this paper, we consider the problem of $\sum_{i=1}^{n}w_{i}\exp(\langle x_{i},y\rangle/d)$ where $y,x_{i}\in[0,R]^{d}$ and $w\in[-R_{w},R_{w}]^{n}$ (Definition 4.1).
-Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorihtm 1 and 2).
-From Lemma 6.3 and the way we initialize DPTree in Algorithm 6, we can see that the sensitivity $\Delta$ of this problem is $2R_{w}$.
+Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorihtm 2).
+From Lemma C.3 and the way we initialize DPTree in Algorithm 5, we see that the sensitivity $\Delta$ of this problem is $2R_{w}$.
 A simple method for adding noise involves adding $n$ noises to a length $n$ array, with each item $w_{i}\exp(\langle x_{i},y\rangle/d)$ for $i\in[n]$.
-However, this approach increases the error by a factor of $n$ by basic composition (Fact 3.10) and also makes the model dependent on the number of queries.
+However, this approach increases the error by a factor of $n$ by basic composition (Fact B.6) and also makes the model dependent on the number of queries.
 Besides, it only supports a single query and requires rebuilding the tree for each new query, rendering it impractical.
-In contrast, our current noise-adding technique (Lines 11 and 17 of Algorithm 1) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$.
+In contrast, our current noise-adding technique (Lines 9 and 15 of Algorithm 2) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$.
 This method also supports multiple queries, eliminating the need to rebuild the tree each time.
+How to remove the relative error parameter $\alpha$?
+The relative error parameter $\alpha$ in Theorem 3.1 appears because of the $(1+\alpha)$-approximation introduced in Algorithm 5 (Remark F.3) to reduce the number of required iterations from naive $O(n)$ to $O(\log(n)/\alpha)$.
+However, we notice that a recent work [59] does not utilize $(1+\alpha)$-approximation and still achieves $O(\log n)$ iteration number.
+They introduce a new tree node representation where each node stores the sum of distances from one point to multiple points, enabling the answer to be divided into only $\log n$ values, each combining two distance values, two count values, and $y$ itself.
+Our DPTree algorithms can be integrated with their method, thus removing parameter $\alpha$.
 6 Conclusion
@@
 The appendix is organized as follows.
-In Section A, we give the preliminary of our paper.
-In Section B, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee.
-In Section C, we show how to solve weighted distance problem.
-In Section D, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{1}$ distance problem with DP and accuracy guarantee.
-In Section E, we present the analysis of our DPTreeHighDim (Algorithm 8) data structure, which can address the high-dimensional $\ell_{1}$ distance problem while ensuring differential privacy and accuracy guarantees.
-In Section F, we show how we can handle adaptive query.
-In Section G, we show how to extend our algorithm to softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 4).
-Appendix A More Preliminary
-In Section A.1, we give the probability tools we use in the paper. In Section A.2, we provide the facts we use.
-A.1 Probability Tools
+In Section A, we provide more related works.
+In Section B, we give the preliminary of our paper.
+In Section C, we offer an outline of our proof techniques.
+In Section D, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee.
+In Section E, we show how to solve weighted distance problem.
+In Section F, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{1}$ distance problem with DP and accuracy guarantee.
+In Section G, we present the analysis of our DPTreeHighDim (Algorithm 7) data structure, which can address the high-dimensional $\ell_{1}$ distance problem while ensuring differential privacy and accuracy guarantees.
+In Section H, we show how we can handle adaptive query.
+In Section I, we show how to extend our algorithm to Softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 1).
+Appendix A More Related Work
+Differential Privacy Guarantee Analysis.
+Ever since DMNS [19] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [22, 67, 115, 84, 111].
+DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [50, 24, 3, 61, 51, 37, 10, 12, 25, 16, 46, 78, 79, 56, 62, 28, 29, 60, 17, 13, 18, 27, 40, 39, 64, 38, 23, 95, 21, 107, 94, 42, 70, 66].
+Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [20].
+For example, truncated Laplace mechanism [32] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions.
+Cross-Attention in System Prompt, RAG, Stable Diffusion and More.
+Cross-attention [97], first introduced in language translation, is a widely used technique in many advanced AI systems.
+For example, Stable Diffusion [87, 73, 98, 103, 106] and SORA [81] employ cross-attention as a core module for a text-to-image conditional generation.
+This technique is also utilized by other multimodal models [72], including Imagen [90] and Diffusion Transformer [85].
+In the realm of text-to-image editing, HMT+ [49] analyzes and controls the cross-attention module to enable editing without requiring additional training.
+Furthermore, YYB+ [114] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas.
+In addition, Retrieval Augmented Generation (RAG) [68, 11, 43], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module.
+Cross-attention also has other applications.
+ORST [82] demonstrates that the prompt-tuning [71] task can be formulated as cross-attention, while CFP [14] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation.
+Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [104], such as spatial reasoning [101], mathematical reasoning [65], in-context learning ability [93], compositional ability [109], few-shot adaptation ability [89, 110], and so on.
+There are some other works that used cross attention in Hopfield Models [52, 100, 48, 108, 99, 44, 45].
+Appendix B More Preliminary
+In Section B.1, we give the probability tools we use in the paper.
+In Section B.2, we provide the algebraic facts we use.
+In Section B.3, we give the DP facts we use in the paper.
+In Section B.4, we compare between popular DP mechanisms.
+B.1 Probability Tools
 In this section, we give several probability lemmas.
-Lemma A.1 (Markov’s inequality).
+Lemma B.1 (Markov’s inequality).
 If $x$ is a nonnegative random variable and $t>0$, we have
 $\Pr[x\geq t]\leq\frac{\operatorname*{{\mathbb{E}}}[x]}{t}.$
-Lemma A.2 (Chernoff bound, [13]).
+Lemma B.2 (Chernoff bound, [15]).
 Let $x_{i}$ be a Bernoulli random variable with probability $p_{i}$ of being equal to $1$ and $1-p_{i}$ of being equal to 00, and all $x_{i}$ for $i\in[n]$ are independent.
@@
 $\Pr[x\leq(1-\delta)\mu]\leq\exp(-\delta^{2}\mu/2).$
-Lemma A.3 (Chebyshev’s inequality).
+Lemma B.3 (Chebyshev’s inequality).
 Let $x$ (integrable) be a random variable with finite non-zero variance $\sigma^{2}$ (and thus finite expected value $\mu$). Then for any real number $k>0$,
 $\Pr[|x-\mu|\geq k\sigma]\leq\frac{1}{k^{2}}.$
-A.2 Facts
-Fact A.4 (Upper bound of exponential, Fact C.9 in [41]).
+B.2 Algebraic Facts
+Fact B.4 (Upper bound of exponential, Fact C.9 in LSSZ24b [73]).
 For $a\in\mathbb{R}$, $b\in\mathbb{R}$, $a,b\leq R$, where $R\geq 0$, we have
 $|\exp(a)-\exp(b)|\leq\exp(R)|a-b|.$
-Appendix B DPTree Algorithm
-In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 1 and 2).
-In Section B.1, we give the theorem (Theorem B.1) of our data structure that can answer summation problem.
-In Section B.2, we improve our data structure from constant probability to high probability by applying Chernoff bound.
-In Section B.3, we give the analysis.
-In Section B.4, we show some results of our data structure if the input queries are disjoint.
-B.1 Single Data Structure
-We give the theorem of our DPTree data structure that can answer the summation problem with DP, accuracy, runtime guarantee.
-Theorem B.1 (DPTree data structure, formal version of Theorem 6.1).
-There is a data structure (see DPTree in Algorithm 1, 2) that uses $O(n)$ spaces to support the following operations:
+B.3 DP Facts
+In this section, we present several facts about differential privacy (DP). We first state the post-processing property, which means, in an algorithm, if one step is DP, all the following steps are DP.
+Fact B.5 (Post-processing, see Fact 2.1 in GKK+ [37]).
+Let ${\mathcal{A}}_{1}$ be an $(\epsilon,\delta)$-DP algorithm and ${\mathcal{A}}_{2}$ be a (randomized) post-processing algorithm.
+Then the algorithm ${\mathcal{A}}(X)={\mathcal{A}}_{2}({\mathcal{A}}_{1}(X))$ is still an $(\epsilon,\delta)$-DP algorithm.
+If we have many DP algorithms, we need a composition rule. The most straightforward composition is the basic/sequential composition rule.
+Fact B.6 (Basic composition, see Fact 2.3 in GKK+ [37]).
+Let ${\mathcal{A}}_{1}$ be an $(\epsilon_{1},\delta_{1})$-DP algorithm and ${\mathcal{A}}_{2}$ be an $(\epsilon_{2},\delta_{2})$-DP algorithm.
+Then ${\mathcal{A}}(X)=({\mathcal{A}}_{1}(X),{\mathcal{A}}_{2}({\mathcal{A}}_{1}(X),X))$ is an $(\epsilon_{1}+\epsilon_{2},\delta_{1}+\delta_{2})$-DP algorithm.
+We can do much better if we know that the inputs are disjoint.
+Fact B.7 (Parallel composition, see Fact 2.4 in GKK+ [37]).
+Let ${\mathcal{A}}_{1}$ be an $(\epsilon_{1},\delta_{1})$-DP algorithm and ${\mathcal{A}}_{2}$ be an $(\epsilon_{2},\delta_{2})$-DP algorithm.
+Assume ${\mathcal{A}}_{1}$ and ${\mathcal{A}}_{2}$ depend on disjoint subsets of input coordinates.
+Then the algorithm ${\mathcal{A}}(X)=({\mathcal{A}}_{1}(X),{\mathcal{A}}_{2}({\mathcal{A}}_{1}(X),X))$ is a $(\max\{\epsilon_{1},\epsilon_{2}\},$ $\max\{\delta_{1},\delta_{2}\})$-DP algorithm.
+In addition, we have the advanced composition, which improves the dependence of the number of DP algorithms to square root but compromises the term $\delta^{\prime}$.
+Theorem B.8 (Advanced composition, see Theorem 3.20 in DR [20]).
+For all $\epsilon,\delta,\delta^{\prime}\geq 0$, the class of $(\epsilon,\delta)$-differentially private mechanisms satisfies $(\epsilon^{\prime},k\delta+\delta^{\prime})$-differential privacy under $k$-fold adaptive composition for:
+$\epsilon^{\prime}=k\epsilon(e^{\epsilon}-1)+\epsilon\sqrt{2k\log(1/\delta^{\prime})}.$
+B.4 Comparison of Truncated Laplace, Gaussian, and Laplace Mechanisms
+We first define the Laplace mechanism as below:
+Definition B.9 (Laplace distribution).
+We use $\mathrm{Lap}(b)$ to denote the pdf: $p(z)=\frac{1}{2b}\exp(-\frac{|z|}{b}).$
+Fact B.10.
+For $z\sim\mathrm{Lap}(b)$, $\operatorname*{{\mathbb{E}}}[z]=0$, and $\operatorname*{\mathrm{Var}}[z]=2b^{2}$.
+Furthermore, if $b=\Delta/\epsilon$, we have $\operatorname*{\mathrm{Var}}[z]=2\Delta^{2}/\epsilon^{2}$.
+In this paper, we use the Chebyshev inequality to bound the error, and from GDGK [32], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
+The variance of Gaussian mechanism in Theorem 3.22 in DR [20]:
+$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}\log(1.25/\delta)}{\epsilon^{2}}.$
+The variance of Laplace mechanism in Fact B.10:
+$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}}{\epsilon^{2}}.$
+The variance of truncated Laplace mechanism in Fact 2.5, for $c\in(0,1]$:
+$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}c}{\epsilon^{2}}.$
+Thus, since it has the minimum variance, we choose the truncated Laplace mechanism to design our algorithms among these popular mechanisms.
+Appendix C Proof Outline
+This section provides the proof outline of our paper.
+In Section C.1, we analyze our DPTree data structure.
+In Section C.2, we show the sensitivity of summation problem.
+In Section C.3, we explain the high-level idea behind the weighted $\ell_{p}^{p}$ distance query.
+In Section C.4, we show how to answer one-dimensional weighted $\ell_{1}$ distance query.
+In Section C.5, we show how to answer Softmax distance query using previous algorithms.
+In Section C.6, we show how to handle adaptive query.
+By combining the results from these sections, we prove the main results in Section 4.
+C.1 Summation Segment Tree
+First, in order to solve the weighted distance problem, we need to have a basic DP algorithm (Algorithm 2) that can answer simple summation queries.
+After analyzing its DP and error in Section D, we state the data structure theorem.
+Theorem C.1 (DPTree data structure, informal version of Theorem D.1).
+There is a data structure (see DPTree in Algorithm 2) that uses $O(n)$ spaces to support the following operations:
   • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
   • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
+    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
+    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
+    • it holds with probability $0.99$.
+During the design of the data structure, we found an interesting property based on the parallel composition rule of DP Fact B.7. We will now state the lemma, whose proof is provided in Section D.
+Lemma C.2 (Weighted sum of disjoint interval queries, informal version of Lemma D.8).
+If the following conditions hold that:
+  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
+  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
+  • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
+Then, we have Alg. 2 is $(\epsilon,\delta)$-DP and output $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ with the error upper bounded by
+$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
+with probability $0.99$.
+From Lemma C.2, we can see that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries, as long as the sum of squared weights is finite.
+C.2 Sensitivity for Range Summation Problem
+Our DP summation tree data structure DPTree (Algorithm 2) requires sensitivity parameter $\Delta$.
+In this section, we show that for the summation problem, we have the sensitivity $\Delta=2R$ if the input $X\in[-R,R]^{n}$.
+Lemma C.3 (Sensitivity of summation).
+Let $X\in[-R,R]^{n}$. We have the sensitivity $\Delta=2R$ for DPTree.Init in Algorithm 2.
+Proof.
+Let’s say two neighboring datasets $X$ and $X^{\prime}$ differ in $x_{i}$ and $x^{\prime}_{i}$ for some $i$ in the array $X$. Then for a summation problem, i.e. $f(X):=\sum_{i=1}^{n}x_{i}$, we have
+$\Delta=$ $~{}\max_{X,X^{\prime}}|f(X)-f(X^{\prime})|=\max_{X,X^{\prime}}|x_{i}-x^{\prime}_{i}|=2R.$
+where the first step follows from Definition 2.2, the second step follows from $X,X^{\prime}$ differ in $x_{i},x^{\prime}_{i}$, and the last step follows from each coordinate of the dataset is bounded in $[-R,R]$.
+∎
+C.3 Weighted $\ell_{p}^{p}$ Distance Problem
+In this section, we introduce the intuition behind the method for handling the weighted $\ell_{p}^{p}$ distance problem. The formal lemmas and proofs can be found in Section E.
+Given a dataset and a query point in $d$ dimensions, we round each coordinate of the data points and the query point to the nearest multiple of a small interval.
+We then aggregate the weights of data points that have been rounded to the same position.
+Finally, we compute the sum of these aggregated weights multiplied by the distances between the query point and the data points over the rounded positions.
+This approach makes the computation more efficient while maintaining sufficient accuracy.
+We provide an example of weighted $\ell_{1}$-distance of a one-dimensional dataset consisting of 10 data points, i.e., $X\in[0,1]^{10}$ and a query $y=0$ in Figure 1.
+Lemma C.4 (Weighted $\ell_{p}^{p}$-distance high dimension, informal version of Lemma E.2).
+If the following conditions hold:
+  • Let data $X\in[0,R]^{n\times d}$ and $x_{i}^{\top}\in[0,R]^{d}$ be the $i$-th row of $x$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]^{d}$.
+  • We round each dimension of $X$ and $y$ to an integer multiple of $R/n$.
+  • Let $x_{i,k},y_{k}$ denote the $k$-th coordinates of $x_{i},y$ for $k\in[d]$.
+  • Let $c_{j,k}:=\sum_{j_{0}\in S_{j,k}}w_{j_{0}}$ where the set $S_{j,k}$ is the set of index $i$ such that the corresponding $x_{i,k}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
+  • After rounding, we assume that $y_{k}$ is in the $l_{k}R/n$ position for $l_{k}\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
+For the weighted problem, we have
+$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}.$
+where $O(R/n)$ is the rounding error for each data point.
+Remark C.5.
+In Lemma C.4, we first round the dataset.
+This rounding simplifies the calculation by reducing the number of possible positions to consider, from real values in $[0,R]^{d}$ to the total $O(nd)$ spots.
+However, it also introduces an error $O(R/n)$ for one data point.
+Then, for one spot in the rounded dataset, we sum over the weights of that spot and multiply the corresponding distance raised to the power of $p$.
+Additionally, since we are dealing with $\ell_{p}^{p}$ distance, the rounding error is also raised to the power of $p$.
+Figure 1: The visualization of how to build $c_{j}$ of rounded dataset $X\in[0,1]^{10}$ and compute the weighted $\ell_{1}$ distance.
+The number above each $x_{i}$ is $w_{i}$.
+See Algorithm 5 for details.
+Suppose $y=0$.
+Then $\sum_{i=1}^{n}w_{i}|y-x_{i}|=0.1\cdot 2.2+0.3\cdot 3.1+0.3\cdot(-2)+0.3\cdot(-3)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5+0.9\cdot(-1)+0.9\cdot 1=4.4$.
+And $\sum_{j=0}^{n}|k-j|c_{j}/n=0.1\cdot 2.2+0.3\cdot(-1.9)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5=4.4$.
+See details in Lemma E.1.
+C.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
+Based on previous discussions in Section C.1 and C.3, we can now describe our one-dimensional weighted $\ell_{1}$ distance data structure, DPTreeDistance, presented in Algorithm 5 and 6, which generalizes the results from BLM+ [10].
+Drawing from the intuition in Section C.3, the initialization process is as follows: first, we round each data point in the dataset to the nearest multiple of a small interval and build an array that aggregates the corresponding weights.
+This array is then fed into our DPTree data structure in Algorithm 2.
+At query time, we query the DPTree to obtain the aggregated weights within a small interval and multiply these weights by the distance to the query point.
+Furthermore, we also introduce a relative error parameter $\alpha$ to reduce the total number of queries to $O(\log(n)/\alpha)$ instead of querying all $n$ positions.
+We also analyze the DP and the error bound; see details in Section F.
+Theorem C.6 (DPTreeDistance data structure, informal version of Theorem F.6).
+There is a data structure DPTreeDistance (Algorithm 5 and 6) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
+  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
+  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 6) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
+    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
+    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$,
+    • it holds with probability $0.99$.
+C.5 Softmax Activation
+We then describe how we extend the previous results to Softmax activation, i.e.
+exponential inner product function (Definition 4.1).
+From AS [6], we know that Softmax activation can be approximated by polynomial kernel function $P(\cdot)$ with a certain error.
+The following lemma shows that we can transform weighted Softmax queries into polynomial kernels.
+More specifically, we have one term that computes the weighted $\ell_{2}^{2}$ distance, which is the place where we add DP noises.
+Because of the decomposability of the $\ell_{p}^{p}$ distance, i.e.
+$\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p}$, we can easily extend the results of Section C.4 to handle the $\ell_{2}^{2}$ distance query.
+After that, we compute the term for the weighted $\ell_{2}^{2}$ norms of approximation kernel exactly.
+Summing all these terms, with a certain error, we can answer the Softmax query.
+Related proofs can be found in Section I.
+Lemma C.7 (Weighted Softmax approximation, informal version of Lemma I.6).
+Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
+Let $R\geq 1$.
+Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
+Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma I.5.
+Then, we can approximate the exponential inner product using the polynomial kernel function:
+$w^{\top}\exp(Xy/d)=$ $~{}-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
+$+~{}O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
+Moreover, the vectors $P(\cdot)$ can be computed in $O(r)$ time.
+C.6 Adaptive Query
+We introduce how we can modify our algorithm to solve the adaptive query problem using some tools in QRS+ [86].
+Our approach is based on proving that our algorithm can handle any query within the query space with a certain error.
+Since adaptive queries must lie within this space, our algorithm can effectively handle them.
+In Section C.5, we demonstrate our algorithm’s capability to answer weighted Softmax distance queries with constant probability.
+We then use the Chernoff bound to boost the constant probability of our algorithm to a high probability.
+Next, we apply the notion of an $\epsilon_{0}$-net to bound all query points within the net using the union bound.
+Finally, we bound all points in the query space by utilizing the Lipschitz property of the weighted Softmax distance function and introducing an additive error.
+See the proofs in Sections H and I.
+Lemma C.8 (Adaptive Softmax, informal version of Lemma I.10).
+If the following conditions hold:
+  • Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and let $|N|$ be the size of the net $N$.
+  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
+  • Let the relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
+  • We create $l=O(\log((R/\epsilon_{0})^{r}/p_{f}))$ independent copies of the data structure $\{\textsc{DPTreeSoftmax}_{j}\}_{j=1}^{l}$ (Algorithm 3) and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
+  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
+  • Let $Z(y):=w^{\top}\exp(Xy/d)$.
+  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
+Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
+$|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
+Appendix D DPTree Algorithm
+In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 2).
+In Section D.1, we give the theorem (Theorem D.1) of our data structure that can answer summation problem.
+In Section D.2, we improve our data structure from constant probability to high probability by applying Chernoff bound.
+In Section D.3, we give the analysis.
+In Section D.4, we show some results of our data structure if the input queries are disjoint.
+D.1 Single Data Structure
+We give the theorem of our DPTree data structure that can answer the summation problem with DP, accuracy, runtime guarantee.
+Theorem D.1 (DPTree data structure, formal version of Theorem C.1).
+There is a data structure (see DPTree in Algorithm 2) that uses $O(n)$ spaces to support the following operations:
+  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
+  • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
+    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
     • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
@@
 Proof.
-The proofs follow from combining Lemma B.3 (running time of initialization), Lemma B.4 (running time of query), Lemma B.5 (DP of query), and Lemma B.6 (error of query) together. ∎
-B.2 Boost the Constant Probability to High Probability
+The proofs follow from combining Lemma D.3 (running time of initialization), Lemma D.4 (running time of query), Lemma D.5 (DP of query), and Lemma D.6 (error of query) together. ∎
+D.2 Boost the Constant Probability to High Probability
+We can use Chernoff bound to boost the high probability by repeating the data structure multiple times.
+Theorem D.2 (High-probability).
+There is a data structure (see DPTreeHighProb in Algorithm 4) that uses $O(n\log(1/\delta_{\mathrm{fail}}))$ spaces to support the following operations
+  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1),\delta_{\mathrm{fail}}\in(0,0.01))$.
+It takes $O(n\log(1/\delta_{\mathrm{fail}}))$ time to initialize the data structure.
+  • Query$(x\in[n],y\in[n])$. It takes $O(\log(n)\cdot\log(1/\delta_{\mathrm{fail}}))$ time to output a number $z$ such that
+    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
+    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$,
+    • it holds with probability $1-\delta_{\mathrm{fail}}$ for failure probability $\delta_{\mathrm{fail}}\in(0,0.01)$.
+Proof.
+Note that our data structure (Theorem D.1) succeeds with probability $0.99$.
+The success of the algorithm (Theorem D.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma B.2).
+By repeating the data structure $O(\log(1/\delta_{\mathrm{fail}}))$ times and taking the median of the outputs, we boost the success probability.
+The details are following.
+To boost the success probability, we assume the query is repeated $l$ times.
+Let $i\in[l]$, and let $z_{i}$ denote the indicator random variable for the success of the $i$-th instance of the data structure for a single query.
+Let $z=\sum_{i=1}^{l}z_{i}$ be the total success times.
+Since $p=\Pr[z_{i}=1]=0.99$, we can have $\mu=\operatorname*{{\mathbb{E}}}[z]=\sum_{i=1}^{l}p=lp$.
+Note that $p=0.99$.
+By setting $\delta=0.1$ and using Chernoff bound from Lemma B.2, we can show
+$\Pr[z\leq l/2]\leq\Pr[z\leq(1-\delta)lp]\leq\exp(-\delta^{2}lp/2).$
+Note that we want $z>l/2$ (since we want at least half to succeed so we could take the median),
+$\Pr[z>l/2]\geq 1-\exp(-\delta^{2}lp/2).$
+To ensure that failure probability is $\delta_{\mathrm{fail}}$, we have
+$\exp(-\delta^{2}lp/2)=\delta_{\mathrm{fail}}.$
+We can make this hold by choosing $l=O(\log(1/\delta_{\mathrm{fail}}))$.
+By the DP basic composition rule (Fact B.6), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem D.1.
+∎
 Algorithm 4 Boost constant probability
-1:datastructure DPTreeHighProb $\triangleright$ Theorem B.2
+1:datastructure DPTreeHighProb $\triangleright$ Theorem D.2
 2:members
-3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(\log(1/\delta_{\mathrm{fail}}))}:\textsc{DPTree}$triangleright Alg. 1
+3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(\log(1/\delta_{\mathrm{fail}}))}:\textsc{DPTree}$triangleright Alg. 2
 4:end members
@@
 17:end datastructure
-We can use Chernoff bound to boost the high probability by repeating the data structure multiple times.
-Theorem B.2 (High-probability).
-There is a data structure (see DPTreeHighProb in Algorithm 5) that uses $O(n\log(1/\delta_{\mathrm{fail}}))$ spaces to support the following operations
-  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1),\delta_{\mathrm{fail}}\in(0,0.01))$.
-It takes $O(n\log(1/\delta_{\mathrm{fail}}))$ time to initialize the data structure.
-  • Query$(x\in[n],y\in[n])$. It takes $O(\log(n)\cdot\log(1/\delta_{\mathrm{fail}}))$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i=x}^{y}a_{i}$,
-    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$,
-    • it holds with probability $1-\delta_{\mathrm{fail}}$ for failure probability $\delta_{\mathrm{fail}}\in(0,0.01)$.
-Proof.
-Note that our data structure (Theorem B.1) succeeds with probability $0.99$.
-The success of the algorithm (Theorem B.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma A.2).
-By repeating the data structure $O(\log(1/\delta_{\mathrm{fail}}))$ times and taking the median of the outputs, we boost the success probability.
-The details are following.
-To boost the success probability, we assume the query is repeated $l$ times.
-Let $i\in[l]$, and let $z_{i}$ denote the indicator random variable for the success of the $i$-th instance of the data structure for a single query.
-Let $z=\sum_{i=1}^{l}z_{i}$ be the total success times.
-Since $p=\Pr[z_{i}=1]=0.99$, we can have $\mu=\operatorname*{{\mathbb{E}}}[z]=\sum_{i=1}^{l}p=lp$.
-Note that $p=0.99$.
-By setting $\delta=0.1$ and using Chernoff bound from Lemma A.2, we can show
-$\Pr[z\leq l/2]\leq\Pr[z\leq(1-\delta)lp]\leq\exp(-\delta^{2}lp/2).$
-Note that we want $z>l/2$ (since we want at least half to succeed so we could take the median),
-$\Pr[z>l/2]\geq 1-\exp(-\delta^{2}lp/2).$
-To ensure that failure probability is $\delta_{\mathrm{fail}}$, we have
-$\exp(-\delta^{2}lp/2)=\delta_{\mathrm{fail}}.$
-We can make this hold by choosing $l=O(\log(1/\delta_{\mathrm{fail}}))$.
-By the DP basic composition rule (Fact 3.10), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem B.1.
-∎
-B.3 Algorithm of Data Structure
-In this section, we analyze the accuracy, DP, and runtime of Algorithm 1 and 2.
+D.3 Algorithm of Data Structure
+In this section, we analyze the accuracy, DP, and runtime of Algorithm 2.
 We first analyze the runtime.
-Lemma B.3 (Runtime of initialization, Algorithm 1).
+Lemma D.3 (Runtime of initialization, Algorithm 2).
 For the initialization, we have the time complexity of Algorithm 2 is $O(n)$.
@@
 All the computations are dominated by $O(n)$ time. ∎
-Lemma B.4 (Runtime of query, Algorithm 2).
+Lemma D.4 (Runtime of query, Algorithm 2).
 For each query, we have the time complexity of Algorithm 2 is $O(\log n)$.
@@
 We now analyze the DP.
-Lemma B.5 (Privacy of query, Algorithm 2).
-The output of Query (see Algorithm 2) is $(\epsilon,\delta)$-DP.
+Lemma D.5 (Privacy of query, Algorithm 2).
+The output process of Query (see Algorithm 2) is $(\epsilon,\delta)$-DP.
 Proof.
 Suppose that our dataset is $X\in[-R,R]^{n}$.
-Note that we only add noise into the pre-processing stage.
+Note that we only add noise in the pre-processing stage.
 There is no noise in the query stage.
-Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma 6.3).
+Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma C.3).
 Since we add noise to each node in the tree, and each leaf node count will contribute to $\log n$ nodes, it is equivalent to our output function being in $\log n$ dimension.
 We will then blow up the DP parameter by $\log n$ factor.
-Thus, using the basic composition rule (Fact 3.10), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP.
+Thus, using the basic composition rule (Fact B.6), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP.
 ∎
 We now analyze the accuracy.
-Lemma B.6 (Accuracy of query, Algorithm 2).
-Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 3.7, we have the error of Query(see Algorithm 2) output is upper bounded by:
+Lemma D.6 (Accuracy of query, Algorithm 2).
+Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 2.5, we have the error of Query(see Algorithm 2) output is upper bounded by:
 $O(\epsilon^{-1}\Delta\log^{3/2}n).$
@@
 We know $\mu=\operatorname*{{\mathbb{E}}}[z]=0$ since $\operatorname*{{\mathbb{E}}}[z_{i}]=0$.
-From Fact 3.7, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
+From Fact 2.5, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
 Therefore, we can show
@@
 Note that we wish to bound $|z|=|\textsc{Query}(S)-\textsc{TrueQuery}(S)|$ as our error.
-Using Lemma A.3, we can have
+Using Lemma B.3, we can have
 $\Pr[|z|\geq k\sigma]\leq\frac{1}{k^{2}}.$
@@
 Thus, we conclude that error is bounded by $O(c^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)=O(\epsilon^{-1}\Delta\log^{3/2}n)$ (since $c\in(0,2]$) with probability $0.99$. ∎
-B.4 Disjoint Intervals
+D.4 Disjoint Intervals
 In this section, we show some interesting results for our DPTree data structure if the input queries are disjoint.
-Lemma B.7 (Sum of disjoint interval queries).
+Lemma D.7 (Sum of disjoint interval queries).
 If the following conditions hold that:
@@
   • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
-Then, we have $\sum_{j=1}^{t}\textsc{Query}(S_{j})$ (see Algorithm 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
+Then, we have Algorithm 2 is $(\epsilon,\delta)$-DP and outputs $\sum_{j=1}^{t}\textsc{Query}(S_{j})$ with the error upper bounded by
 $O(t^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
@@
 Proof.
-From Lemma B.5, we know that $\mathrm{DPTree}$.Query returns a $(\epsilon,\delta)$-DP output.
-Then, from Fact 3.11 and the disjoint intervals in Algorithm 7, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
+From Lemma D.5, we know that $\mathrm{DPTree}$.Query is $(\epsilon,\delta)$-DP.
+Then, from Fact B.7 and the disjoint intervals in Algorithm 6, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
 Let $\textsc{TrueQuery}(S_{j})$ denote the true interval query answer returned by DPTree.TrueQuery in Algorithm 2 for interval $S_{j}$.
@@
 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}z_{j}$.
-From the proof of Lemma B.6, we know $z_{j}$ is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$ and the variance is bounded by
+From the proof of Lemma D.6, we know $z_{j}$ is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$ and the variance is bounded by
 $\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
@@
 Note that we wish to bound $|z|$ as our error.
-Using Lemma A.3, we can have error bounded by
+Using Lemma B.3, we can have error bounded by
 $O(t^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
@@
 Moreover, this can be generalized to weighted sum of queries.
-Lemma B.8 (Weighted sum of disjoint interval queries, formal version of Lemma 6.2).
+Lemma D.8 (Weighted sum of disjoint interval queries, formal version of Lemma C.2).
 If the following conditions hold that:
@@
   • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
-Then, we have $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ (see Algorithm 2) is $(\epsilon,\delta)$-DP and the error is upper bounded by
+Then, we have Alg. 2 is $(\epsilon,\delta)$-DP and output $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ with the error upper bounded by
 $O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
@@
 Proof.
-The DP proof is the same as in the proof of Lemma B.7.
-Let $\textsc{TrueQuery}(S_{j})$ and $\textsc{Query}(S_{j})$ be same in the proof of Lemma B.7 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}a_{j}z_{j}$.
-From the proof of Lemma B.7, we know the variance of $z_{j}$ is bounded by
+The DP proof is the same as in the proof of Lemma D.7.
+Let $\textsc{TrueQuery}(S_{j})$ and $\textsc{Query}(S_{j})$ be same in the proof of Lemma D.7 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}a_{j}z_{j}$.
+From the proof of Lemma D.7, we know the variance of $z_{j}$ is bounded by
 $\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
@@
 Note that we wish to bound $|z|$ as our error.
-Using Lemma A.3, we can have error bounded by
+Using Lemma B.3, we can have error bounded by
 $O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
 with probability $0.99$. ∎
-Appendix C Weighted $\ell_{p}^{p}$ Distance
+Appendix E Weighted $\ell_{p}^{p}$ Distance
 In this section, we introduce how to handle weighted $\ell_{p}^{p}$ distance problem in the high level idea.
-In Section C.1, we show how to solve one dimensional weighted problem.
-In Section C.2, we show how to solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset.
+In Section E.1, we show how to solve one dimensional weighted problem.
+In Section E.2, we show how to solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset.
 Suppose we have the original data $X\in[0,R]^{n}$ and weight $w\in\mathbb{R}^{n}$ and query $y\in[0,R]$. We want to compute the weighted $\ell_{1}$-distance, i.e.
@@
 Therefore, we can solve one dimension problem first, and then the high dimension case can be solved automatically.
-C.1 One Dimensional Weighted Distance
+E.1 One Dimensional Weighted Distance
 Now we can give the lemma for weighted distance of dataset.
-Lemma C.1 (Weighted distance one dimension).
+Lemma E.1 (Weighted distance one dimension).
 If the following conditions hold:
@@
 Moreover, we have
-$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(|k-j|R/n+O(R/n))^{p}c_{j}.$
+$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(|k-j|R/n+O(R/n))^{p}c_{j}$
 where $O(R/n)$ is the rounding error for each data point.
@@
 $\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(\frac{|k-j|R}{n}+O(\frac{R}{n}))^{p}c_{j}.$
-C.2 High Dimensional Weighted Distance
+E.2 High Dimensional Weighted Distance
 Finally, we can solve the problem of weighted distance for $d$-dimensional dataset.
-Lemma C.2 (Weighted $\ell_{p}^{p}$-distance high dimension, formal version of Lemma 6.4).
+Lemma E.2 (Weighted $\ell_{p}^{p}$-distance high dimension, formal version of Lemma C.4).
 If the following conditions hold:
@@
 For the weighted problem, we have
-$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}.$
+$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}$
 where $O(R/n)$ is the rounding error for each data point.
@@
 $=$ $~{}\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}$
-where the first step follows from decomposability of $\ell_{p}^{p}$-distance by dimension, the second step follows from Lemma C.1.
-Appendix D One-Dimensional Weighted $\ell_{1}$ Distance Query
-In this section, we generalize the algorithms in [8] to weighted distance.
+where the first step follows from decomposability of $\ell_{p}^{p}$-distance by dimension, the second step follows from Lemma E.1.
+Appendix F One-Dimensional Weighted $\ell_{1}$ Distance Query
+In this section, we generalize the algorithms in BLM+ [10] to weighted distance.
 Here, we compute the problem of one-dimensional weighted $\ell_{1}$ distance query i.e.
 $\sum_{i\in[n]}w_{i}|y-x_{i}|$ for a given query $y\in[0,R]$, weights $w\in[-R_{w},R_{w}]^{n}$ and dataset $X\subset[0,R]$ and $n=|X|$.
-In Section D.1, we analyze the runtime of our algorithm.
-In Section D.2, we analyze the DP and accuracy of our algorithm.
-In Section D.3, we give the theorem for our DPTreeDistance data structure.
+In Section F.1, we analyze the runtime of our algorithm.
+In Section F.2, we analyze the DP and accuracy of our algorithm.
+In Section F.3, we give the theorem for our DPTreeDistance data structure.
 Algorithm 5 Pre-processing data structure
-1:datastructure DPTreeDistance $\triangleright$ Theorem D.6
+1:datastructure DPTreeDistance $\triangleright$ Theorem F.6
 2:members
-3:     ${\mathcal{D}}:$ DPTree $\triangleright$ Alg. 1
+3:     ${\mathcal{D}}:$ DPTree $\triangleright$ Alg. 2
 4:     $X:[0,R]^{n}$
@@
 6:end members
-7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma C.1
+7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma E.1
 8:     $X,w,a\leftarrow X,w,0^{n+1}$
@@
 12:     end for
-13:     ${\mathcal{D}}$.$\textsc{Init}(a,n+1,2R_{w},\epsilon,\delta)$triangleright Alg. 1, Lemma 6.3
+13:     ${\mathcal{D}}$.$\textsc{Init}(a,n+1,2R_{w},\epsilon,\delta)$triangleright Alg. 2, Lemma C.3
 14:end procedure
@@
 Algorithm 6 One dimensional weighted $\ell_{1}$ distance query
-1:datastructure DPTreeDistance $\triangleright$ Theorem D.6
-2:procedure DistanceQuery($y\in[0,R]$, $\alpha\in(0,1)$) $\triangleright$ Lemma D.2, Lemma D.4, Lemma D.5
-3:     $y\leftarrow\textsc{Round}(y,n)\cdot(R/n)$triangleright Alg. 6
+1:datastructure DPTreeDistance $\triangleright$ Theorem F.6
+2:procedure DistanceQuery($y\in[0,R]$, $\alpha\in(0,1)$) $\triangleright$ Lemma F.2, Lemma F.4, Lemma F.5
+3:     $y\leftarrow\textsc{Round}(y,n)\cdot(R/n)$triangleright Alg. 5
 4:     Value $\leftarrow 0$
@@
 17:end datastructure
-D.1 Runtime Analysis
+F.1 Runtime Analysis
 We first analyze the runtime.
-Lemma D.1 (Runtime of initialization, Algorithm 6).
+Lemma F.1 (Runtime of initialization, Algorithm 5).
 For the initialization, we have the time complexity of Init in Algorithm 5 is $O(n)$.
@@
 In the initialization of Init, the computations need $O(n)$ time to compute the count and $O(\log n)$ time to build the tree. Thus, total time is $O(n)$. ∎
-Lemma D.2 (Runtime of DistanceQuery, Algorithm 7).
+Lemma F.2 (Runtime of DistanceQuery, Algorithm 6).
 For the $\ell_{1}$ distance query, we have the time complexity of DistanceQuery in Algorithm 6 is $O(\alpha^{-1}\log^{2}n)$.
@@
 In DistanceQuery, the computations need $O(\log n)$ time to compute one value from DPTree.Query and this process need to be repeated $O(\alpha^{-1}\log n)$ times. ∎
-Remark D.3.
-In Line 8 and 13 of Algorithm 7, we use $R/(1+\alpha)^{j}$ to approximate the distance of each data point to the query in Lemma C.1, i.e.
+Remark F.3.
+In Line 8 and 13 of Algorithm 6, we use $R/(1+\alpha)^{j}$ to approximate the distance of each data point to the query in Lemma E.1, i.e.
 $|k-j|R/n$.
 This will introduce $\alpha$ relative error but also reduce the numbers of iteration from $O(n)$ to $O(\log(n)/\alpha)$.
-D.2 Privacy and Accuracy Analysis
+F.2 Privacy and Accuracy Analysis
 We show the DP.
-Lemma D.4 (Privacy of DistanceQuery, Algorithm 7).
-The output of DistanceQuery (Algorithm 7) is $(\epsilon,\delta)$-DP.
+Lemma F.4 (Privacy of DistanceQuery, Algorithm 6).
+The output process of DistanceQuery (Algorithm 6) is $(\epsilon,\delta)$-DP.
 Proof.
-From Lemma B.5, we know that $\mathrm{DPTree}$.Query returns a $(\epsilon,\delta)$-DP output.
+From Lemma D.5, we know that $\mathrm{DPTree}$.Query is $(\epsilon,\delta)$-DP output.
 We observe that intervals in Algorithm 6 are disjoint.
-Then, following the same logic in the proof of Lemma B.8, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
+Then, following the same logic in the proof of Lemma D.8, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
 ∎
 We now analyze the accuracy of the algorithm.
-Lemma D.5 (Accuracy of DistanceQuery, Algorithm 7).
+Lemma F.5 (Accuracy of DistanceQuery, Algorithm 6).
 If the following conditions are satisfied:
@@
 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$, which from Algorithm 2 we can see this is the sum of $O(\log n)$ independent truncated Laplace random variables.
-From Lemma B.8, we only need to show that the series $\frac{1}{(1+\alpha)^{j}}$ for $j\in\{0,1,\dots,O(\log(n)/\alpha)\}$ square converges to $1/\alpha$, since $R$ is a constant.
+From Lemma D.8, we only need to show that the series $\frac{1}{(1+\alpha)^{j}}$ for $j\in\{0,1,\dots,O(\log(n)/\alpha)\}$ square converges to $1/\alpha$, since $R$ is a constant.
 We can show
@@
 where the first step follows from we extend the finite sum to infinite sum, the second step follows from $\frac{1}{(1+\alpha)^{2j}}\leq\frac{1}{(1+\alpha)^{j}}$, the third step follows from the closed form of geometric sum, the fourth step follows from simple algebra, and the last step follows from $\alpha\in(0,1)$.
-Then from the proof of Lemma B.8, we can know that the variance is given by
+Then from the proof of Lemma D.8, we can know that the variance is given by
 $O(\frac{R^{2}R_{w}^{2}\log^{3}n}{\alpha\epsilon^{2}})$ (5)
-since the sensitivity $\Delta=2R_{w}$ from Lemma 6.3.
-Using Lemma A.3, we can have additive error bounded by
+since the sensitivity $\Delta=2R_{w}$ from Lemma C.3.
+Using Lemma B.3, we can have additive error bounded by
 $O(\frac{R\cdot R_{w}\log^{3/2}n}{\epsilon\sqrt{\alpha}}).$
 with probability $0.99$. ∎
-D.3 One Dimension Single Data Structure
+F.3 One Dimension Single Data Structure
 We therefore have the data structure that can solve weighted $\ell_{1}$-distance problem.
-Theorem D.6 (DPTreeDistance data structure, formal version of Theorem 6.6).
-There is a data structure DPTreeDistance (Algorithm 6,7) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
+Theorem F.6 (DPTreeDistance data structure, formal version of Theorem C.6).
+There is a data structure DPTreeDistance (Algorithm 5,6) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
   • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
   • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 6) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
+    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
     • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$,
@@
 Proof.
-The proofs follow from combining Lemma D.1 (running time of initialization), Lemma D.2 (running time of query), Lemma D.4 (DP of query), and Lemma D.5 (error of query) together. ∎
-Appendix E High-Dimensional Weighted $\ell_{1}$ Query
-In this section, we show how we can solve the high dimensional weighted $\ell_{1}$ distance problem, generalizing results from [8].
-In Section E.1, we give the analysis of Algorithm 8.
-In Section E.2, we give the theorem of our DPTreeHighDim data structure.
-Algorithm 6,7 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{1}$ distance function.
+The proofs follow from combining Lemma F.1 (running time of initialization), Lemma F.2 (running time of query), Lemma F.4 (DP of query), and Lemma F.5 (error of query) together. ∎
+Appendix G High-Dimensional Weighted $\ell_{1}$ Query
+In this section, we show how we can solve the high dimensional weighted $\ell_{1}$ distance problem, generalizing results from BLM+ [10].
+In Section G.1, we give the analysis of Algorithm 7.
+In Section G.2, we give the theorem of our DPTreeHighDim data structure.
+Algorithm 5,6 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{1}$ distance function.
 We construct $d$ separate one-dimensional distance query data structures, each corresponding to a coordinate projection of the dataset.
-E.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
+G.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
 We now give the analysis of our Algorithm 7 for high dimensional weighted $\ell_{1}$-distance query.
 Algorithm 7 High-dimensional weighted $\ell_{1}$ distance query
-1:datastrucutre DPTreeHighDim $\triangleright$ Theorem E.3
+1:datastrucutre DPTreeHighDim $\triangleright$ Theorem G.3
 2:members
-3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$triangleright Alg. 6
+3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$triangleright Alg. 5
 4:     $X:[0,R]^{n\times d}$
@@
 6:end members
-7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$) $\triangleright$ Lemma C.2
+7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$) $\triangleright$ Lemma E.2
 8:     $X\leftarrow X$
@@
 10:     for $i=1\to d$ do
-11:         $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 6
+11:         $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 5
 12:     end for
 13:end procedure
-14:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma E.1, Lemma E.2
+14:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma G.1, Lemma G.2
 15:     Value $\leftarrow 0$
 16:     for $i=1\to d$ do
-17:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(y_{i},\alpha)$triangleright Alg. 7
+17:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(y_{i},\alpha)$triangleright Alg. 6
 18:     end for
@@
 21:end datastrucutre
-Lemma E.1 (Privacy of DistanceQuery, Algorithm 8).
+Lemma G.1 (Privacy of DistanceQuery, Algorithm 7).
 If the following conditions hold
@@
   • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ represent the true distance query value.
-  • Let $\epsilon\leq O(\log(1/\delta^{\prime}))$.
-Then, we have the output of DistanceQuery (Algorithm 8) is $(\epsilon,\delta+\delta^{\prime})$-DP.
+  • Let $\epsilon=O(\log(1/\delta^{\prime}))$.
+Then, we have the output process of DistanceQuery (Algorithm 7) is $(\epsilon,\delta+\delta^{\prime})$-DP.
 Proof.
-The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem 3.12.
+The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem B.8.
 Our algorithm instantiate each one-dimensional data structure with $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP total $d$ times.
-From advanced composition in Theorem 3.12, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
+From advanced composition in Theorem B.8, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
 $O(c\epsilon\sqrt{2d\log(1/\delta^{\prime})}/\sqrt{d\log(1/\delta^{\prime})})=O(\epsilon)$
@@
 $d\delta/d+\delta^{\prime}=\delta+\delta^{\prime}.$
-Lemma E.2 (Accuracy of DistanceQuery, Algorithm 8).
+Lemma G.2 (Accuracy of DistanceQuery, Algorithm 7).
 If the following conditions hold
@@
 Observe that $A_{*}=\sum_{i=1}^{d}A_{*,i}$ and $A=\sum_{i=1}^{d}A_{i}$.
-We follow the similar idea in the proof of Lemma D.5.
-Let $z_{j,i}$ be the random variables that represent $z_{j}$ (used in the proof of Lemma D.5) for the $i$-th coordinate.
+We follow the similar idea in the proof of Lemma F.5.
+Let $z_{j,i}$ be the random variables that represent $z_{j}$ (used in the proof of Lemma F.5) for the $i$-th coordinate.
 We can observe that the overall error across $d$ coordinates can be upper bounded by
@@
 $O(\alpha^{-1}\epsilon^{-2}d^{2}R^{2}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
-Finally, from Lemma A.3, we have the additive error given by
+Finally, from Lemma B.3, we have the additive error given by
 $O(\alpha^{-1/2}\epsilon^{-1}dRR_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
-E.2 High Dimension Single Data Structure
+G.2 High Dimension Single Data Structure
 We have the data structure that can solve weighted $\ell_{1}$-distance problem in $d$-dimensional data.
-Theorem E.3 (DPTreeHighDim data structure).
+Theorem G.3 (DPTreeHighDim data structure).
 There is a data structure DPTreeHighDim (Algorithm 7) that uses $O(nd)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
@@
   • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 7) It takes $O(\alpha^{-1}d\log^{2}n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$,
+    • the process of output $z$ satisfies is $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$,
     • $|z-\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}|\leq\alpha\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}d\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
 For the runtime analysis, since we loop data structure DPTreeDistance $d$ times, an additional $d$ factor will appear for both initialization and query time complexity.
-The DP is proved by Lemma E.1.
-The accuracy is proved by Lemma E.2.
+The DP is proved by Lemma G.1.
+The accuracy is proved by Lemma G.2.
 ∎
-Appendix F Adaptive Query
-In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from [75].
+Appendix H Adaptive Query
+In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from QRS+ [86].
 Our idea is that, if we can prove that our algorithm can solve any query in the query space with certain error.
 Then, since adaptive query must lie in this space, we can handle adaptive query.
-In Section F.1, we show how we can boost the constant probability of our algorithm to high probability.
-In Section F.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net.
-In Section F.3, we show how we can bound all points in the query space by introducing an additive error.
-In Section F.4, we examine the effects of different norms on our adaptive query proof.
-First, from Theorem E.3, given query $y\in[0,R]^{d},\alpha\in(0,1)$ we have DistanceQuery$(y,\alpha)$ that can solve $d$-dimension weighted $\ell_{1}$-distance problem with constant probability $0.99$.
+In Section H.1, we show how we can boost the constant probability of our algorithm to high probability.
+In Section H.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net.
+In Section H.3, we show how we can bound all points in the query space by introducing an additive error.
+In Section H.4, we examine the effects of different norms on our adaptive query proof.
+First, from Theorem G.3, given query $y\in[0,R]^{d},\alpha\in(0,1)$ we have DistanceQuery$(y,\alpha)$ that can solve $d$-dimension weighted $\ell_{1}$-distance problem with constant probability $0.99$.
 Now we show how to improve it to solve adaptive query problem.
-F.1 Boost the Constant Probability to High Probability
-We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma A.2.
-Lemma F.1 (Using Chernoff bound to boost the probability).
+H.1 Boost the Constant Probability to High Probability
+We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma B.2.
+Lemma H.1 (Using Chernoff bound to boost the probability).
 If the following conditions hold:
@@
   • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
-Then for each fixed query point $y$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by
-$\alpha A_{*}+B$
-with probability $1-p_{f}$.
+Then for each fixed query point $y$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}$.
 Proof.
-By basic composition Fact 3.10, we prove the DP. Similar to the proof of Theorem B.2, we prove the error by Chernoff bound (Lemma A.2). ∎
-F.2 From Each Fixed Query Point to All On-net Points
+By basic composition Fact B.6, we prove the DP. Similar to the proof of Theorem D.2, we prove the error by Chernoff bound (Lemma B.2). ∎
+H.2 From Each Fixed Query Point to All On-net Points
 In this section, we build $\epsilon_{0}$-net and generalize from each fixed query point to all on-net points.
-Definition F.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in [85]).
+Definition H.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in Ver [96]).
 We define $N$ be $\ell_{p}$ $\epsilon_{0}$-net of :=${\mathcal{B}}:=\{q\in[0,R]^{d}\}$ such that, for every point $q$ in ${\mathcal{B}}$, there exists $y\in N$ satisfying $\|y-q\|_{p}\leq\epsilon_{0}$.
-Fact F.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
+Fact H.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
 Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$. We have $|N|\leq(5R/\epsilon_{0})^{d}$.
-Fact F.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in [88]).
+Fact H.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in Woo [102]).
 Let $N$ be the $\ell_{2}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$. We have $|N|\leq(5R/\epsilon_{0})^{d}$.
-Fact F.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in [43]).
+Fact H.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in GS [41]).
 Let $N$ be the $\ell_{1}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and $|N|$ be the size of net $N$. We have $|N|\leq(5R\sqrt{d}/\epsilon_{0})^{d}$.
-Lemma F.6 (From for each query point to for all points in net).
+Lemma H.6 (From for each query point to for all points in net).
 If the following conditions hold:
@@
   • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
-Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by
-$\alpha A_{*}+B.$
+Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $\alpha A_{*}+B$.
 Proof.
-By basic composition Fact 3.10, we prove the DP. From Lemma F.1, we know for each $y\in N$, the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}/|N|$.
+By basic composition Fact B.6, we prove the DP. From Lemma H.1, we know for each $y\in N$, the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}/|N|$.
 Then, by union bound, with probability $1-p_{f}$, the error of all $|N|$ query points in the net $y\in N$ is upper bounded by $\alpha A_{*}+B$. ∎
-F.3 From Net Points to All Points
+H.3 From Net Points to All Points
 In this section, we show how to generalize points from net to all points in the query space. Since adaptive query must lie in this space, we complete the proof of adaptive query.
-Lemma F.7 (Lipschitz of query function).
+Lemma H.7 (Lipschitz of query function).
 If the following conditions hold:
@@
 ∎
-Lemma F.8 (From points in net to all points in query space).
+Lemma H.8 (From points in net to all points in query space).
 If the following conditions hold:
@@
   • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
-Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
+Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
 $|f(y)-Z(q)|\leq\alpha Z(q)+B+2Ld\epsilon_{0}.$
 Proof.
-By basic composition Fact 3.10, we prove the DP.
+By basic composition Fact B.6, we prove the DP.
 We define an event $E$ such that:
@@
 $~{}|f(y)-Z(y)|\leq\alpha Z(y)+B.$
-From Lemma F.1, with $l=O(\log(|N|/p_{f}))$ we know
+From Lemma H.1, with $l=O(\log(|N|/p_{f}))$ we know
 $\Pr[\mathrm{event}~{}E~{}\mathrm{holds}]\geq 1-p_{f}$
@@
 $=$ $~{}O(\log((R/\epsilon_{0})^{d}/p_{f})$
-where the first step follows from definition of $l$, the second step follows from Fact F.3.
-We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition F.2), for each $q\notin N$, there exists $y\in N$ such that
+where the first step follows from definition of $l$, the second step follows from Fact H.3.
+We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition H.2), for each $q\notin N$, there exists $y\in N$ such that
 $\|y-q\|_{\infty}\leq\epsilon_{0}$ (6)
@@
 $\leq$ $~{}L\cdot d\epsilon_{0}$ (7)
-where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (3).
+where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (6).
 Using the on-net query $y$ to answer the off-net query $q$, for any $q\notin N$, we have
@@
 $\leq$ $~{}\alpha Z(q)+B+2L\cdot d\cdot\epsilon_{0}$ (8)
-where the first step follows from triangular inequality, the second step follows from Eq. (F.3), the third step follows from Lemma F.6, and the last step follows from Eq. (F.3).
+where the first step follows from triangular inequality, the second step follows from Eq. (H.3), the third step follows from Lemma H.6, and the last step follows from Eq. (H.3).
 Thus, we complete the proof. ∎
 Therefore, even adaptive queries can be answered accurately, since any adaptive query can be assumed in ${\mathcal{B}}$.
-F.4 Effect of Different Norms on the Result
-In the above proof, we have two different measure spaces, i.e. $\ell_{\infty}$ distance of $\epsilon_{0}$-net (Definition F.2) and $\ell_{1}$ Lipschitz (Lemma F.7).
+H.4 Effect of Different Norms on the Result
+In the above proof, we have two different measure spaces, i.e. $\ell_{\infty}$ distance of $\epsilon_{0}$-net (Definition H.2) and $\ell_{1}$ Lipschitz (Lemma H.7).
 One might ask, will the norm we choose in two spaces have an impact on the final result? We can show that the norm we choose currently is sufficient to use.
-For different norms, the only differences in the proofs will be Lipschitz smoothness in Eq. (F.3) and the cardinality of $\epsilon_{0}$-net, i.e. $|N|$ in Fact F.3.
-Lemma F.9.
-If we use $\ell_{\infty}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
-$l=O(d\log(nR/p_{f}))$
+For different norms, the only differences in the proofs will be Lipschitz smoothness in Eq. (H.3) and the cardinality of $\epsilon_{0}$-net, i.e. $|N|$ in Fact H.3.
+Lemma H.9.
+If we use $\ell_{\infty}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(nR/p_{f}))$.
 Proof.
-If we use $\ell_{\infty}$ to bound the distance to net, Eq. (F.3) is:
+If we use $\ell_{\infty}$ to bound the distance to net, Eq. (H.3) is:
 $|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
@@
 $\leq$ $~{}nR_{w}\cdot d\epsilon_{0}$
-where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{\infty}$ $\epsilon_{0}$-net.
-Then, Eq. (F.3) is
+where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{\infty}$ $\epsilon_{0}$-net.
+Then, Eq. (H.3) is
 $|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot d\cdot\epsilon_{0}$
-For $\ell_{\infty}$ distance, we have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact F.3.
-We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot d\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
+For $\ell_{\infty}$ distance, we have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact H.3.
+We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot d\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
 $l=$ $~{}O(\log(|N|/p_{f})$
@@
 where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
-Lemma F.10.
-If we use $\ell_{2}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
-$l=O(d\log(nR/p_{f}))$
+Lemma H.10.
+If we use $\ell_{2}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(nR/p_{f}))$.
 Proof.
-If we use $\ell_{2}$ to bound the distance to net, Eq. (F.3) changes to be:
+If we use $\ell_{2}$ to bound the distance to net, Eq. (H.3) changes to be:
 $|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
@@
 $\leq$ $~{}nR_{w}\cdot\epsilon_{0}\sqrt{d}$
-where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq\sqrt{d}\cdot\|x\|_{2}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{2}$ $\epsilon_{0}$-net.
-Then, Eq. (F.3) changes to be
+where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq\sqrt{d}\cdot\|x\|_{2}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{2}$ $\epsilon_{0}$-net.
+Then, Eq. (H.3) changes to be
 $|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot\epsilon_{0}\sqrt{d}$
-For $\ell_{2}$ distance, we also have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact F.4.
-We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\sqrt{d}\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
+For $\ell_{2}$ distance, we also have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact H.4.
+We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\sqrt{d}\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
 $l=$ $~{}O(\log(|N|/p_{f})$
@@
 where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
-Lemma F.11.
-If we use $\ell_{1}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma F.8, we have copies of data structure
-$l=O(d\log(nRd/p_{f}))$
+Lemma H.11.
+If we use $\ell_{1}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(ndR/p_{f}))$.
 Proof.
-If we use $\ell_{1}$ to bound the distance to net, Eq. (F.3) changes to be:
+If we use $\ell_{1}$ to bound the distance to net, Eq. (H.3) changes to be:
 $|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
 $\leq$ $~{}nR_{w}\cdot\epsilon_{0}$
-where the first step follows from Lemma F.7, and the last step follows from $\ell_{1}$ $\epsilon_{0}$-net.
-Then, Eq. (F.3) changes to be
+where the first step follows from Lemma H.7, and the last step follows from $\ell_{1}$ $\epsilon_{0}$-net.
+Then, Eq. (H.3) changes to be
 $|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot\epsilon_{0}$
 For $\ell_{1}$ distance, we have $|N|\leq(5R\sqrt{d}/\epsilon_{0})^{d}$.
-We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\epsilon_{0}$ term in $B$ in Lemma F.8. Thus,
+We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
 $l=$ $~{}O(\log(|N|/p_{f})$
@@
 From the above analysis, we can show that $\ell_{\infty}$ or $\ell_{2}$ $\epsilon_{0}$-net is slightly better than $\ell_{1}$ $\epsilon_{0}$-net.
-  • $\ell_{\infty}$ $\epsilon_{0}$-net, Lemma F.9: we have $l=O(d\log(nR/p_{f}))$.
-  • $\ell_{2}$ $\epsilon_{0}$-net, Lemma F.10: we have $l=O(d\log(nR/p_{f}))$.
-  • $\ell_{1}$ $\epsilon_{0}$-net, Lemma F.11: we have $l=O(d\log(nRd/p_{f}))$.
+  • $\ell_{\infty}$ $\epsilon_{0}$-net, Lemma H.9: we have $l=O(d\log(nR/p_{f}))$.
+  • $\ell_{2}$ $\epsilon_{0}$-net, Lemma H.10: we have $l=O(d\log(nR/p_{f}))$.
+  • $\ell_{1}$ $\epsilon_{0}$-net, Lemma H.11: we have $l=O(d\log(nRd/p_{f}))$.
 Thus, the norm we choose for $\epsilon_{0}$-net is sufficient good.
-Appendix G Softmax Activation
+Appendix I Softmax Activation
 In this section, we introduce how we extend previous $\ell_{1}$ distance results to the Softmax activation function, which is the most widely used distance measure in attention mechanism based models.
-In order to have softmax activation, we need the distance measure function $\phi$ in Definition 1.1 to be the softmax function in Definition 4.1.
-In Section G.1, we show how to extend to the softmax distance function in Lemma G.6.
-In Section G.2, we show how to adjust our algorithms.
-In Section G.3, we extend our algorithm to be robust to adaptive query.
-In Section G.4, we give the proof of our main result Theorem 5.1.
-G.1 Exponential Inner Product
-In this section, we show how we obtain the softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from [6].
-Definition G.1 (Definition 3.1 in [6]).
+In Section I.1, we show how to extend to the Softmax distance function in Lemma I.6.
+In Section I.2, we show how to adjust our algorithms.
+In Section I.3, we extend our algorithm to be robust to adaptive query.
+In Section I.4, we give the proof of our main result Theorem 3.1.
+I.1 Exponential Inner Product
+In this section, we show how we obtain the Softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from AS [6].
+Definition I.1 (Definition 3.1 in AS [6]).
 Let $r\geq 1$ denote a positive integer.
@@
   • $|\widetilde{A}_{i,j}-A_{i,j}|\leq\epsilon\cdot A_{i,j}$ for all $(i,j)\in[n]^{2}$.
-Lemma G.2 (Lemma 3.4 in [6]).
+Lemma I.2 (Lemma 3.4 in AS [6]).
 Suppose $Q,K\in\mathbb{R}^{n\times d}$, with $\|Q\|_{\infty}\leq R$, and $\|K\|_{\infty}\leq R$.
@@
 $r\leq{2s+2d\choose 2s}$ (10)
-such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition G.1) of $A\in\mathbb{R}^{n\times n}$.
+such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition I.1) of $A\in\mathbb{R}^{n\times n}$.
 Furthermore, the matrices $U_{1}$ and $U_{2}$ defining $\widetilde{A}$ can be computed in $O(n\cdot r)$ time.
-Here we consider the vector version of Lemma G.2.
-Definition G.3.
+Here we consider the vector version of Lemma I.2.
+Definition I.3.
 We define $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$.
-Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in [6].
-Remark G.4.
+Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in AS [6].
+Remark I.4.
 We use $\Gamma_{R,s}$ to denote the value range of our polynomial kernel methods function, i.e., $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$.
@@
 We take the maximum overall $s$ order approximation terms to get the upper bound of our value range.
-Lemma G.5 (Polynomial approximation).
-For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in [6] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
+We use the polynomial approximation method, which has been applied to accelerate Transformer model extensively [6, 7, 8, 72, 69].
+Lemma I.5 (Polynomial approximation).
+For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in AS [6] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
 Then, for any $x,y\in[0,R]^{d}$, we have
@@
 Proof.
-Let $n=1$. The proof follows from directly applying Lemma G.2. ∎
-Using the results from [6] above, we can extend our results to softmax activation.
-Lemma G.6 (Weighted softmax approximation, formal version of Lemma 6.7).
+Let $n=1$. The proof follows from directly applying Lemma I.2. ∎
+Using the results from AS [6] above, we can extend our results to Softmax activation.
+Lemma I.6 (Weighted Softmax approximation, formal version of Lemma C.7).
 Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
@@
 Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma G.5.
+Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma I.5.
 Then we can approximate exponential inner product using polynomial kernel function:
@@
 Proof.
-From Lemma G.5, we can use polynomial kernel to approximate the softmax function:
+From Lemma I.5, we can use polynomial kernel to approximate the Softmax function:
 $w^{\top}\exp(Xy/d)=$ $~{}\sum_{i\in[n]}w_{i}P(x_{i})^{\top}P(y)+O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
-The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma G.5.
+The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma I.5.
 Then, we can show
@@
 where the first step follows from $\|x-y\|_{2}^{2}=\|x\|_{2}^{2}+\|y\|_{2}^{2}-2\langle x,y\rangle$, and the second step follows $\|x\|_{2}^{2}=\sum_{j=1}^{d}|x_{j}|^{2}$ for $x\in\mathbb{R}^{d}$. ∎
-G.2 Algorithm Modifications
-Based on Lemma G.6, we can now extend our DP algorithms to handle softmax activation.
+I.2 Algorithm Modifications
+Based on Lemma I.6, we can now extend our DP algorithms to handle Softmax activation.
 First, we need to construct $P(y)$ and $P(x_{i})$ for $i\in[n]$, each costing $O(r)$ time.
-Then, for the second term in Lemma G.6, i.e.
+Then, for the second term in Lemma I.6, i.e.
 $\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$, we don’t need to add DP noises in it; instead, we calculate this term exactly, preprocess it, and store the results in the algorithm.
@@
 $\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p},$
-we can compute $\ell_{2}^{2}$ norm easily (see details in Lemma C.2).
-We then show how to extend our one dimensional $\ell_{1}$ distance algorithm (Algorithm 6 and 7) to $\ell_{2}^{2}$ distance with minor modifications.
-Theorem G.7 (DPTreeDistance $\ell_{2}^{2}$ distance).
+we can compute $\ell_{2}^{2}$ norm easily (see details in Lemma E.2).
+We then show how to extend our one dimensional $\ell_{1}$ distance algorithm (Algorithm 5 and 6) to $\ell_{2}^{2}$ distance with minor modifications.
+Theorem I.7 (DPTreeDistance $\ell_{2}^{2}$ distance).
 With $\alpha$ scaled down by a factor of $2$ and all Query instead multiplied by $R^{2}/(1+\alpha/2)^{2j}$ in Lines 8 and 13 of Algorithm 6, i.e., from
@@
   • Lines 8 and 13: Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R^{2}}{(1+\alpha/2)^{2j}}$.
-The data structure DPTreeDistance (Algorithm 6,7) uses $O(n)$ spaces to solve weighted $\ell_{2}^{2}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
+The data structure DPTreeDistance (Algorithm 5,6) uses $O(n)$ spaces to solve weighted $\ell_{2}^{2}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
   • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
@@
     It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta)$-DP private of $\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}$,
+    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}$,
     • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}|\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}+O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$,
@@
 Proof.
-The proof is similar to that of Theorem D.6, except that now our additive error includes $R$ increased by a power of $2$, i.e., from $O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$ to $O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$.
+The proof is similar to that of Theorem F.6, except that now our additive error includes $R$ increased by a power of $2$, i.e., from $O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$ to $O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$.
 ∎
 Now we can give our result that can answer Softmax query.
-Theorem G.8 (Softmax query, formal version of Theorem 4.2).
+Theorem I.8 (Softmax query, formal version of Theorem 4.2).
 Let $R\geq 1$.
 Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}$ be defined in Definition G.3.
+Let $\Gamma_{R,s}$ be defined in Definition I.3.
 Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
@@
   • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 3) It takes $O(\alpha^{-1}r\log^{2}n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
+    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
     • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
@@
 Proof.
-The DP proof is the same as the proof of Lemma E.1.
+The DP proof is the same as the proof of Lemma G.1.
 We then show the time complexity.
-From Lemma G.6, we know that constructing $P(\cdot)$ requires $O(r)$ time.
+From Lemma I.6, we know that constructing $P(\cdot)$ requires $O(r)$ time.
 In the first for loop of Init, the dominating time consumption is $O(nr)$.
@@
 The space complexity is $O(nr)$, since storing the $n\times r$ matrix $P$ is the dominating factor.
-The proof of the error follows from the triangle inequality by combining the errors in Lemma G.6 and Theorem G.7.
-G.3 Adaptive Softmax
+The proof of the error follows from the triangle inequality by combining the errors in Lemma I.6 and Theorem I.7. ∎
+I.3 Adaptive Softmax
 In this section, we show how to make Algorithm 3 robust to adaptive query.
-We follow the same idea from Section F.
-We notice that, in the softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$-distance in Section F.
+We follow the same idea from Section H.
+We notice that, in the Softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$-distance in Section H.
 Therefore, we need to recalculate Lipschitz constant first.
-Lemma G.9 (Lipschitz of weighted softmax).
+Lemma I.9 (Lipschitz of weighted Softmax).
 If the following conditions hold:
@@
 $\leq$ $~{}nd^{-1/2}RR_{w}\exp(R^{2})\|y-\widetilde{y}\|_{1}$
-where the first step follows from definition of $Z(y),Z(\widetilde{y})$, the second step follows from triangular inequality, the third step follows from Fact A.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$.
+where the first step follows from definition of $Z(y),Z(\widetilde{y})$, the second step follows from triangular inequality, the third step follows from Fact B.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$.
 ∎
 Then we can show how to extend our algorithm to be robust to adaptive query.
-Lemma G.10 (Adaptive softmax, formal version of Lemma 6.8).
+Lemma I.10 (Adaptive Softmax, formal version of Lemma C.8).
 If the following conditions hold:
@@
   • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
-Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
+Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
 $|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
 Proof.
-The proof follows from the same idea as the proof of Lemma F.8, except that we use Theorem G.8 and the Lipschitz in Lemma G.9. ∎
-Theorem G.11 (Adaptive query softmax data structure, formal version of Theorem 4.4).
+The proof follows from the same idea as the proof of Lemma H.8, except that we use Theorem I.8 and the Lipschitz in Lemma I.9. ∎
+Theorem I.11 (Adaptive query Softmax data structure, formal version of Theorem 4.4).
 Let $R\geq 1$.
 Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}$ be defined in Definition G.3.
+Let $\Gamma_{R,s}$ be defined in Definition I.3.
 Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
@@
   • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 1) It takes $O(\alpha^{-1}ld\log^{2}n)$ time to output a number $z$ such that
-    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
+    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
     • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
@@
 Proof.
-We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma G.10.
+We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma I.10.
 We know the additive error introduced by adaptive query is $E_{a}:=O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0})$ and the relative error introduced by polynomial kernel approximation is $E_{p}:=w^{\top}\exp(Xy/d)\cdot\epsilon_{s}$.
@@
 ∎
-G.4 Proof of main result
+I.4 Proof of Main Result
 In this section, we give the proof of our main result of Theorem 3.1.
-Theorem G.12 (Cross-attention, formal version of Theorem 5.1).
-Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1.
+Theorem I.12 (Softmax cross-attention, formal version of Theorem 3.1).
+Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1.
 Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter.
-Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma 6.7).
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
+Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma C.7).
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
 Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$.
-Let $\phi$ be softmax function (Definition 4.1).
 There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
@@
 It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
-    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP,
-    • $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
-    • $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP,
+    • the process of output $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
+    • the process of output $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
     • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
@@
 Proof.
-From Section 3.4, we know that we can make matrix $AV$ in cross-attention computation DP.
+From Section 3, we know that we can ensure matrix $AV$ in cross-attention computation satisfies DP.
 Next, from Theorem 4.4, for $i\in[m],j\in[n],k\in[d]$, we have $(AV)_{i,k}$ is $(\epsilon,\delta+\delta^{\prime})$-DP and also robust to adaptive query.
 For $D$, we compute the exact true value.
-By the post-processing property of DP (Fact 3.9), our cross-attention is DP.
+By the post-processing property of DP (Fact B.5), our cross-attention process is DP.
 Let $(AV)_{i,k}$ be the true value and $\widetilde{(AV)}_{i,k}$ be the noisy value. Then, from Theorem 4.4, we have

### v3

@@
-Differential Privacy of Cross-Attention with Provable Guarantee
-Differential Privacy of Cross-Attention with Provable Guarantee
-Yingyu Liang yingyul@hku.hk.
-The University of Hong Kong.
-yliang@cs.wisc.edu.
-University of Wisconsin-Madison.
-Zhenmei Shi zhmeishi@cs.wisc.edu.
-University of Wisconsin-Madison.
-Zhao Song magic.linuxkde@gmail.com.
+Provable Differentially Private Computation of the Cross-Attention Mechanism
+Provable Differentially Private Computation of the Cross-Attention Mechanism
+Yekun Ke and Yingyu Liang and Zhenmei Shi and Zhao Song and Jiahao Zhang The University of Hong Kong.University of Wisconsin-Madison.
+magic.linuxkde@gmail.com.
 The Simons Institute for the Theory of Computing at the University of California, Berkeley.
-Yufa Zhou yufazhou@seas.upenn.edu.
-University of Pennsylvania.
-Cross-attention has become a fundamental module nowadays in many important artificial intelligence applications, e.g., retrieval-augmented generation (RAG), system prompt, guided stable diffusion, and many more.
-Ensuring cross-attention privacy is crucial and urgently needed because its key and value matrices may contain sensitive information about model providers and their users.
-In this work, we design a novel differential privacy (DP) data structure to address the privacy security of cross-attention with a theoretical guarantee.
-In detail, let $n$ be the input token length of system prompt/RAG data, $d$ be the feature dimension, $0<\alpha\leq 1$ be the relative error parameter, $R$ be the maximum value of the query and key matrices, $R_{w}$ be the maximum value of the value matrix, and $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods.
-Then, our data structure requires $\widetilde{O}(ndr^{2})$ memory consumption with $\widetilde{O}(nr^{2})$ initialization time complexity and $\widetilde{O}(\alpha^{-1}r^{2})$ query time complexity for a single token query.
-In addition, our data structure can guarantee that the process of answering user query satisfies $(\epsilon,\delta)$-DP with $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}R^{2s}R_{w}r^{2})$ additive error and $n^{-1}(\alpha+\epsilon_{s})$ relative error between our output and the true answer.
-Furthermore, our result is robust to adaptive queries in which users can intentionally attack the cross-attention system.
-To our knowledge, this is the first work to provide DP for cross-attention and is promising to inspire more privacy algorithm design in large generative models (LGMs).
+Cross-attention has emerged as a cornerstone module in modern artificial intelligence, underpinning critical applications such as retrieval-augmented generation (RAG), system prompting, and guided stable diffusion.
+However, this is a rising concern about securing the privacy of cross-attention, as the underlying key and value matrices frequently encode sensitive data or private user information.
+In this work, we introduce a novel data structure designed to enforce differential privacy (DP) for cross-attention mechanisms, accompanied by provable theoretical guarantees.
+Specifically, letting $n$ denote the input sequence length, $d$ the feature dimension, $R$ the maximum magnitude of query and key matrices, $R_{w}$ the maximum magnitude of the value matrix, and $r,s,\epsilon_{s}$ the parameters for polynomial kernel methods, our proposed structure achieves $\widetilde{O}(ndr^{2})$ space and initialization complexity, with a query time of $\widetilde{O}(dr^{2})$ per token.
+Moreover, we demonstrate that our mechanism satisfies $(\epsilon,\delta)$-DP, incurring an additive error of $\widetilde{O}((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}R^{2s}R_{w}r^{2})$ and a relative error of $2\epsilon_{s}/(1-\epsilon_{s})$ with respect to the ground truth.
+Crucially, our framework maintains robustness against adaptive queries, ensuring security even in adversarial settings.
+To the best of our knowledge, this constitutes the first approach providing provable differential privacy for cross-attention, establishing a foundation for future privacy-preserving algorithms in large generative models (LGMs).
 1 Introduction
-The development of Artificial Intelligence (AI) has four stages: (1) prediction AI, e.g., ResNet [54] in image classification; (2) generation AI, e.g., ChatGPT [1] in language generation; (3) autonomous agent AI, Voyager [105] autonomously plays Minecraft game [30]; (4) Artificial Generalization Intelligence (AGI).
-Humans have made rapid progress in generative AI, and we are excitingly heading to the third stage, the era of AI agent [75].
-One prevalent application of AI agents is customized large generative models (LGMs) agents [80], e.g., AgentGPT [33], SuperAGI [36], MetaGPT [53, 47], GPT Researcher [35] and many so on.
-In particular, recently, Apple Inc.
-introduced Apple Intelligence [5], signaling the integration of LGMs into physical devices.
-This innovation allows devices to use personal information for real-life assistance, such as entering passport numbers when booking flights or informing users of their latest meetings.
-With increased AI capabilities, privacy concerns become significant, as the more personal information devices handle, the greater the potential privacy risks.
-One fundamental technique used in LGMs is cross-attention [97], which is an essential module in retrieval-augmented generation (RAG) [68], system prompt, guided stable diffusion, and many so on.
-In RAG, to be more professional, the LGMs answer user input queries by using a domain-specific database under cross-attention, which may contain specific privacy data and knowledge so that the LGMs gain additional power.
-For system prompts, based on cross-attention, some customized long prompts, e.g., user information or concrete rules, are concatenated before user input to follow human instructions better, which are commonly used in ChatGPT [34], Claude3 [4] and other commercial LGMs.
-Consequently, protecting the privacy of domain-specific data in RAG or system prompts is crucial as they contain sensitive information about users and companies.
+The proliferation of Large Generative Models (LGMs) is rapidly extending beyond general-purpose chatbots towards highly personalized AI agents [86, 63].
+This evolution is marked by the deep integration of LGMs into personal devices and enterprise systems, exemplified by recent initiatives like Apple Intelligence [3].
+Such agents are designed to leverage personal and proprietary information to provide customized assistance, such as accessing private documents, managing schedules, or utilizing confidential company data.
+This personalization is frequently enabled by techniques like Retrieval-Augmented Generation (RAG) [53] and the use of detailed system prompts.
+At the technical core of these capabilities lies the cross-attention mechanism [78].
+Cross-attention is the fundamental module that allows the model to weigh the importance of and selectively focus on information from these external sources.
+While powerful, this mechanism also introduces a significant privacy vulnerability.
+The context that provides the LGM with its enhanced capabilities, including sensitive user data, proprietary business logic, and confidential documents, is directly processed by the cross-attention layers.
+Given this, protecting the privacy of domain-specific data in RAG or system prompts becomes essential.
 These data and prompts are the core assets of many start-ups.
@@
 Our work demonstrates that the Softmax cross-attention computation is equivalent to computing the weighted distance problem.
-Definition 1.1 (Softmax cross-attention).
+Definition 1.1 (Softmax cross-attention, [78]).
 Let $n$ and $m$ be the token length of the data and input query, respectively.
 Let $d$ be the feature dimension.
-Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Softmax Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is
+Given fixed key matrix $K\in[0,R]^{n\times d}$ and fixed value matrix $V\in[-R_{w},R_{w}]^{n\times d}$, $R_{w}\geq 1$, for any input query matrix $Q\in[0,R]^{m\times d}$, the goal of the Softmax Cross-Attention Computation is to get the matrix $\mathrm{Attn}(Q,K,V)\in\mathbb{R}^{m\times d}$, which is
 $\mathrm{Attn}(Q,K,V):=D^{-1}AV,$
 where $A\in\mathbb{R}^{m\times n}$ satisfies $A_{i,j}:=\exp(\langle Q_{i},K_{j}\rangle/d)$ for any $i\in[m],j\in[n]$ ($Q_{i}$ and $K_{j}$ denote the $i$-th and $j$-th rows of $Q$ and $K$, respectively) and $D:=\mathrm{diag}(A{\bf 1}_{n})\in\mathbb{R}^{m\times m}$ is a diagonal matrix.
-Note that $\mathsf{Softmax}(QK^{\top})=D^{-1}A\in\mathbb{R}^{m\times n}$ in Definition 1.1, which is the standard function used in transformers, and usually, we call it as attention matrix.
+Note that $\mathsf{Softmax}(QK^{\top}/d)=D^{-1}A\in\mathbb{R}^{m\times n}$ in Definition 1.1, which is the standard function used in transformers, and usually, we call it as attention matrix.
 Our main theorem, presented below, provides a robust solution of cross-attention, ensuring privacy and accuracy guarantees.
@@
 Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1.
-Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter.
-Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma C.7).
-Let $l=\widetilde{O}(r)$.
-Then, our Algorithm 1 requires $O(lndr)$ memory with $O(lnr)$ initialization time and $\widetilde{O}(\alpha^{-1}lr)$ query time, such that with probability $1-p_{f}$, the output process of cross-attention satisfies $(\epsilon,\delta)$-DP and is robust to adaptive query with relative error $n^{-1}(\alpha+\epsilon_{s})$ and additive error $\widetilde{O}(n^{-1}\epsilon^{-1}\alpha^{-1/2}lR^{2s}R_{w}r)$.
+Let $p_{f}$ be the probability of failure parameter.
+Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma G.6).
+Then, our Algorithm 1 requires $\widetilde{O}(ndr^{2})$ total memory with $\widetilde{O}(ndr^{2})$ total initialization time and $\widetilde{O}(dr^{2})$ query time for a single token query, such that with probability $1-p_{f}$, the output process of cross-attention satisfies $(\epsilon,\delta)$-DP and is robust to adaptive query with relative error $2\epsilon_{s}/(1-\epsilon_{s})$ and additive error $\widetilde{O}((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}R^{2s}R_{w}r^{2})$.
 Our main technique in Theorem 1.2 ensures that cross-attention is differentially private by using the polynomial kernel approximation method and transforming it into a weighted distance problem.
@@
 We build a data structure for weighted Softmax queries in Section 5.3, and we extend this data structure to handle adaptive queries using the $\epsilon_{0}$-net/metric entropy argument in Section 5.4.
-Furthermore, our error decreases as the input token length grows, diminishing both the relative and additive errors to zero.
+Furthermore, our additive error decreases as the input token length grows, diminishing to zero.
 Our contributions are as follows:
@@
 The strength of DP lies in its strong, unambiguous, and concrete definition of privacy, enabling algorithm designs with provable privacy and accuracy analysis.
-Therefore, we believe that the theoretical aspects of DP applications in LGMs remain a highly impactful direction, and we aim to pave the way for further exploration in this area.
-1.1 Related Work
+Therefore, we believe that the theoretical aspects of DP applications in LGMs remain a highly impactful direction, and we aim to pave the way for further exploration.
+2 Related Works
+Differential Privacy Guarantee Analysis.
+Ever since [16] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [19, 51, 95, 68, 91].
+DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [37, 20, 2, 47, 38, 27, 8, 11, 21, 14, 33, 61, 62, 42, 48, 23, 24, 46, 15].
+Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [17].
+For example, truncated Laplace mechanism [26] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions.
+Despite extensive research on differential privacy, existing studies on differential privacy have not yet proposed methods specifically designed to enhance the privacy of cross-attention mechanisms.
 Differential Privacy in Data Structure and Attention.
 Differential privacy (DP) is a flourishing and powerful technique that has enormous applications in the topic of private machine learning.
-In the era of Large Generative Models (LGMs), there are three primary approaches to ensuring privacy: (1) during the pre-training stage: to protect training data [2, 84], (2) during the adaptation stage: to protect target data [9, 88, 63, 113, 74, 92], (3) during the inference stage: to protect user/system prompts [26] and RAG data [68].
-To protect training data, DP-SGD [2] uses DP optimizer to ensure data privacy, severing as the traditional baseline method.
-Recently, numerous works have aimed to improve this method by integrating DP in both the pre-training and fine-tuning stages of LGMs [113, 74, 31, 9, 92, 76, 88, 116, 63].
-To protect user/system prompts, EW [26] provides a survey on both DP and non-DP methods.
-In the use of LGMs, prompting methods almost become a standard way for inference [91].
-Given the billions of prompt interactions daily, ensuring privacy is essential [77].
-We refer readers to Appendix A for more related works.
-Roadmap.
-In Section 2, we present the preliminary of differential privacy (DP) and cross-attention.
-In Section 3, we present the main result of our cross-attention theorem (Theorem 3.1).
-In Section 4, we outline the main results of our algorithms.
-In Section 5, we discuss DP-related topics and potential extensions.
-In Section 6, we conclude our paper.
+In the era of Large Generative Models (LGMs), there are three primary approaches to ensuring privacy: (1) during the pre-training stage: to protect training data [1, 68], (2) during the adaptation stage: to protect target data [7, 72, 49, 93, 58, 75, 39], (3) during the inference stage: to protect user/system prompts [60, 22], RAG data [53], and style images [67].
+To protect training data, DP-SGD [1] uses DP optimizer to ensure data privacy, serving as the traditional baseline method.
+Recently, numerous works have aimed to improve this method by integrating DP in both the pre-training and fine-tuning stages of LGMs [93, 58, 25, 7, 75, 59, 72, 49].
+However, DP-SGD confines differential privacy to the optimizer.
+In contrast, we propose a novel approach that integrates DP directly into the attention mechanism, supported by strong theoretical analysis and guarantees.
+Given the resource-intensive nature of training LGMs, our technique offers a practical alternative for models trained with standard SGD, which lack inherent privacy guarantees.
+In such cases, applying DP-SGD would require retraining the models, which is computationally expensive, whereas our method avoids this additional cost.
+Additionally, there are several recent efforts on differential privacy (DP) in attention mechanisms.
+For instance, [29] approximates the attention matrix with DP guarantees but targets general attention without provable bounds for weighted cross-attention Softmax.
+Similarly, [18] analyzes DP training challenges, such as attention distraction from noise, but focuses on full-model optimization rather than mechanism-specific privatization.
+Another relevant study, [10], explores DP pretraining for linear attention, differing from our polynomial kernel-based approach for full Softmax in cross-attention.
+Our framework is the first to provide provable DP guarantees tailored to cross-attention, offering high-probability accuracy and resilience to adaptive attacks.
+Cross-Attention in System Prompt, RAG, Stable Diffusion and More.
+Cross-attention [78], first introduced in language translation, is a widely used technique in many advanced AI systems.
+For example, Stable Diffusion [71, 57, 79, 84, 87] and SORA [64] employ cross-attention as a core module for a text-to-image conditional generation.
+This technique is also utilized by other multimodal models [56], including Imagen [74] and Diffusion Transformer [69].
+In the realm of text-to-image editing, [35] analyzes and controls the cross-attention module to enable editing without requiring additional training.
+Furthermore, [94] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas.
+In addition, Retrieval Augmented Generation (RAG) [53, 9, 30], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module.
+Cross-attention also has other applications.
+[65] demonstrates that the prompt-tuning [55] task can be formulated as cross-attention, while [12] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation.
+Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [85], such as spatial reasoning [82], mathematical reasoning [50], in-context learning ability [76], compositional ability [89], few-shot adaptation ability [73, 90], and so on.
+There are some other works that used cross attention in Hopfield Models [40, 81, 34, 88, 80, 31, 32].
 3 Preliminary
-In this section, we give the preliminary of differential privacy (DP) and cross-attention. In Section 2.1, we describe the notations. In Section 2.2, we give definitions related to DP.
-2.1 Notations
+In this section, we describe the notations we use and give a preliminary of differential privacy (DP) and cross-attention.
+Notations.
 We use $\Pr[]$ to denote the probability.
@@
 We use $\|x\|_{p}$ to denote the $\ell_{p}$ norm of a vector $x\in\mathbb{R}^{n}$, i.e., $\|x\|_{1}:=\sum_{i=1}^{n}|x_{i}|$, $\|x\|_{2}:=(\sum_{i=1}^{n}x_{i}^{2})^{1/2}$, and $\|x\|_{\infty}:=\max_{i\in[n]}|x_{i}|$.
-2.2 Differential Privacy Definitions
-In this section, we give several definitions related to differential privacy (DP). We refer the reader to DR [20] for more background and details on DP.
+We denote polynomial time complexity with respect to $n$ as $\operatorname{poly}(n)$.
+For a function $f$, we use $\widetilde{O}(f)$ to represent $f$ multiplied by a polylogarithmic factor, i.e., $f\cdot\operatorname{poly}(\log f)$.
+This notation, known as soft-$O$ or tilde notation, simplifies expressions by omitting logarithmic factors, focusing on the dominant term’s growth rate.
+Differential Privacy Definitions. To move on, we give several definitions related to differential privacy (DP). We refer the reader to [17] for more background and details on DP.
 Definition 3.1 (Neighboring dataset).
-We define the two neighboring datasets as $X,X^{\prime}\in\mathbb{R}^{n}$ such that $\|X-X^{\prime}\|_{1}\leq 1$, i.e., they differ on a single data point.
+Two datasets $X,X^{\prime}\in[0,R]^{n\times d}$ are neighboring if they differ in exactly one row, i.e., there exists $i\in[n]$ such that $X_{i,*}\neq X^{\prime}_{i,*}$ and $X_{j,*}=X^{\prime}_{j,*}$ for all $j\neq i$.
 Definition 3.2 (Sensitivity).
-The sensitivity of a function $f:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is defined by: $\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n},\|X-X^{\prime}\|_{1}=1}\|f(X)-f(X^{\prime})\|_{1}.$
+The sensitivity of a function $f:\mathbb{R}^{n\times d}\to\mathbb{R}^{n\times d^{\prime}}$ is: $\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n\times d}}\|f(X)-f(X^{\prime})\|_{1},$ where $X,X^{\prime}$ are neighboring datasets and $\|\cdot\|_{1}$ is the entry-wise $\ell_{1}$-norm.
 Definition 3.3 ($(\epsilon,\delta)$-DP).
-For $\epsilon>0,\delta\geq 0$, a randomized algorithm $\mathcal{A}$ is $(\epsilon,\delta)$-DP, if for all $\mathcal{S}\subseteq\mathrm{Range}({\mathcal{A}})$ and for all $X,X^{\prime}$ such that $\|X-X^{\prime}\|_{1}\leq 1$:
+For $\epsilon>0,\delta\geq 0$, a randomized algorithm $\mathcal{A}$ is $(\epsilon,\delta)$-DP, if for all $\mathcal{S}\subseteq\mathrm{Range}({\mathcal{A}})$ and for all neighboring datasets $X,X^{\prime}$:
 $\Pr[{\mathcal{A}}(X)\in\mathcal{S}]\leq\exp(\epsilon)\Pr[{\mathcal{A}}(X^{\prime})\in{\mathcal{S}}]+\delta.$
 When $\delta=0$, the algorithm is said to have pure differential privacy.
-We mainly use the truncated Laplace mechanism, which has the following definitions.
-Definition 2.4 (Truncated Laplace distribution).
-We use $\mathrm{TLap}(\Delta,\epsilon,\delta)$ to denote the Truncated Laplace distribution with pdf proportional to $\exp(-\epsilon|z|/\Delta)$ on the region $[-B,B]$, where $B=\frac{\Delta}{\epsilon}\cdot\log(1+\frac{\exp(\epsilon)-1}{2\delta})$.
-Fact 2.5 (Theorem 3 in GDGK [32]).
-Let $z$ denote a $\mathrm{TLap}(\Delta,\epsilon,\delta)$ random variable. Then we have $\operatorname*{{\mathbb{E}}}[z]=0$, and
-$\operatorname*{\mathrm{Var}}[z]=\frac{2\Delta^{2}}{\epsilon^{2}}(1-\delta\cdot\frac{\log^{2}(1+\frac{e^{\epsilon}-1}{2\delta})+2\log(1+\frac{e^{\epsilon}-1}{2\delta})}{e^{\epsilon}-1}).$
+To obtain the DP guarantee, we employ the truncated Laplace mechanism, which is a standard practice in differential privacy.
+Definition 3.4 (Truncated Laplace distribution, [26]).
+We use $\mathrm{TLap}(\Delta,\epsilon,\delta)$ to denote the Truncated Laplace distribution with pdf proportional to $\exp(-\epsilon|z|/\Delta)$ on the region [B,B][-B,B], where $B=\frac{\Delta}{\epsilon}\cdot\log(1+\frac{\exp(\epsilon)-1}{2\delta})$.
+Fact 3.5 (Theorem 3 in [26]).
+Let $z$ denote a $\mathrm{TLap}(\Delta,\epsilon,\delta)$ random variable. Then we have $\operatorname*{{\mathbb{E}}}[z]=0$ and
+$~\operatorname*{\mathrm{Var}}[z]$
+$=$ $~=\frac{2\Delta^{2}}{\epsilon^{2}}(1-\delta\cdot\frac{\log^{2}(1+\frac{e^{\epsilon}-1}{2\delta})+2\log(1+\frac{e^{\epsilon}-1}{2\delta})}{e^{\epsilon}-1}).$
 Furthermore, if $\delta=0$, we have $\operatorname*{\mathrm{Var}}[z]=2\Delta^{2}/\epsilon^{2}$, meaning truncated Laplacian mechanism will be reduced to the standard Laplacian mechanism.
-Lemma 2.6 (Laplace mechanism, [20, 32], see Lemma 2.2 in AIMN [3]).
+Lemma 3.6 (Laplace mechanism, [17, 26], see Lemma 2.2 in [2]).
 Given a numeric function $f$ that takes a dataset $X$ as the input, and has sensitivity $\Delta$, the mechanism that outputs $f(X)+z$ where $z\sim\mathrm{Lap}(\Delta/\epsilon)$ is $(\epsilon,0)$-DP.
@@
 Moreover, the truncated Laplace mechanism is always accuracy up to error $B$.
-Algorithm 1 Adaptive query data structure
-1:datastructure DPTreeSoftmaxAdaptive $\triangleright$ Theorem 4.4
-2:members
-3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$triangleright Algorithm 3
-4:end members
-[…]
-6:     $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
-7:     for $i=1\to l$ do
-8:         ${\mathcal{D}}_{i}$.Init$(X,n,w,\epsilon/l,\delta/l,\delta^{\prime}/l,c,\epsilon_{s})$
-9:     end for
-10:end procedure
-11:procedure DistanceQuery($y\in[0,R]^{d},\alpha\in(0,1)$)
-12:     $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
-13:     $r\leftarrow 0^{l}$
-14:     for $i=1\to l$ do
-15:         $r_{i}\leftarrow{\mathcal{D}}_{i}.\textsc{DistanceQuery}(y,\alpha)$
-16:     end for
-17:     return Median of $r$
-18:end procedure
-19:end datastructure
 4 Main Results: Cross-Attention
@@
 For a typical weighted distance problem, we define the following: Let $w\in\mathbb{R}^{n}$ be the weights, $X\in\mathbb{R}^{n\times d}$ be the data matrix, where $x_{i}^{\top}$ is the $i$-th row of $X$ for $i\in[n]$, and let $y\in\mathbb{R}^{d}$ be the query.
-Suppose we need to answer $\ell_{1}$-distance query.
+Suppose we need to answer $\ell_{1}$ distance query.
 We have
 $\sum_{i\in[n]}\underbrace{w_{i}}_{\mathrm{weight}}\|\underbrace{y}_{\mathrm{query}}-\underbrace{x_{i}}_{\mathrm{data}}\|_{1}.$
+Algorithm 1 DP cross-attention algorithm
+1:datastrucutre DPCrossAttention $\triangleright$ Theorem 4.1
+2:members
+3:   ${\mathcal{D}}_{0},{\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeSoftmaxAdaptive}$ $\triangleright$ Algorithm 7
+4:end members
+5:procedure Init($K\in[0,R]^{n\times d}$, $V\in[-R_{w},R_{w}]^{n\times d}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01)$) $\triangleright$ $n=|K|$
+6:  for $k=1\to d$ do
 […]
+8:  end for
+[…]
+10:end procedure
+11:procedure Query($Q_{i}\in[0,R]^{d}$)
+12:  $O\leftarrow 0^{d}$
+[…]
+14:  for $k=1\to d$ do
+[…]
+16:  end for
+17:  return $O$
+18:end procedure
+19:end datastrucutre
+Now, we introduce cross-attention.
+Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1.
+In a standard cross-attention process, $K$ and $V$ are accessible before inference, while the user input $Q$ becomes available only when the user provides it.
+Here, $K$ and $V$ represent values stored in memory or disks and are considered private assets protected within the model, whereas $Q$ is treated as public.
+For the cross-attention mechanism $\mathrm{Attn}$ (Definition 1.1), we aim to ensure that the matrix $AV$ satisfies DP guarantee.
+Let $A_{i,j}=\exp(\langle Q_{i},K_{j}\rangle/d)$ for $i\in[m],j\in[n]$.
+Let $V_{j,k}\in\mathbb{R}$ be the $(j,k)$-th entry of $V$, for $j\in[n],k\in[d]$.
+The $(i,k)$-th entry of $AV$ for each $i\in[m],k\in[d]$ is computed by
 $(AV)_{i,k}=$ $~\sum_{j=1}^{n}\underbrace{V_{j,k}}_{\mathrm{weight}}\exp(\langle\underbrace{Q_{i}}_{\mathrm{query}},\underbrace{K_{j}}_{\mathrm{data}}\rangle/d),$ (1)
@@
 Therefore, we need to initialize a total of $d$ different data structures, each with weights $V_{*,k}$ for $k\in[d]$.
+For computing $D$, we treat $V={\bf 1}_{n}$, which can be interpreted as a weighted Softmax problem with weight ${\bf 1}_{n}$.
 Here, we present our main result below.
-Theorem 3.1 (Softmax cross-attention, informal version of Theorem I.12).
+Theorem 4.1 (Softmax cross-attention, informal version of Theorem G.11).
 Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1.
-Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter.
-Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma C.7).
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
+Let $p_{f}$ be the probability of failure parameter.
+Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma G.6).
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
 Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$.
-There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnrd)$ spaces to ensure cross-attention satisfies DP and supports the following operations:
-  • We initialize $d$ data structures using Init$(K,n,V_{*,k},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$ (Algorithm 1), for $k\in[d]$.
-It takes $O(lnr)$ time to initialize one data structure.
-  • At query time, for user input $Q$, we process one token at a time by passing the $i$-th row of $Q$, denoted $Q_{i}\in\mathbb{R}^{d}$, to DistanceQuery$(Q_{i},\alpha\in(0,1))$ (Algorithm 1) for each $i\in[m]$.
-It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
+Assume the input context length $n$ is large enough.
+Then there is a data structure DPTreeCrossAttention (Algorithm 1) that uses $O(lnrd)$ spaces to ensure cross-attention DP and supports the following operations:
+[…]
+[…]
     • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP,
-    • the process of output $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
-    • the process of output $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • the process of output $z$ has relative error $2\epsilon_{s}/(1-\epsilon_{s})$ and additive error $O((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n),$
     • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
     • it is robust to adaptive query.
-In Theorem 3.1, we use our DPTreeSoftmaxAdaptive (Algorithm 1) and guarantee that, for each query token of cross-attention, the output process satisfies $(\epsilon,\delta)$-DP with $n^{-1}(\alpha+\epsilon_{s})$ relative error and $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query.
-More specifically, the algorithm creates $d$ DPTreeSoftmaxAdaptive data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time.
-Notably, our error is inversely proportional to $n$, meaning that as the input token length increases, both the relative and approximate errors approach zero.
-This is achieved by the normalizing matrix $D$ (Definition 1.1).
-We refer the reader to Section I for proof details.
+Remark 4.2.
+Notice in Theorem 4.1 that we ensure the process of computing each entry is $(\epsilon,\delta+\delta^{\prime})$-DP.
+To guarantee that the overall output vector of length $d$ is DP, we initialize each $\mathcal{D}_{i}$ for $i\in\{0,1,2,\dots,d\}$ with parameters scaled from $\epsilon/2,\delta/2,\delta^{\prime}/2$ to $\epsilon/(d+1),\delta/(d+1),\delta^{\prime}/(d+1)$.
+Then, by the basic composition property (Fact A.8), the output vector is $(\epsilon,\delta+\delta^{\prime})$-DP, with the additive error increasing by a factor of $\widetilde{O}(d)$.
+In Theorem 4.1, we use our DPTreeCrossAttention (Algorithm 1) and guarantee that, for each query token of cross-attention, the output process satisfies $(\epsilon,\delta+\delta^{\prime})$-DP with $2\epsilon_{s}/(1-\epsilon_{s})$ relative error and $O((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(ldr\log n)$ running time under adaptive query.
+More specifically, the algorithm creates $d+1$ DPTreeSoftmaxAdaptive (Algorithm 7) data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time.
+Notably, our additive error is inversely proportional to $n$, meaning that as the input token length increases, the additive error approaches zero.
+We refer the reader to Section G for proof details.
 Thus, our algorithm theoretically protects system prompts/RAG data in cross-attention as discussed in Section 1.
@@
 Algorithm 2 DPTree initialization and query
-1:datastructure DPTree $\triangleright$ Theorem C.1
+1:datastructure DPTree $\triangleright$ Theorem B.1
 2:members
-3:     $b:\mathbb{R}^{2n-1}$
-4:     $c:\mathbb{R}^{2n-1}$
-5:end members
-6:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{R},\epsilon\in(0,1),\delta\in(0,1)$) $\triangleright$ Lemma D.3, Lemma C.3
-7:     $b[n,2n-1]\leftarrow a$
-8:     for $i=n\to 2n-1$ do
-9:         $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
-10:     end for
-11:     for $i=(\log n)\to 1$ do
-12:         for $j=1\to 2^{i-1}$ do
-13:              $k\leftarrow 2^{i-1}+j-1$
-14:              $b[k]\leftarrow b[2k]+b[2k+1]$
-15:              $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
+3:   $c:\mathbb{R}^{2n-1}$
+4:end members
+[…]
+6:  $b[n,2n-1]\leftarrow a$
+7:  for $i=n\to 2n-1$ do
+8:   $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
+9:  end for
+10:  for $i=(\log n)\to 1$ do
+11:   for $j=1\to 2^{i-1}$ do
+12:     $k\leftarrow 2^{i-1}+j-1$
+13:     $b[k]\leftarrow b[2k]+b[2k+1]$
+14:     $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
+15:   end for
 16:  end for
-17:     end for
-18:end procedure
-19:procedure Query($x\in[n],y\in[n]$) $\triangleright$ Lemma D.4, D.5, D.6
-20:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $c$) by using at most $2\log n$ nodes on the path.
-Let Value be the above summation.
-21:     return Value
-22:end procedure
-23:procedure TrueQuery($x\in[n],y\in[n]$)
-24:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $b$) by using at most $2\log n$ nodes on the path, where the height of the tree is $\log n$, and we need left and right boundary points.
-Let Value be the above summation.
-25:     return Value
-26:end procedure
-27:end datastructure
+17:end procedure
+18:procedure Query($y\in[0,R]$)
+19:  $c_{\rm left},c_{\rm right}\leftarrow 0,0$
+20:  for $i=1\to\log n$ do
+21:   Let node $j\in[2^{i}]$ of layer $i$ denotes the integer such that $y\in[(j-1)R/2^{i},jR/2^{i})$
+22:   if $j$ is even then $\triangleright$ Node $j$ is the right child of its parent
+23:     $c_{\rm left}\leftarrow c_{\rm left}+c[2^{i}+j-2]$ $\triangleright$ Add the value of left sibling node
+[…]
+25:     $c_{\rm right}\leftarrow c_{\rm right}+c[2^{i}+j]$ $\triangleright$ Add the value of right sibling node
+26:   end if
+27:  end for
+28:  return $c_{\rm left},c_{\rm right}$
+29:end procedure
+30:end datastructure
 5 Key Data Structure: DPTree
-This section provides our key data structures: DPTree (Algorithm 2), DPTreeDistance (Algorithm 5 and 6), DPTreeHighDim (Algorithm 7), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 1).
-In Section 4.1, we provide our high-level proof insights.
-In Section 4.2, we give our basic building block algorithms DPTree, DPTreeDistance and DPTreeHighDim.
-In Section 4.3, we present our DPTreeSoftmax algorithm that solves the weighted Softmax problem.
-In Section 4.4, we present our DPTreeSoftmaxAdaptive algorithm that enables DPTreeSoftmax to handle adaptive query problem.
+This section provides our key data structures: DPTree (Algorithm 2), DPTreeDistance (Algorithm 4 and 5), DPTreeHighDim (Algorithm 6), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 7).
 5.1 Technique Overview
@@
 (1) is not a typical distance measure like $\ell_{1}$ or $\ell_{2}$, but by using polynomial kernel method techniques, we transform it into a distance measure.
-AS [6] states that the exponential inner product can be approximated by polynomial kernel function $P(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{r}$, i.e., $P(x)^{\top}P(y)\approx\exp(x^{\top}y/d)$ for two vector $x,y\in\mathbb{R}^{d}$, with a relative error.
+[4] states that the exponential inner product can be approximated by polynomial kernel function $P(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{r}$, i.e., $P(x)^{\top}P(y)\approx\exp(x^{\top}y/d)$ for two vector $x,y\in\mathbb{R}^{d}$, with a relative error.
 Then, by the Law of Cosines, we transform the inner product of polynomial kernel functions into a distance measure, i.e.,
-$2P(x)^{\top}P(y)=$ $~{}-\|P(x)-P(y)\|_{2}^{2}+\|P(x)\|^{2}_{2}+\|P(y)\|^{2}_{2}.$ (2)
-After transforming Eq. (1) into a distance measure, we design the DPTree series data structures to provide cross-attention DP guarantee.
-In summary, we first design the data structure DPTree (Algorithm 2) that builds a binary segment tree with truncated Laplace noise added in the leaf nodes to ensure DP guarantee.
-Then, based on this data structure, we design DPTreeDistance (Algorithm 5 and 6) to answer one dimensional weighted distance queries $\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|$, which utilizes DPTree to store and return noised weights $w_{i}$ multiplied with the approximated distances between the query $y$ and data $x_{i}$.
-We further decompose high dimensional $\ell_{p}^{p}$-distance problem into one dimensional $\ell_{1}$-distance problems using
+$~2P(x)^{\top}P(y)$
+$=$ $~-\|P(x)-P(y)\|_{2}^{2}+\|P(x)\|^{2}_{2}+\|P(y)\|^{2}_{2}.$ (2)
+After transforming Eq. (1) into a distance measure, we design the DPTree series data structures to provide a cross-attention DP guarantee.
+In summary, we first design the data structure DPTree (Algorithm 2) that builds a binary segment tree with truncated Laplace noise added in the nodes to ensure DP guarantee.
+Then, based on this data structure, we design DPTreeDistance (Algorithm 4 and 5) to answer one dimensional weighted $\ell_{p}^{p}$ distance queries $\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}$.
+We further decompose high dimensional $\ell_{p}^{p}$ distance problem into one dimensional $\ell_{p}^{p}$ distance problems using
 $\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=$ $~\sum_{k=1}^{d}\sum_{i=1}^{n}w_{i}\cdot|y_{k}-x_{i,k}|^{p}.$ (3)
-Based on this decomposition, we design DPTreeHighDim (Algorithm 7) which is capable of answering high dimension queries.
+Based on this decomposition, we design DPTreeHighDim (Algorithm 6), which is capable of answering high-dimension queries.
 Then, using Eq.
-(2) and DPTreeHighDim, we design DPTreeSoftmax (Algorithm 3) to answer Softmax queries.
-By building multiple copies of this data structure, we boost the success probability such that it can answer any query (including adaptive query) with an additive error, establishing the final data structure DPTreeSoftmaxAdaptive (Algorithm 1).
-See Section C for a more detailed outline of algorithms and proof techniques.
+(5.1) and DPTreeHighDim, we design DPTreeSoftmax (Algorithm 3) to answer Softmax queries.
+By building multiple copies of this data structure, we boost the success probability such that it can answer any query (including adaptive query) with an additive error, establishing the final data structure DPTreeCrossAttention (Algorithm 1).
 5.2 DPTree, DPTreeDistance, and DPTreeHighDim
+The unweighted distance query has been explored in prior works [36, 8, 45].
+Specifically, [36] leverages online learning techniques to approximate the sum of distances, while [8] introduces a DP data structure based on a node-contaminated balanced binary tree.
+Furthermore, [45] presents a new data representation in tree nodes, where each node stores the sum of distances from one point to multiple points.
+In contrast, we focus on the weighted distance query, generalizing their results.
 We design a basic data structure DPTree (Algorithm 2) that answers summation queries by a summation segment tree with truncated Laplace noise (Definition 3.4).
 The algorithm first builds a binary summation tree in an array and then adds truncated Laplace noises to each node.
-In query time, we first trace from bottom nodes to find their lowest common ancestor, then report the summation by using at most $2\log n$ nodes on the path (Algorithm 2).
-Based on the parallel composition rule of DP (Fact B.7), we find that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries (Lemma D.8).
-See more details in Section D.
-We then design DPTreeDistance, a one-dimensional weighted $\ell_{1}$ distance data structure detailed in Algorithm 5 and 6.
-Initialization involves rounding each data point to the nearest multiple of a small interval and aggregating their weights into an array (illustrated in Figure 1), which is then input into our DPTree.
-At query time, we retrieve aggregated weights within small intervals and multiply them by their distances to the query point.
-We introduce a relative error parameter $\alpha$ to reduce the number of iterations to $O(\log(n)/\alpha)$, improving efficiency.
-Guided by Eq.(3), we design DPTreeHighDim (Algorithm 7), which extends DPTreeDistance to higher dimension by constructing independent data structures for each coordinate.
-See details in Section F and G.
+During a query, the algorithm traverses each layer of the binary structure based on the input $y$, aggregating values from sibling nodes by accessing at most $O(\log n)$ nodes along the path.
+It then returns the accumulated left and right sums as the query result (Algorithm 2).
+See more details in Section B.
+We then design DPTreeDistance, a one-dimensional weighted $\ell_{p}^{p}$ distance data structure detailed in Algorithm 4 and 5.
+Initialization involves assigning each data point to the nearest bin and aggregating their weighted polynomial terms into multiple arrays (illustrated in Figure 1), which are then used to initialize several instances of our DPTree.
+At query time, the algorithm retrieves aggregated weights from each DPTree corresponding to the query point and combines them using binomial coefficients and distance powers to compute the one-dimensional weighted $\ell_{p}^{p}$ distance.
+Guided by Eq.
+(3), we design DPTreeHighDim (Algorithm 6), which extends DPTreeDistance to higher dimension by constructing independent data structures for each coordinate.
+See details in Section D and E.
 Algorithm 3 Softmax query
@@
 2:members
-3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$triangleright Algorithm 5, Theorem I.7
-4:     $P:[0,\Gamma_{R,s}]^{n\times r}$triangleright Definition I.3 for $\Gamma_{R,s}$, Eq. (9) for $s$, Eq. (10) for $r$
+3:   ${\mathcal{D}}_{0},{\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$ $\triangleright$ Algorithm 4, Theorem D.1
+4:   $P:[0,\Gamma_{R,s}]^{n\times r}$ $\triangleright$ Definition G.3 for $\Gamma_{R,s}$, Eq. (9) for $s$, Eq. (10) for $r$
 5:   $w:[-R_{w},R_{w}]^{n}$
@@
 7:end members
-8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma C.7
+8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma G.6
 9:  $\epsilon_{s},~w,~P,~P_{wx},~s_{w}\leftarrow\epsilon_{s},~w,~0^{n\times r},~0,~0$
 10:  for $j=1\to n$ do
-11:         Compute $P(x_{j})$triangleright Polynomial kernel function $P(\cdot)$, Lemma I.5
-12:         Compute $w_{j}\|P(x_{j})\|_{2}^{2}$
-13:         $P_{wx}\leftarrow P_{wx}+w_{j}\|P(x_{j})\|_{2}^{2}$
-14:         $s_{w}\leftarrow s_{w}+w_{j}$
-15:         $P_{j,:}\leftarrow P(x_{j})$
-16:     end for
-17:     for $i=1\to r$ do
-18:         $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{r\log(1/\delta^{\prime})}$, $\delta/r$) $\triangleright$ Algorithm 5
-19:     end for
+11:   Compute $P(x_{j})$ $\triangleright$ Polynomial kernel function $P(\cdot)$, Lemma G.5
+12:   $P_{j,:}\leftarrow P(x_{j})$
+13:  end for
+14:  for $i=1\to r$ do
+15:   $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $\frac{c\epsilon}{3\sqrt{r\log(2/\delta^{\prime})}}$, $\frac{\delta}{3r}$) $\triangleright$ Algorithm 4
+16:   $P_{wx}\leftarrow P_{wx}+\mathcal{D}_{i}$.DistanceQuery(0)(0)
+17:  end for
+18:  $\mathcal{D}_{0}.$Init(${\bf 1}_{n}$, $n$, $w$, $\epsilon/3$, $\delta/3$)
+19:  $s_{w}\leftarrow s_{w}+\mathcal{D}_{0}$.DistanceQuery(0)(0)
 20:end procedure
-21:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma C.7
+21:procedure DistanceQuery($y\in[0,R]^{d}$) $\triangleright$ Lemma G.6
 22:  Value $\leftarrow 0$
@@
 25:  for $i=1\to r$ do
-26:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(P(y)_{i},\alpha)$triangleright Algorithm 6
+[…]
 27:  end for
@@
 $\sum_{i\in[n]}w_{i}\exp(\langle x_{i},y\rangle/d)=w^{\top}\exp(Xy/d).$
-Building on Definition 4.1, we develop a novel algorithm to answer differentially private weighted Softmax queries using the polynomial kernel method from AS [6].
-Specifically, in Eq.(2), there is a term that computes the weighted $\ell_{2}^{2}$ distance, which we calculate using DPTreeHighDim.
-We then compute the exact term for the weighted $\ell_{2}^{2}$ norms of the approximation kernel.
+Building on Definition 5.1, we develop a novel algorithm to answer differentially private weighted Softmax queries using the polynomial kernel method from [4].
+Specifically, in Eq.(5.1), the three terms compute the weighted $\ell_{2}^{2}$ distance, which we calculate using DPTreeHighDim.
 By summing these terms with a controlled error, we extend DPTreeHighDim to answer the Softmax query efficiently.
-More details can be found in Section I.
-Theorem 4.2 (Softmax query, informal version of Theorem I.8).
+More details can be found in Section G.
+Theorem 5.2 (Softmax query, informal version of Theorem G.7).
 Let $R\geq 1$.
-Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
 Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
-Our data structure DPTreeSoftmax (Algorithm 3) uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support following operations:
-  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1))$.
-(Algorithm 3) It takes $O(nr)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 3) It takes $O(\alpha^{-1}r\log^{2}n)$ time to output a number $z$ such that
+Let $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$.
+Let $r={2s+2d\choose 2s}$.
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
+Our data structure DPTreeSoftmax (Algorithm 3) uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
+[…]
+[…]
     • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
-    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
-      $+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|+O(\epsilon^{-1}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
     • it holds with probability at least $0.99$.
-Remark 4.3.
-In Theorem 4.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section C.5.
+In Theorem 5.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section G.
 Besides, note that the error bound in Theorem 5.2 does not depend on $\delta$ but depends on $\delta^{\prime}$.
-The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ reduces the error by a small constant (Fact 2.5).
-In practice, we set $\delta$ as a small positive constant close to 00.
-Please refer to the Lemma D.6 for more details.
+The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ reduces the error by a small constant (Fact 3.5).
+In practice, we set $\delta$ as a small positive constant close to 0.
+Please refer to the Lemma B.7 for more details.
 5.4 Adaptive Query Data Structure
@@
 By proving it can handle any query within the query space with a certain error, we ensure it effectively processes adaptive queries.
-We first boost the constant probability to high probability using the Chernoff bound (Lemma B.2).
+We first boost the constant probability to high probability using the Chernoff bound (Lemma A.2).
 Employing an $\epsilon_{0}$-net argument and the union bound, we bound all query points within the net.
 Finally, we use the Lipschitz property of the weighted Softmax distance function with an additive error to bound all points in the query space.
-The corresponding proofs can be found in Section H and Section I.
-Theorem 4.4 (Adaptive query Softmax data structure, informal version of Theorem I.11).
+The corresponding proofs can be found in Section F and Section G.
+Theorem 5.3 (Adaptive query Softmax data structure, informal version of Theorem G.10).
 Let $R\geq 1$.
-Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
 Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
-Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the failure probability parameter.
-Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$.
+Let $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$.
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
+Let $r={2s+2d\choose 2s}$.
+Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, and $p_{f}$ be the failure probability parameter.
+Let $l=O(r\log(dR/(\epsilon_{s}p_{f}))).$.
 There is a data structure DPTreeSoftmaxAdaptive (Algorithm 7) that uses $O(lnr)$ spaces to solve the weighted Softmax query problem for the dataset $X\subset[0,R]^{d}$ and supports the following operations:
-  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$.
-It takes $O(lnr)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output a number $z$ such that
+[…]
+[…]
     • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
-    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
-      $+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • we have $|z-w^{\top}\exp(Xy/d)|\leq|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|+O(\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
     • it holds with probability at least $1-p_{f}$ (where $p_{f}$ is used in $l$),
     • it is robust to adaptive query.
-Remark 4.5.
 We describe the parallelization of our algorithms.
@@
 In addition, the for loops in Algorithm 7 are also parallelizable since the $l=O(r\log(dR/(\epsilon_{s}p_{f})))$ copies are independent.
-After parallelization, we have the final time complexity of Init to be $O(nr)$ and DistanceQuery to be $O(\alpha^{-1}\log^{2}n)$ in Algorithm 1 with $O(lr)$ GPU process.
+After parallelization, we have the final time complexity of Init to be $O(nr)$ and DistanceQuery to be $O(\log n)$ in Algorithm 7 with $O(lr)$ GPU process.
 6 Discussion
-How do we extend to self-attention?
+How do we extend to self-attention and other data structures?
 As self-attention is a more fundamental module in LGMs, we would like to extend our data structure to this setting.
@@
 How we can solve this challenge is crucial, and we leave it as our future direction.
+Moreover, we observe that [52] introduces the DP matrix mechanism, which offers an alternative to our currently used binary tree data structure.
+A preliminary idea for extending this is as follows: consider $A=\exp(QK^{\top}/d)$ as defined in Definition 1.1, where $Q$ of size $m\times d$ represents the query matrix with $m$ linear queries, and $K$ serves as the database.
+Leveraging the results from [52], we could design an alternative algorithm to enhance the current binary tree data structure, DPTree.
+We leave this exploration for future work.
 Why not add noise to some other places?
@@
 In this paper, we consider the problem of $\sum_{i=1}^{n}w_{i}\exp(\langle x_{i},y\rangle/d)$ where $y,x_{i}\in[0,R]^{d}$ and $w\in[-R_{w},R_{w}]^{n}$ (Definition 5.1).
-Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorihtm 2).
-From Lemma C.3 and the way we initialize DPTree in Algorithm 5, we see that the sensitivity $\Delta$ of this problem is $2R_{w}$.
+Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorithm 2).
+From Lemma B.3 and the way we initialize DPTree in Algorithm 4, we see that the sensitivity $\Delta$ of this problem is $2R_{w}$.
 A simple method for adding noise involves adding $n$ noises to a length $n$ array, with each item $w_{i}\exp(\langle x_{i},y\rangle/d)$ for $i\in[n]$.
-However, this approach increases the error by a factor of $n$ by basic composition (Fact B.6) and also makes the model dependent on the number of queries.
+However, this approach increases the error by a factor of $n$ by basic composition (Fact A.8) and also makes the model dependent on the number of queries.
 Besides, it only supports a single query and requires rebuilding the tree for each new query, rendering it impractical.
-In contrast, our current noise-adding technique (Lines 9 and 15 of Algorithm 2) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$.
+In contrast, our current noise-adding technique (Lines 8 and 14 of Algorithm 2) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$.
 This method also supports multiple queries, eliminating the need to rebuild the tree each time.
-How to remove the relative error parameter $\alpha$?
-The relative error parameter $\alpha$ in Theorem 3.1 appears because of the $(1+\alpha)$-approximation introduced in Algorithm 5 (Remark F.3) to reduce the number of required iterations from naive $O(n)$ to $O(\log(n)/\alpha)$.
-However, we notice that a recent work [59] does not utilize $(1+\alpha)$-approximation and still achieves $O(\log n)$ iteration number.
-They introduce a new tree node representation where each node stores the sum of distances from one point to multiple points, enabling the answer to be divided into only $\log n$ values, each combining two distance values, two count values, and $y$ itself.
-Our DPTree algorithms can be integrated with their method, thus removing parameter $\alpha$.
 7 Conclusion
-To our knowledge, we are the first work to provide differential privacy for cross-attention.
+To the best of our knowledge, this paper represents the first attempt to incorporate differential privacy into the cross-attention mechanisms.
 This paper presents the DPTree data structures, which provide a differential privacy guarantee for the cross-attention module in large generative models.
@@
 Our results may inspire more privacy algorithm design in large generative models.
-Acknowledgement
-Research is partially supported by the National Science Foundation (NSF) Grants 2023239-DMS, CCF-2046710, and Air Force Grant FA9550-18-1-0166.
-Appendix
+Checklist
+  1. 1.
+    For all models and algorithms presented, check if you include:
+    1. (a)
+      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
+    2. (b)
+      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
+    3. (c)
+      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Not Applicable]
+  2. 2.
+    For any theoretical claim, check if you include:
+    1. (a)
+      Statements of the full set of assumptions of all theoretical results. [Yes]
+    2. (b)
+      Complete proofs of all theoretical results. [Yes]
+    3. (c)
+      Clear explanations of any assumptions. [Yes]
+  3. 3.
+    For all figures and tables that present empirical results, check if you include:
+    1. (a)
+      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Not Applicable]
+    2. (b)
+      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Not Applicable]
+    3. (c)
+      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Not Applicable]
+    4. (d)
+      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
+    1. (a)
+      Citations of the creator If your work uses existing assets. [Not Applicable]
+    2. (b)
+      The license information of the assets, if applicable. [Not Applicable]
+    3. (c)
+      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
+    4. (d)
+      Information about consent from data providers/curators. [Not Applicable]
+    5. (e)
+      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects, check if you include:
+    1. (a)
+      The full text of instructions given to participants and screenshots. [Not Applicable]
+    2. (b)
+      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
+    3. (c)
+      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
 Roadmap.
 The appendix is organized as follows.
-In Section A, we provide more related works.
-In Section B, we give the preliminary of our paper.
-In Section C, we offer an outline of our proof techniques.
-In Section D, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee.
-In Section E, we show how to solve weighted distance problem.
-In Section F, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{1}$ distance problem with DP and accuracy guarantee.
-In Section G, we present the analysis of our DPTreeHighDim (Algorithm 7) data structure, which can address the high-dimensional $\ell_{1}$ distance problem while ensuring differential privacy and accuracy guarantees.
-In Section H, we show how we can handle adaptive query.
-In Section I, we show how to extend our algorithm to Softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 1).
-Appendix A More Related Work
-Differential Privacy Guarantee Analysis.
-Ever since DMNS [19] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [22, 67, 115, 84, 111].
-DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [50, 24, 3, 61, 51, 37, 10, 12, 25, 16, 46, 78, 79, 56, 62, 28, 29, 60, 17, 13, 18, 27, 40, 39, 64, 38, 23, 95, 21, 107, 94, 42, 70, 66].
-Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [20].
-For example, truncated Laplace mechanism [32] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions.
-Cross-Attention in System Prompt, RAG, Stable Diffusion and More.
-Cross-attention [97], first introduced in language translation, is a widely used technique in many advanced AI systems.
-For example, Stable Diffusion [87, 73, 98, 103, 106] and SORA [81] employ cross-attention as a core module for a text-to-image conditional generation.
-This technique is also utilized by other multimodal models [72], including Imagen [90] and Diffusion Transformer [85].
-In the realm of text-to-image editing, HMT+ [49] analyzes and controls the cross-attention module to enable editing without requiring additional training.
-Furthermore, YYB+ [114] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas.
-In addition, Retrieval Augmented Generation (RAG) [68, 11, 43], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module.
-Cross-attention also has other applications.
-ORST [82] demonstrates that the prompt-tuning [71] task can be formulated as cross-attention, while CFP [14] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation.
-Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [104], such as spatial reasoning [101], mathematical reasoning [65], in-context learning ability [93], compositional ability [109], few-shot adaptation ability [89, 110], and so on.
-There are some other works that used cross attention in Hopfield Models [52, 100, 48, 108, 99, 44, 45].
-Appendix B More Preliminary
-In Section B.1, we give the probability tools we use in the paper.
-In Section B.2, we provide the algebraic facts we use.
-In Section B.3, we give the DP facts we use in the paper.
-In Section B.4, we compare between popular DP mechanisms.
-B.1 Probability Tools
+In Section A, we give the preliminary of our paper.
+In Section B, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee.
+In Section C, we show how to solve weighted distance problem.
+In Section D, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{p}^{p}$ distance problem with DP and accuracy guarantee.
+In Section E, we present the analysis of our DPTreeHighDim (Algorithm 6) data structure, which can address the high-dimensional $\ell_{p}^{p}$ distance problem while ensuring differential privacy and accuracy guarantees.
+In Section F, we show how we can handle adaptive query.
+In Section G, we show how to extend our algorithm to Softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 7).
+Appendix A More Preliminary
+In Section A.1, we give the probability tools we use in the paper.
+In Section A.2, we provide the algebraic facts we use.
+In Section A.3, we give the DP facts we use in the paper.
+In Section A.4, we compare between popular DP mechanisms.
+A.1 Probability Tools
 In this section, we give several probability lemmas.
-Lemma B.1 (Markov’s inequality).
+Lemma A.1 (Markov’s inequality).
 If $x$ is a nonnegative random variable and $t>0$, we have
 $\Pr[x\geq t]\leq\frac{\operatorname*{{\mathbb{E}}}[x]}{t}.$
-Lemma B.2 (Chernoff bound, [15]).
-Let $x_{i}$ be a Bernoulli random variable with probability $p_{i}$ of being equal to $1$ and $1-p_{i}$ of being equal to 00, and all $x_{i}$ for $i\in[n]$ are independent.
+Lemma A.2 (Chernoff bound, [13]).
+Let $x_{i}$ be a Bernoulli random variable with probability $p_{i}$ of being equal to $1$ and $1-p_{i}$ of being equal to 0, and all $x_{i}$ for $i\in[n]$ are independent.
 Let $x=\sum_{i=1}^{n}x_{i}$.
@@
 $\Pr[x\geq(1+\delta)\mu]\leq\exp(-\delta^{2}\mu/3),$
-and for all $10<\delta<$ 1
+and for all $0<\delta<1$
 $\Pr[x\leq(1-\delta)\mu]\leq\exp(-\delta^{2}\mu/2).$
-Lemma B.3 (Chebyshev’s inequality).
+Lemma A.3 (Chebyshev’s inequality).
 Let $x$ (integrable) be a random variable with finite non-zero variance $\sigma^{2}$ (and thus finite expected value $\mu$). Then for any real number $k>0$,
 $\Pr[|x-\mu|\geq k\sigma]\leq\frac{1}{k^{2}}.$
-B.2 Algebraic Facts
-Fact B.4 (Upper bound of exponential, Fact C.9 in LSSZ24b [73]).
+A.2 Algebraic Facts
+Fact A.4 (Upper bound of exponential, Fact C.9 in [57]).
 For $a\in\mathbb{R}$, $b\in\mathbb{R}$, $a,b\leq R$, where $R\geq 0$, we have
 $|\exp(a)-\exp(b)|\leq\exp(R)|a-b|.$
-B.3 DP Facts
-In this section, we present several facts about differential privacy (DP). We first state the post-processing property, which means, in an algorithm, if one step is DP, all the following steps are DP.
-Fact B.5 (Post-processing, see Fact 2.1 in GKK+ [37]).
+A.3 DP Facts
+In this section, we present several facts about differential privacy (DP).
+We first define vector neighboring dataset and sensitivity.
+Definition A.5 (Vector neighboring dataset).
+We define the two neighboring datasets as $X,X^{\prime}\in\mathbb{R}^{n}$ such that $\|X-X^{\prime}\|_{1}\leq 1$, i.e., they differ on a single data point.
+Definition A.6 (Vector sensitivity).
+The sensitivity of a function $f:\mathbb{R}^{n}\to\mathbb{R}^{d}$ is defined by: $\Delta:=\max_{X,X^{\prime}\in\mathbb{R}^{n},\|X-X^{\prime}\|_{1}=1}\|f(X)-f(X^{\prime})\|_{1}.$
+We state the post-processing property, which means, in an algorithm, if one step is DP, all the following steps are DP.
+Fact A.7 (Post-processing, see Fact 2.1 in [27]).
 Let ${\mathcal{A}}_{1}$ be an $(\epsilon,\delta)$-DP algorithm and ${\mathcal{A}}_{2}$ be a (randomized) post-processing algorithm.
@@
 If we have many DP algorithms, we need a composition rule. The most straightforward composition is the basic/sequential composition rule.
-Fact B.6 (Basic composition, see Fact 2.3 in GKK+ [37]).
+Fact A.8 (Basic composition, see Fact 2.3 in [27]).
 Let ${\mathcal{A}}_{1}$ be an $(\epsilon_{1},\delta_{1})$-DP algorithm and ${\mathcal{A}}_{2}$ be an $(\epsilon_{2},\delta_{2})$-DP algorithm.
@@
 We can do much better if we know that the inputs are disjoint.
-Fact B.7 (Parallel composition, see Fact 2.4 in GKK+ [37]).
+Fact A.9 (Parallel composition, see Fact 2.4 in [27]).
 Let ${\mathcal{A}}_{1}$ be an $(\epsilon_{1},\delta_{1})$-DP algorithm and ${\mathcal{A}}_{2}$ be an $(\epsilon_{2},\delta_{2})$-DP algorithm.
@@
 In addition, we have the advanced composition, which improves the dependence of the number of DP algorithms to square root but compromises the term $\delta^{\prime}$.
-Theorem B.8 (Advanced composition, see Theorem 3.20 in DR [20]).
+Theorem A.10 (Advanced composition, see Theorem 3.20 in [17]).
 For all $\epsilon,\delta,\delta^{\prime}\geq 0$, the class of $(\epsilon,\delta)$-differentially private mechanisms satisfies $(\epsilon^{\prime},k\delta+\delta^{\prime})$-differential privacy under $k$-fold adaptive composition for:
 $\epsilon^{\prime}=k\epsilon(e^{\epsilon}-1)+\epsilon\sqrt{2k\log(1/\delta^{\prime})}.$
-B.4 Comparison of Truncated Laplace, Gaussian, and Laplace Mechanisms
+A.4 Comparison of Truncated Laplace, Gaussian, and Laplace Mechanisms
 We first define the Laplace mechanism as below:
-Definition B.9 (Laplace distribution).
+Definition A.11 (Laplace distribution).
 We use $\mathrm{Lap}(b)$ to denote the pdf: $p(z)=\frac{1}{2b}\exp(-\frac{|z|}{b}).$
-Fact B.10.
+Fact A.12.
 For $z\sim\mathrm{Lap}(b)$, $\operatorname*{{\mathbb{E}}}[z]=0$, and $\operatorname*{\mathrm{Var}}[z]=2b^{2}$.
 Furthermore, if $b=\Delta/\epsilon$, we have $\operatorname*{\mathrm{Var}}[z]=2\Delta^{2}/\epsilon^{2}$.
-In this paper, we use the Chebyshev inequality to bound the error, and from GDGK [32], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
-The variance of Gaussian mechanism in Theorem 3.22 in DR [20]:
+In this paper, we use the Chebyshev inequality to bound the error, and from [26], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
+The variance of Gaussian mechanism in Theorem 3.22 in [17]:
 $\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}\log(1.25/\delta)}{\epsilon^{2}}.$
-The variance of Laplace mechanism in Fact B.10:
+The variance of Laplace mechanism in Fact A.12:
 $\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}}{\epsilon^{2}}.$
-The variance of truncated Laplace mechanism in Fact 2.5, for $c\in(0,1]$:
+The variance of truncated Laplace mechanism in Fact 3.5, for $c\in(0,1]$:
 $\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}c}{\epsilon^{2}}.$
 Thus, since it has the minimum variance, we choose the truncated Laplace mechanism to design our algorithms among these popular mechanisms.
-Appendix C Proof Outline
-This section provides the proof outline of our paper.
-In Section C.1, we analyze our DPTree data structure.
-In Section C.2, we show the sensitivity of summation problem.
-In Section C.3, we explain the high-level idea behind the weighted $\ell_{p}^{p}$ distance query.
-In Section C.4, we show how to answer one-dimensional weighted $\ell_{1}$ distance query.
-In Section C.5, we show how to answer Softmax distance query using previous algorithms.
-In Section C.6, we show how to handle adaptive query.
-By combining the results from these sections, we prove the main results in Section 4.
-C.1 Summation Segment Tree
-First, in order to solve the weighted distance problem, we need to have a basic DP algorithm (Algorithm 2) that can answer simple summation queries.
-After analyzing its DP and error in Section D, we state the data structure theorem.
-Theorem C.1 (DPTree data structure, informal version of Theorem D.1).
+Appendix B DPTree Algorithm
+In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 2).
+B.1 Single Data Structure
+We give the theorem of our DPTree data structure that can answer the summation problem with DP, accuracy, runtime guarantee.
+Theorem B.1 (DPTree data structure ).
 There is a data structure (see DPTree in Algorithm 2) that uses $O(n)$ spaces to support the following operations:
-  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
-  • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
-    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
-    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
+[…]
+[…]
+    • the process satisfies $(\epsilon,\delta)$-DP,
+    • $|z_{1}-\sum_{\{k|x_{k}\leq y\}}a_{k}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$ and $|z_{2}-\sum_{\{k|x_{k}\geq y\}}a_{k}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
     • it holds with probability $0.99$.
-During the design of the data structure, we found an interesting property based on the parallel composition rule of DP Fact B.7. We will now state the lemma, whose proof is provided in Section D.
-Lemma C.2 (Weighted sum of disjoint interval queries, informal version of Lemma D.8).
-If the following conditions hold that:
-  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
-  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
-  • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
-Then, we have Alg. 2 is $(\epsilon,\delta)$-DP and output $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ with the error upper bounded by
-$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
-with probability $0.99$.
-From Lemma C.2, we can see that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries, as long as the sum of squared weights is finite.
-C.2 Sensitivity for Range Summation Problem
-Our DP summation tree data structure DPTree (Algorithm 2) requires sensitivity parameter $\Delta$.
-In this section, we show that for the summation problem, we have the sensitivity $\Delta=2R$ if the input $X\in[-R,R]^{n}$.
-Lemma C.3 (Sensitivity of summation).
-Let $X\in[-R,R]^{n}$. We have the sensitivity $\Delta=2R$ for DPTree.Init in Algorithm 2.
 Proof.
-Let’s say two neighboring datasets $X$ and $X^{\prime}$ differ in $x_{i}$ and $x^{\prime}_{i}$ for some $i$ in the array $X$. Then for a summation problem, i.e. $f(X):=\sum_{i=1}^{n}x_{i}$, we have
-$\Delta=$ $~{}\max_{X,X^{\prime}}|f(X)-f(X^{\prime})|=\max_{X,X^{\prime}}|x_{i}-x^{\prime}_{i}|=2R.$
-where the first step follows from Definition 2.2, the second step follows from $X,X^{\prime}$ differ in $x_{i},x^{\prime}_{i}$, and the last step follows from each coordinate of the dataset is bounded in $[-R,R]$.
-∎
-C.3 Weighted $\ell_{p}^{p}$ Distance Problem
-In this section, we introduce the intuition behind the method for handling the weighted $\ell_{p}^{p}$ distance problem. The formal lemmas and proofs can be found in Section E.
-Given a dataset and a query point in $d$ dimensions, we round each coordinate of the data points and the query point to the nearest multiple of a small interval.
-We then aggregate the weights of data points that have been rounded to the same position.
-Finally, we compute the sum of these aggregated weights multiplied by the distances between the query point and the data points over the rounded positions.
-This approach makes the computation more efficient while maintaining sufficient accuracy.
-We provide an example of weighted $\ell_{1}$-distance of a one-dimensional dataset consisting of 10 data points, i.e., $X\in[0,1]^{10}$ and a query $y=0$ in Figure 1.
-Lemma C.4 (Weighted $\ell_{p}^{p}$-distance high dimension, informal version of Lemma E.2).
-If the following conditions hold:
-  • Let data $X\in[0,R]^{n\times d}$ and $x_{i}^{\top}\in[0,R]^{d}$ be the $i$-th row of $x$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]^{d}$.
-  • We round each dimension of $X$ and $y$ to an integer multiple of $R/n$.
-  • Let $x_{i,k},y_{k}$ denote the $k$-th coordinates of $x_{i},y$ for $k\in[d]$.
-  • Let $c_{j,k}:=\sum_{j_{0}\in S_{j,k}}w_{j_{0}}$ where the set $S_{j,k}$ is the set of index $i$ such that the corresponding $x_{i,k}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
-  • After rounding, we assume that $y_{k}$ is in the $l_{k}R/n$ position for $l_{k}\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
-For the weighted problem, we have
-$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}.$
-where $O(R/n)$ is the rounding error for each data point.
-Remark C.5.
-In Lemma C.4, we first round the dataset.
-This rounding simplifies the calculation by reducing the number of possible positions to consider, from real values in $[0,R]^{d}$ to the total $O(nd)$ spots.
-However, it also introduces an error $O(R/n)$ for one data point.
-Then, for one spot in the rounded dataset, we sum over the weights of that spot and multiply the corresponding distance raised to the power of $p$.
-Additionally, since we are dealing with $\ell_{p}^{p}$ distance, the rounding error is also raised to the power of $p$.
-Figure 1: The visualization of how to build $c_{j}$ of rounded dataset $X\in[0,1]^{10}$ and compute the weighted $\ell_{1}$ distance.
+The proofs follow from combining Lemma B.4 (running time of initialization), Lemma B.5 (running time of query), Lemma B.6 (DP of query), and Lemma B.7 (error of query) together. ∎
+Figure 1: The visualization of how to compute the weighted $\ell_{1}$ distance for rounded dataset $X\in[0,1]^{10}$.
 The number above each $x_{i}$ is $w_{i}$.
@@
 Then $\sum_{i=1}^{n}w_{i}|y-x_{i}|=0.1\cdot 2.2+0.3\cdot 3.1+0.3\cdot(-2)+0.3\cdot(-3)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5+0.9\cdot(-1)+0.9\cdot 1=4.4$.
-And $\sum_{j=0}^{n}|k-j|c_{j}/n=0.1\cdot 2.2+0.3\cdot(-1.9)+0.4\cdot 2+0.6\cdot 6+0.7\cdot 0.5=4.4$.
-See details in Lemma E.1.
-C.4 One-Dimensional Weighted $\ell_{1}$ Distance Data Structure
-Based on previous discussions in Section C.1 and C.3, we can now describe our one-dimensional weighted $\ell_{1}$ distance data structure, DPTreeDistance, presented in Algorithm 5 and 6, which generalizes the results from BLM+ [10].
-Drawing from the intuition in Section C.3, the initialization process is as follows: first, we round each data point in the dataset to the nearest multiple of a small interval and build an array that aggregates the corresponding weights.
-This array is then fed into our DPTree data structure in Algorithm 2.
-At query time, we query the DPTree to obtain the aggregated weights within a small interval and multiply these weights by the distance to the query point.
-Furthermore, we also introduce a relative error parameter $\alpha$ to reduce the total number of queries to $O(\log(n)/\alpha)$ instead of querying all $n$ positions.
-We also analyze the DP and the error bound; see details in Section F.
-Theorem C.6 (DPTreeDistance data structure, informal version of Theorem F.6).
-There is a data structure DPTreeDistance (Algorithm 5 and 6) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
-  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 6) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
-    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
-    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$,
-    • it holds with probability $0.99$.
-C.5 Softmax Activation
-We then describe how we extend the previous results to Softmax activation, i.e.
-exponential inner product function (Definition 4.1).
-From AS [6], we know that Softmax activation can be approximated by polynomial kernel function $P(\cdot)$ with a certain error.
-The following lemma shows that we can transform weighted Softmax queries into polynomial kernels.
-More specifically, we have one term that computes the weighted $\ell_{2}^{2}$ distance, which is the place where we add DP noises.
-Because of the decomposability of the $\ell_{p}^{p}$ distance, i.e.
-$\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p}$, we can easily extend the results of Section C.4 to handle the $\ell_{2}^{2}$ distance query.
-After that, we compute the term for the weighted $\ell_{2}^{2}$ norms of approximation kernel exactly.
-Summing all these terms, with a certain error, we can answer the Softmax query.
-Related proofs can be found in Section I.
-Lemma C.7 (Weighted Softmax approximation, informal version of Lemma I.6).
-Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
-Let $R\geq 1$.
-Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
-Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma I.5.
-Then, we can approximate the exponential inner product using the polynomial kernel function:
-$w^{\top}\exp(Xy/d)=$ $~{}-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
-$+~{}O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
-Moreover, the vectors $P(\cdot)$ can be computed in $O(r)$ time.
-C.6 Adaptive Query
-We introduce how we can modify our algorithm to solve the adaptive query problem using some tools in QRS+ [86].
-Our approach is based on proving that our algorithm can handle any query within the query space with a certain error.
-Since adaptive queries must lie within this space, our algorithm can effectively handle them.
-In Section C.5, we demonstrate our algorithm’s capability to answer weighted Softmax distance queries with constant probability.
-We then use the Chernoff bound to boost the constant probability of our algorithm to a high probability.
-Next, we apply the notion of an $\epsilon_{0}$-net to bound all query points within the net using the union bound.
-Finally, we bound all points in the query space by utilizing the Lipschitz property of the weighted Softmax distance function and introducing an additive error.
-See the proofs in Sections H and I.
-Lemma C.8 (Adaptive Softmax, informal version of Lemma I.10).
-If the following conditions hold:
-  • Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of ${\mathcal{B}}$, and let $|N|$ be the size of the net $N$.
-  • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
-  • Let the relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
-  • We create $l=O(\log((R/\epsilon_{0})^{r}/p_{f}))$ independent copies of the data structure $\{\textsc{DPTreeSoftmax}_{j}\}_{j=1}^{l}$ (Algorithm 3) and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
-  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
-  • Let $Z(y):=w^{\top}\exp(Xy/d)$.
-  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
-Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
-$|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
-Appendix D DPTree Algorithm
-In this section, we give the analysis of privacy, accuracy and runtime of our DPTree (Algorithm 2).
-In Section D.1, we give the theorem (Theorem D.1) of our data structure that can answer summation problem.
-In Section D.2, we improve our data structure from constant probability to high probability by applying Chernoff bound.
-In Section D.3, we give the analysis.
-In Section D.4, we show some results of our data structure if the input queries are disjoint.
-D.1 Single Data Structure
-We give the theorem of our DPTree data structure that can answer the summation problem with DP, accuracy, runtime guarantee.
-Theorem D.1 (DPTree data structure, formal version of Theorem C.1).
-There is a data structure (see DPTree in Algorithm 2) that uses $O(n)$ spaces to support the following operations:
-  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1))$. It takes $O(n)$ time to initialize the data structure.
-  • Query$(x\in[n],y\in[n])$. It takes $O(\log n)$ time to output a number $z$ such that
-    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
-    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}n)$,
-    • it holds with probability $0.99$.
-Proof.
-The proofs follow from combining Lemma D.3 (running time of initialization), Lemma D.4 (running time of query), Lemma D.5 (DP of query), and Lemma D.6 (error of query) together. ∎
-D.2 Boost the Constant Probability to High Probability
-We can use Chernoff bound to boost the high probability by repeating the data structure multiple times.
-Theorem D.2 (High-probability).
-There is a data structure (see DPTreeHighProb in Algorithm 4) that uses $O(n\log(1/\delta_{\mathrm{fail}}))$ spaces to support the following operations
-  • Init$(a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1),\delta_{\mathrm{fail}}\in(0,0.01))$.
-It takes $O(n\log(1/\delta_{\mathrm{fail}}))$ time to initialize the data structure.
-  • Query$(x\in[n],y\in[n])$. It takes $O(\log(n)\cdot\log(1/\delta_{\mathrm{fail}}))$ time to output a number $z$ such that
-    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i=x}^{y}a_{i}$,
-    • $|z-\sum_{i=x}^{y}a_{i}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$,
+See more details in Lemma C.1.
+B.2 Boost the Constant Probability to High Probability
+By applying the Chernoff bound, we can increase the probability of obtaining a correct result.
+This is achieved by replicating the data structure multiple times, generating several independent results, and then reporting the median of these results.
+Taking the median helps mitigate the effect of outliers and ensures that the final answer is reliable with high probability.
+Theorem B.2 (High-probability).
+There is a data structure that uses $O(n\log(1/\delta_{\mathrm{fail}}))$ spaces to support the following operations
+[…]
+[…]
+    • the process satisfies $(\epsilon,\delta)$-DP,
+    • $|z_{1}-\sum_{\{k|x_{k}\leq y\}}a_{k}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$ and $|z_{2}-\sum_{\{k|x_{k}\geq y\}}a_{k}|\leq O(\epsilon^{-1}\Delta\log^{3/2}(n)\cdot\log(1/\delta_{\mathrm{fail}}))$,
     • it holds with probability $1-\delta_{\mathrm{fail}}$ for failure probability $\delta_{\mathrm{fail}}\in(0,0.01)$.
 Proof.
-Note that our data structure (Theorem D.1) succeeds with probability $0.99$.
-The success of the algorithm (Theorem D.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma B.2).
+Note that our data structure (Theorem B.1) succeeds with probability $0.99$.
+The success of the algorithm (Theorem B.1) can be viewed as a Bernoulli random variable, to which we apply the Chernoff bound (Lemma A.2).
 By repeating the data structure $O(\log(1/\delta_{\mathrm{fail}}))$ times and taking the median of the outputs, we boost the success probability.
@@
 Note that $p=0.99$.
-By setting $\delta=0.1$ and using Chernoff bound from Lemma B.2, we can show
+By setting $\delta=0.1$ and using Chernoff bound from Lemma A.2, we can show
 $\Pr[z\leq l/2]\leq\Pr[z\leq(1-\delta)lp]\leq\exp(-\delta^{2}lp/2).$
@@
 We can make this hold by choosing $l=O(\log(1/\delta_{\mathrm{fail}}))$.
-By the DP basic composition rule (Fact B.6), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem D.1.
+By the DP basic composition rule (Fact A.8), we need to choose $\epsilon=\epsilon^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ and $\delta=\delta^{\prime}/O(\log(1/\delta_{\mathrm{fail}}))$ where $\epsilon^{\prime},\delta^{\prime}$ are the $\epsilon,\delta$ in Theorem B.1.
 ∎
-Algorithm 4 Boost constant probability
-1:datastructure DPTreeHighProb $\triangleright$ Theorem D.2
-2:members
-3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(\log(1/\delta_{\mathrm{fail}}))}:\textsc{DPTree}$triangleright Alg. 2
-4:end members
-5:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{N}_{+},\epsilon\in(0,1),\delta\in(0,1),\delta_{\mathrm{fail}}\in(0,0.01)$)
-6:     for $i=1\to O(\log(1/\delta_{\mathrm{fail}}))$ do
-7:         ${\mathcal{D}}_{i}$.Init$(a,n,\Delta,\epsilon/O(\log(1/\delta_{\mathrm{fail}})),\delta/O(\log(1/\delta_{\mathrm{fail}})))$
-8:     end for
-9:end procedure
-10:procedure Query($x\in[n],y\in[n]$)
-11:     $r\leftarrow 0^{O(\log(1/\delta_{\mathrm{fail}}))}$
-12:     for $i=1\to O(\log(1/\delta_{\mathrm{fail}}))$ do
-13:         $r_{i}\leftarrow{\mathcal{D}}_{i}.\textsc{Query}(x,y)$
-14:     end for
-15:     return Median of $r$
-16:end procedure
-17:end datastructure
-D.3 Algorithm of Data Structure
+B.3 Sensitivity for Summation Problem
+Our DP summation tree data structure DPTree (Algorithm 2) requires sensitivity parameter $\Delta$.
+In this section, we show that for the summation problem, we have the sensitivity $\Delta=2R$ if the input $X\in[-R,R]^{n}$.
+Lemma B.3 (Sensitivity of summation).
+Let $X\in[-R,R]^{n}$. We have the sensitivity $\Delta=2R$ for DPTree.Init in Algorithm 2.
+Proof.
+Let’s say two neighboring datasets $X$ and $X^{\prime}$ differ in $x_{i}$ and $x^{\prime}_{i}$ for some $i$ in the array $X$. Then for a summation problem, i.e. $f(X):=\sum_{i=1}^{n}x_{i}$, we have
+$\Delta=$ $~\max_{X,X^{\prime}}|f(X)-f(X^{\prime})|=\max_{X,X^{\prime}}|x_{i}-x^{\prime}_{i}|=2R.$
+where the first step follows from Definition A.6, the second step follows from $X,X^{\prime}$ differ in $x_{i},x^{\prime}_{i}$, and the last step follows from each coordinate of the dataset is bounded in [R,R][-R,R].
+∎
+B.4 Algorithm of Data Structure
 In this section, we analyze the accuracy, DP, and runtime of Algorithm 2.
 We first analyze the runtime.
-Lemma D.3 (Runtime of initialization, Algorithm 2).
+Lemma B.4 (Runtime of initialization, Algorithm 2).
 For the initialization, we have the time complexity of Algorithm 2 is $O(n)$.
@@
 All the computations are dominated by $O(n)$ time. ∎
-Lemma D.4 (Runtime of query, Algorithm 2).
+Lemma B.5 (Runtime of query, Algorithm 2).
 For each query, we have the time complexity of Algorithm 2 is $O(\log n)$.
@@
 We now analyze the DP.
-Lemma D.5 (Privacy of query, Algorithm 2).
+Lemma B.6 (Privacy of query, Algorithm 2).
 The output process of Query (see Algorithm 2) is $(\epsilon,\delta)$-DP.
@@
 There is no noise in the query stage.
-Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma C.3).
+Since the problem we care about is summation, if we change one leaf node, the sensitivity $\Delta=2R$ (see Lemma B.3).
 Since we add noise to each node in the tree, and each leaf node count will contribute to $\log n$ nodes, it is equivalent to our output function being in $\log n$ dimension.
 We will then blow up the DP parameter by $\log n$ factor.
-Thus, using the basic composition rule (Fact B.6), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP.
+Thus, using the basic composition rule (Fact A.8), the DP guarantee for the whole tree data structure is $((\epsilon/\log n)\cdot\log n,(\delta/\log n)\cdot\log n)$ which is $(\epsilon,\delta)$-DP.
 ∎
 We now analyze the accuracy.
-Lemma D.6 (Accuracy of query, Algorithm 2).
-Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 2.5, we have the error of Query(see Algorithm 2) output is upper bounded by:
+Lemma B.7 (Accuracy of query, Algorithm 2).
+Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Then, using Chebyshev’s inequality and Fact 3.5, we have the error of Query(see Algorithm 2) output is upper bounded by:
 $O(\epsilon^{-1}\Delta\log^{3/2}n).$
@@
 Proof.
-[…]
+Let $y\in[0,R]$ be the query.
+Let $A_{1},A_{2}=\textsc{Query}(y)$ denote the noised query answers returned by DPTree.Query in Algorithm 2.
+Let $A^{*}_{1},A^{*}_{2}$ be the true query answers without noise.
+Let $z:=A_{1}-A_{1}^{*}+A_{2}-A_{2}^{*}$, which from Algorithm 2 we can see this is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$.
+Thus,
 $z=\sum_{i=1}^{O(\log n)}z_{i}$
@@
 We know $\mu=\operatorname*{{\mathbb{E}}}[z]=0$ since $\operatorname*{{\mathbb{E}}}[z_{i}]=0$.
-From Fact 2.5, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
+From Fact 3.5, we know the variance for each $z_{i}$ is $\operatorname*{\mathrm{Var}}[z_{i}]=c\epsilon^{-2}\Delta^{2}\log^{2}n$ where $0<c\leq 2$ and $c=2$ when $\delta=0$.
 Therefore, we can show
@@
 where the first step follows from definition of $z$, the second step follows from every $z_{i}$ are independent to each other, and the last step follows from $\operatorname*{\mathrm{Var}}[z_{i}]=O(c\epsilon^{-2}\Delta^{2}\log^{2}n)$.
-Note that we wish to bound $|z|=|\textsc{Query}(S)-\textsc{TrueQuery}(S)|$ as our error.
-Using Lemma B.3, we can have
+Note that we wish to bound $|z|$ as our error.
+Using Lemma A.3, we can have
 $\Pr[|z|\geq k\sigma]\leq\frac{1}{k^{2}}.$
@@
 Thus, we conclude that error is bounded by $O(c^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)=O(\epsilon^{-1}\Delta\log^{3/2}n)$ (since $c\in(0,2]$) with probability $0.99$. ∎
-D.4 Disjoint Intervals
-In this section, we show some interesting results for our DPTree data structure if the input queries are disjoint.
-Lemma D.7 (Sum of disjoint interval queries).
-If the following conditions hold that:
-  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
-  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
-Then, we have Algorithm 2 is $(\epsilon,\delta)$-DP and outputs $\sum_{j=1}^{t}\textsc{Query}(S_{j})$ with the error upper bounded by
-$O(t^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
-with probability $0.99$.
-Proof.
-From Lemma D.5, we know that $\mathrm{DPTree}$.Query is $(\epsilon,\delta)$-DP.
-Then, from Fact B.7 and the disjoint intervals in Algorithm 6, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
-Let $\textsc{TrueQuery}(S_{j})$ denote the true interval query answer returned by DPTree.TrueQuery in Algorithm 2 for interval $S_{j}$.
-Let $\textsc{Query}(S_{j})$ denote the noised interval query answer returned by DPTree.Query in Algorithm 2 for interval $S_{j}$.
-Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}z_{j}$.
-From the proof of Lemma D.6, we know $z_{j}$ is the sum of $O(\log n)$ independent truncated Laplace random variables each with parameter $\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$ and the variance is bounded by
-$\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
-Since the intervals $S_{j}$ are disjoint, they are independent to each other. Then, we have
-$\operatorname*{\mathrm{Var}}[z]=$ $~{}\operatorname*{\mathrm{Var}}[\sum_{j=1}^{t}z_{j}]$
-$=$ $~{}\sum_{j=1}^{t}\operatorname*{\mathrm{Var}}[z_{j}]$
-$=$ $~{}O(t\epsilon^{-2}\Delta^{2}\log^{3}n)$
-where the first step follows from definition of $z$, the second step follows from the intervals are disjoint, and the last step follows from $\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{2}n)$.
-Note that we wish to bound $|z|$ as our error.
-Using Lemma B.3, we can have error bounded by
-$O(t^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
-with probability $0.99$. ∎
-Moreover, this can be generalized to weighted sum of queries.
-Lemma D.8 (Weighted sum of disjoint interval queries, formal version of Lemma C.2).
-If the following conditions hold that:
-  • Let there be $t$ disjoint intervals, i.e., $S_{j}$ for $j\in[t]$, such that $S_{j}\cap S_{k}=\emptyset$ for all $j\neq k$.
-  • Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$.
-  • Let $a_{j}$ for $j\in[t]$ be a series that square converges to $a$, i.e., $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
-Then, we have Alg. 2 is $(\epsilon,\delta)$-DP and output $\sum_{j=1}^{t}a_{j}\textsc{Query}(S_{j})$ with the error upper bounded by
-$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
-with probability $0.99$.
-Proof.
-The DP proof is the same as in the proof of Lemma D.7.
-Let $\textsc{TrueQuery}(S_{j})$ and $\textsc{Query}(S_{j})$ be same in the proof of Lemma D.7 Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$ and $z=\sum_{j=1}^{t}a_{j}z_{j}$.
-From the proof of Lemma D.7, we know the variance of $z_{j}$ is bounded by
-$\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
-Since the intervals $S_{j}$ are disjoint, they are independent to each other. Then, we have
-$\operatorname*{\mathrm{Var}}[z]=$ $~{}\operatorname*{\mathrm{Var}}[\sum_{j=1}^{t}a_{j}z_{j}]$
-$=$ $~{}\sum_{j=1}^{t}\operatorname*{\mathrm{Var}}[a_{j}z_{j}]$
-$=$ $~{}\sum_{j=1}^{t}a_{j}^{2}\operatorname*{\mathrm{Var}}[z_{j}]$
-$=$ $~{}\sum_{j=1}^{t}a_{j}^{2}\cdot O(\epsilon^{-2}\Delta^{2}\log^{3}n)$
-$=$ $~{}O(a\epsilon^{-2}\Delta^{2}\log^{3}n)$
-where the first step follows from the definition of $z$, the second step follows from the intervals are disjoint, the third step follows from the $\operatorname*{\mathrm{Var}}[az]=a^{2}\operatorname*{\mathrm{Var}}[z]$ for a random variable $z$ and a constant $a$, the fourth step follows from the $\operatorname*{\mathrm{Var}}[z_{j}]=O(\epsilon^{-2}\Delta^{2}\log^{2}n)$, and the last step follows from $\sum_{j=1}^{t}a_{j}^{2}\leq a$.
-Note that we wish to bound $|z|$ as our error.
-Using Lemma B.3, we can have error bounded by
-$O(a^{1/2}\epsilon^{-1}\Delta\log^{3/2}n)$
-with probability $0.99$. ∎
-Appendix E Weighted $\ell_{p}^{p}$ Distance
+Appendix C Weighted $\ell_{p}^{p}$ Distance
 In this section, we introduce how to handle weighted $\ell_{p}^{p}$ distance problem in the high level idea.
-In Section E.1, we show how to solve one dimensional weighted problem.
-In Section E.2, we show how to solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset.
-Suppose we have the original data $X\in[0,R]^{n}$ and weight $w\in\mathbb{R}^{n}$ and query $y\in[0,R]$. We want to compute the weighted $\ell_{1}$-distance, i.e.
-$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|.$
+We can solve high dimensional weighted problem by decomposing each coordinate of the high dimensional dataset.
+Thus, we only need to show how to solve the one-dimensional weighted problem.
 For data in $d$-dimension, due to the decomposability of $\ell_{p}^{p}$ distance, our problem will be: given $x_{i}\in[0,R]^{d}$ and $w_{i}\in\mathbb{R}$ for $i\in[n]$, and $y\in[0,R]^{d}$, we can compute
@@
 where $x_{i,j},y_{j}$ means the $j$-th coordinates of $x_{i},y$ for $j\in[d]$.
-Therefore, we can solve one dimension problem first, and then the high dimension case can be solved automatically.
-E.1 One Dimensional Weighted Distance
 Now we can give the lemma for weighted distance of dataset.
-Lemma E.1 (Weighted distance one dimension).
-If the following conditions hold:
-  • Let data $X\in[0,R]^{n}$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]$.
-  • We round $X$ and $y$ to an integer multiple of $R/n$.
-  • Let $c_{j}=\sum_{j_{0}\in S_{j}}w_{j_{0}}$ where set $S_{j}$ is the set of index $i$ such that the corresponding $x_{i}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$.
-  • After rounding, we assume $y$ is in the $kR/n$ position for $k\in\{0,1,2,\dots,n\}$.
-For the weighted problem, we have
-$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|=\sum_{j=0}^{n}(|k-j|R/n+O(R/n))c_{j}.$
-Moreover, we have
-$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(|k-j|R/n+O(R/n))^{p}c_{j}$
-where $O(R/n)$ is the rounding error for each data point.
+Lemma C.1 (Weighted distance one dimension).
+For a collection of numbers $\{x_{1},x_{2},\cdots,x_{n}\}\subset\mathbb{R}$ and corresponding weights $\{w_{1},w_{2},\cdots,w_{n}\}\subset\mathbb{R}$, and a number $y\in\mathbb{R}$. We define two sets
+$S_{+}:=$ $~\{k\in[n]~:~x_{k}>y\}$
+$S_{-}:=$ $~\{k\in[n]~:~x_{k}<y\},$
+It holds
+$\sum_{k=1}^{n}w_{k}|x_{k}-y|^{p}=\sum_{j=0}^{p}{p\choose j}y^{p-j}((-1)^{p-j}\sum_{k\in S_{+}}w_{k}x_{k}^{j}+(-1)^{j}\sum_{k\in S_{-}}w_{k}x_{k}^{j}),$
+where $p\choose j$} denotes the binomial coefficient that ${p\choose j}=\frac{p!}{j!(p-j)!}$.
 Proof.
-For each $i$, we have:
-$w_{i}\cdot|y-x_{i}|=w_{i}\cdot(\frac{|k-j|R}{n}+O(\frac{R}{n})).$
-where $O(R/n)$ is the rounding error introduced by each data point, since each data point will be at most $O(R/n)$ away from its true position.
-We can construct $c_{j}$ by
-$c_{j}=\sum_{j_{0}\in S_{j}}w_{j_{0}}$
-set $S_{j}$ is the set of index $i$ such that the corresponding $x_{i}$ is rounded to $jR/n$. Moreover, $c_{j}$ can be negative.
-Summing over all $i$ and grouping by $j$, we get:
-$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|=\sum_{j=0}^{n}(\frac{|k-j|R}{n}+O(\frac{R}{n}))c_{j}.$
-The total rounding error will be $O(R)$ because we have $n$ data points, each with an error of at most $O(R/n)$.
-Moreover, we have
-$\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}=\sum_{j=0}^{n}(\frac{|k-j|R}{n}+O(\frac{R}{n}))^{p}c_{j}.$
-E.2 High Dimensional Weighted Distance
-Finally, we can solve the problem of weighted distance for $d$-dimensional dataset.
-Lemma E.2 (Weighted $\ell_{p}^{p}$-distance high dimension, formal version of Lemma C.4).
-If the following conditions hold:
-  • Let data $X\in[0,R]^{n\times d}$ and $x_{i}^{\top}\in[0,R]^{d}$ be the $i$-th row of $x$, weight $w\in\mathbb{R}^{n}$, query $y\in[0,R]^{d}$.
-  • We round each dimension of $X$ and $y$ to an integer multiple of $R/n$.
-  • Let $x_{i,k},y_{k}$ denote the $k$-th coordinates of $x_{i},y$ for $k\in[d]$.
-  • Let $c_{j,k}:=\sum_{j_{0}\in S_{j,k}}w_{j_{0}}$ where set $S_{j,k}$ is the set of index $i$ such that the corresponding $x_{i,k}$ is rounded to $jR/n$ for $j\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
-  • After rounding, we assume $y_{k}$ is in the $l_{k}R/n$ position for $l_{k}\in\{0,1,2,\dots,n\}$ for $k\in[d]$.
-For the weighted problem, we have
-$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}$
-where $O(R/n)$ is the rounding error for each data point.
-Proof.
-We can show
-$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=$ $~{}\sum_{k=1}^{d}\sum_{i=1}^{n}w_{i}\cdot|y_{k}-x_{i,k}|^{p}$
-$=$ $~{}\sum_{k=1}^{d}\sum_{j=0}^{n}(|l_{k}-j|R/n+O(R/n))^{p}c_{j,k}$
-where the first step follows from decomposability of $\ell_{p}^{p}$-distance by dimension, the second step follows from Lemma E.1.
-Appendix F One-Dimensional Weighted $\ell_{1}$ Distance Query
-In this section, we generalize the algorithms in BLM+ [10] to weighted distance.
-Here, we compute the problem of one-dimensional weighted $\ell_{1}$ distance query i.e.
+We show that
+$\sum_{k=1}^{n}w_{k}|x_{k}-y|^{p}=$ $~\sum_{x_{k}\in S_{+}}w_{k}(x_{k}-y)^{p}+\sum_{x_{k}\in S_{-}}w_{k}(y-x_{k})^{p}$
+$=$ $~(\sum_{x_{k}\in S_{+}}w_{k}\sum_{j=0}^{p}(-1)^{p-j}{p\choose j}x_{k}^{j}y^{p-j})+(\sum_{x_{k}\in S_{-}}w_{k}\sum_{j=0}^{p}(-1)^{j}{p\choose j}x_{k}^{j}y^{p-j})$
+$=$ $~\sum_{j=0}^{p}({p\choose j}(-1)^{p-j}y^{p-j}\sum_{k\in S_{+}}w_{k}x_{k}^{j})+\sum_{j=0}^{p}({p\choose j}(-1)^{j}y^{p-j}\sum_{k\in S_{-}}w_{k}x_{k}^{j})$
+$=$ $~\sum_{j=0}^{p}{p\choose j}y^{p-j}((-1)^{p-j}\sum_{k\in S_{+}}w_{k}x_{k}^{j}+(-1)^{j}\sum_{k\in S_{-}}w_{k}x_{k}^{j}).$
+Thus, we complete the proof. ∎
+Appendix D One-Dimensional Weighted $\ell_{p}^{p}$ Distance Query
+In this section, we generalize the algorithms in [8] and [45] to weighted distance.
+Here, we compute the problem of one-dimensional weighted $\ell_{p}^{p}$ distance query i.e.
 $\sum_{i\in[n]}w_{i}|y-x_{i}|$ for a given query $y\in[0,R]$, weights $w\in[-R_{w},R_{w}]^{n}$ and dataset $X\subset[0,R]$ and $n=|X|$.
-In Section F.1, we analyze the runtime of our algorithm.
-In Section F.2, we analyze the DP and accuracy of our algorithm.
-In Section F.3, we give the theorem for our DPTreeDistance data structure.
+In this section, we give the theorem for our DPTreeDistance data structure.
 Algorithm 4 Pre-processing data structure
-1:datastructure DPTreeDistance $\triangleright$ Theorem F.6
+1:datastructure DPTreeDistance $\triangleright$ Theorem D.1
 2:members
-3:     ${\mathcal{D}}:$ DPTree $\triangleright$ Alg. 2
+3:   ${\mathcal{D}}_{0},\dots,{\mathcal{D}}_{p}:$ DPTree $\triangleright$ Alg. 2
 4:   $X:[0,R]^{n}$
@@
 6:end members
-7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma E.1
-8:     $X,w,a\leftarrow X,w,0^{n+1}$
-9:     for $i=1\to n$ do
-10:         $j\leftarrow\textsc{Round}(x_{i},n)$triangleright $x_{i}\in X$ for $i\in[n]$
-11:         $a_{j}\leftarrow a_{j}+w_{i}$
-12:     end for
-13:     ${\mathcal{D}}$.$\textsc{Init}(a,n+1,2R_{w},\epsilon,\delta)$triangleright Alg. 2, Lemma C.3
-14:end procedure
-15:procedure Round($x\in[0,R]$, $n\in\mathbb{N}_{+}$)
-16:     Let $j\in\{0,1,2,\dots n-1\}$ denote the integer such that $jR/n\leq x<(j+1)R/n$
-17:     if $|x-(j+1)R/n|\leq|x-jR/n|$ then
-18:         $j\leftarrow j+1$
-19:     end if
-20:     return $j$
-21:end procedure
-22:end datastructure
-Algorithm 6 One dimensional weighted $\ell_{1}$ distance query
-1:datastructure DPTreeDistance $\triangleright$ Theorem F.6
-2:procedure DistanceQuery($y\in[0,R]$, $\alpha\in(0,1)$) $\triangleright$ Lemma F.2, Lemma F.4, Lemma F.5
-3:     $y\leftarrow\textsc{Round}(y,n)\cdot(R/n)$triangleright Alg. 5
-4:     Value $\leftarrow 0$
-5:     for $j=0,1,...,O(\log(n)/\alpha)$ do
-6:         $l_{j}\leftarrow\textsc{Round}(y+\frac{R}{(1+\alpha)^{j+1}},n)$
-7:         $r_{j}\leftarrow\textsc{Round}(y+\frac{R}{(1+\alpha)^{j}},n)$triangleright Consider the points to the right of $y$
-8:         Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$triangleright Alg. 2
-9:     end for
-10:     for $j=0,1,...,O(\log(n)/\alpha)$ do
-11:         $l_{j}\leftarrow\textsc{Round}(y-\frac{R}{(1+\alpha)^{j}},n)$
-12:         $r_{j}\leftarrow\textsc{Round}(y-\frac{R}{(1+\alpha)^{j+1}},n)$triangleright Consider the points to the left of $y$
-13:         Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$triangleright Alg. 2
+7:procedure Init($X\subset[0,R]$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$ ) $\triangleright$ Lemma C.1
+8:  $X,w,a\leftarrow X,w,0^{n\times(p+1)}$
+9:  for $i=1\to n$ do $\triangleright$ $x_{i}\in X$ for $i\in[n]$
+10:   Let $j\in[n]$ denotes the integer such that $x_{i}\in[(j-1)R/n,jR/n)$
+11:   for $q=0\to p$ do
+12:     $a_{j,q}\leftarrow a_{j,q}+w_{i}x_{i}^{q}$
+13:   end for
 14:  end for
-15:     Return Value
-16:end procedure
-17:end datastructure
-F.1 Runtime Analysis
-We first analyze the runtime.
-Lemma F.1 (Runtime of initialization, Algorithm 5).
-For the initialization, we have the time complexity of Init in Algorithm 5 is $O(n)$.
-Proof.
-In the initialization of Init, the computations need $O(n)$ time to compute the count and $O(\log n)$ time to build the tree. Thus, total time is $O(n)$. ∎
-Lemma F.2 (Runtime of DistanceQuery, Algorithm 6).
-For the $\ell_{1}$ distance query, we have the time complexity of DistanceQuery in Algorithm 6 is $O(\alpha^{-1}\log^{2}n)$.
-Proof.
-In DistanceQuery, the computations need $O(\log n)$ time to compute one value from DPTree.Query and this process need to be repeated $O(\alpha^{-1}\log n)$ times. ∎
-Remark F.3.
-In Line 8 and 13 of Algorithm 6, we use $R/(1+\alpha)^{j}$ to approximate the distance of each data point to the query in Lemma E.1, i.e.
-$|k-j|R/n$.
-This will introduce $\alpha$ relative error but also reduce the numbers of iteration from $O(n)$ to $O(\log(n)/\alpha)$.
-F.2 Privacy and Accuracy Analysis
-We show the DP.
-Lemma F.4 (Privacy of DistanceQuery, Algorithm 6).
-The output process of DistanceQuery (Algorithm 6) is $(\epsilon,\delta)$-DP.
-Proof.
-From Lemma D.5, we know that $\mathrm{DPTree}$.Query is $(\epsilon,\delta)$-DP output.
-We observe that intervals in Algorithm 6 are disjoint.
-Then, following the same logic in the proof of Lemma D.8, we can conclude that the value returned is $(\epsilon,\delta)$-DP.
-∎
-We now analyze the accuracy of the algorithm.
-Lemma F.5 (Accuracy of DistanceQuery, Algorithm 6).
-If the following conditions are satisfied:
-  • Let $X\in[0,R]^{n}$ be a dataset consisting of $n$ one-dimensional numbers, with weights $w\in[-R_{w},R_{w}]^{n}$.
-  • Let $\alpha\in(0,1)$ represent the relative error parameter utilized in Algorithm 6.
-  • Let $\widetilde{A}$ denote the output of the DistanceQuery in Algorithm 6.
-  • Let $A_{*}:=\sum_{i\in[n]}w_{i}|y-x_{i}|$ represent the true distance query value for a specific query $y$.
-Then with probability $0.99$, we have
-$|\widetilde{A}-A_{*}|\leq\alpha A_{*}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n).$
-Proof.
+15:  for $q=0\to p$ do
+16:   ${\mathcal{D}}_{q}$$\textsc{Init}(a_{:,q},n,2R_{w}R^{q},\epsilon/(p+1),\delta/(p+1))$ $\triangleright$ Alg. 2, Lemma B.3
+17:  end for
+18:end procedure
+19:end datastructure
+Algorithm 5 One dimensional weighted $\ell_{p}^{p}$
+distance query
+1:datastructure DPTreeDistance $\triangleright$ Theorem D.1
+2:procedure DistanceQuery($y\in[0,R]$)
+3:  for $q=0\to p$ do
+4:   $c_{\mathrm{left},q},c_{\mathrm{right},q}\leftarrow$ $\mathcal{D}_{q}.$Query($y$)
+5:  end for
+6:  return $\sum_{q=0}^{p}{p\choose q}y^{p-q}((-1)^{p-q}c_{\mathrm{right},q}+(-1)^{q}c_{\mathrm{left},q})$
+7:end procedure
+8:end datastructure
+Theorem D.1 (DPTreeDistance data structure ).
+There is a data structure DPTreeDistance (Algorithm 4,5) that uses $O(np)$ spaces to solve weighted $\ell_{p}^{p}$ distance query problem for dataset $X\subset[0,R]$ and support the following operations:
 […]
-$\widehat{A}:=$ $~{}\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot\textsc{TrueQuery}(S_{j}).$
-Since TrueQuery returns the sum of the corresponding weights, it aligns with the true answer $A_{*}:=\sum_{i\in[n]}w_{i}|y-x_{i}|$. Thus, we have
-$|\widehat{A}-A_{*}|\leq\alpha\cdot A_{*},$
-because for all $j$, the distances between $y$ and different points in $X$ vary only by a multiplicative factor of $(1+\alpha)$.
-Next we show the additive error.
-Let $\textsc{Query}(S_{j})$ denote the noised interval query answer returned by DPTree.Query in Algorithm 2.
-Algorithm 6 outputs $\widetilde{A}=\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot\textsc{Query}(S_{j})$.
-We wish to bound
-$|\widehat{A}-\widetilde{A}|\leq$ $~{}|\sum_{j=0}^{O(\log(n)/\alpha)}\frac{R}{(1+\alpha)^{j}}\cdot(\textsc{TrueQuery}(S_{j})-\textsc{Query}(S_{j}))|.$
-Let $z_{j}:=\textsc{Query}(S_{j})-\textsc{TrueQuery}(S_{j})$, which from Algorithm 2 we can see this is the sum of $O(\log n)$ independent truncated Laplace random variables.
-From Lemma D.8, we only need to show that the series $\frac{1}{(1+\alpha)^{j}}$ for $j\in\{0,1,\dots,O(\log(n)/\alpha)\}$ square converges to $1/\alpha$, since $R$ is a constant.
-We can show
-$\sum_{j=0}^{O(\log(n)/\alpha)}\frac{1}{(1+\alpha)^{2j}}\leq$ $~{}\sum_{j=0}^{\infty}\frac{1}{(1+\alpha)^{2j}}$
-$\leq$ $~{}\sum_{j=0}^{\infty}\frac{1}{(1+\alpha)^{j}}$
-$=$ $~{}\frac{1}{1-\frac{1}{1+\alpha}}$
-$=$ $~{}1+\frac{1}{\alpha}$
-$=$ $~{}O(1/\alpha)$
-where the first step follows from we extend the finite sum to infinite sum, the second step follows from $\frac{1}{(1+\alpha)^{2j}}\leq\frac{1}{(1+\alpha)^{j}}$, the third step follows from the closed form of geometric sum, the fourth step follows from simple algebra, and the last step follows from $\alpha\in(0,1)$.
-Then from the proof of Lemma D.8, we can know that the variance is given by
-$O(\frac{R^{2}R_{w}^{2}\log^{3}n}{\alpha\epsilon^{2}})$ (5)
-since the sensitivity $\Delta=2R_{w}$ from Lemma C.3.
-Using Lemma B.3, we can have additive error bounded by
-$O(\frac{R\cdot R_{w}\log^{3/2}n}{\epsilon\sqrt{\alpha}}).$
-with probability $0.99$. ∎
-F.3 One Dimension Single Data Structure
-We therefore have the data structure that can solve weighted $\ell_{1}$-distance problem.
-Theorem F.6 (DPTreeDistance data structure, formal version of Theorem C.6).
-There is a data structure DPTreeDistance (Algorithm 5,6) that uses $O(n)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
-  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 6) It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
+[…]
     • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|$,
-    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$,
+    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}||\leq O(\epsilon^{-1}pR_{w}(2R)^{p}\log^{3/2}n)$,
     • it holds with probability $0.99$.
 Proof.
-The proofs follow from combining Lemma F.1 (running time of initialization), Lemma F.2 (running time of query), Lemma F.4 (DP of query), and Lemma F.5 (error of query) together. ∎
-Appendix G High-Dimensional Weighted $\ell_{1}$ Query
-In this section, we show how we can solve the high dimensional weighted $\ell_{1}$ distance problem, generalizing results from BLM+ [10].
-In Section G.1, we give the analysis of Algorithm 7.
-In Section G.2, we give the theorem of our DPTreeHighDim data structure.
-Algorithm 5,6 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{1}$ distance function.
+We set the total layers of one tree $L=(\log n)$. There are $p+1$ trees.
+Init Time and Space.
+The total number of nodes on one tree is $O(n)$. There are total $O(pn)$ values stored for $p+1$ trees. Adding the time of iterating all data points, initializing these values takes $O(pn)$ time.
+Query Time.
+Each query iterates through all layers.
+On each layer it takes $O(1)$ time to calculate $c_{\text{left},q}$ and $c_{\text{right},q}$.
+There are $(\log n)$ layers, and $p+1$ trees, so the total query time is $O(p\log n)$.
+Privacy Guarantees.
+For each $\mathcal{D}_{q}$ for $q\in\{0,1,\dots,p\}$, we input $a_{:,q}$.
+Since $X\in[0,R]^{n}$ and $w\in[-R_{w},R_{w}]^{n}$, the input range for $a_{:,q}$ is $R_{w}R^{q},R_{w}R^{q}]$.
+Then from Lemma B.3, sensitivity is $2R_{w}R^{q}$.
+From Lemma B.6, we know each $\mathcal{D}_{q}$ query is $(\epsilon/(p+1),\delta/(p+1))$-DP.
+By basic composition Fact A.8, the total differential privacy parameter is $(\epsilon,\delta)$.
+This completes the proof.
+Error Guarantees.
+The additive error consists of two parts.
+The first part is from the data in the leaf node which contains query $y$. The error is
+$\sum_{x_{k}\in[(j-1)\cdot R/2^{L},j\cdot R/2^{L})}|x_{k}-y|^{p}\leq n\cdot(\frac{R}{2^{L}})^{p}.$
+When $L=\log n$, this error is $O(R^{p}/n^{p-1})$.
+The second part is the Truncated Laplace noise.
+From the proof of Lemma B.7, we have each $\mathcal{D}_{q}$ for $q\in\{0,1,\dots,p\}$ has $O(L)$ independent $\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)$ noises for $L=\log n$ layers.
+Let $A$ be the noisy output of DistanceQuery in Algorithm 5 and $A_{*}=\sum_{k\in[n]}w_{k}|y-x_{k}|$ be the true output. Then, for our Algorithm 4 and 5, the variance is
+$\operatorname*{\mathrm{Var}}[\sum_{i}^{L}\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)]=$ $\sum_{i}^{L}\operatorname*{\mathrm{Var}}[\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)]$
+$=$ $~O(L^{3}\epsilon_{q}^{-2}\Delta_{q}^{2})$
+Replacing $\Delta_{q}=O(R^{q}R_{w})$ and $\epsilon_{q}=O(\epsilon/p)$, using Lemma A.3, with high probability 0.99, we have
+$|\sum_{i}^{L}\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)|\leq O(pR_{w}R^{q}L^{3/2}/\epsilon).$ (5)
+Then we bound the error with this inequality:
+$|A-A^{\prime}|\leq$ $~|\sum_{q=0}^{p}{p\choose q}y^{p-q}\sum_{i=1}^{L}((-1)^{p-q}\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)+(-1)^{q}\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L))|$
+$\leq$ $~\sum_{q=0}^{p}{p\choose q}y^{p-q}|\sum_{i=1}^{L}(\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L)+\mathrm{TLap}(\Delta_{q},\epsilon_{q}/L,\delta_{q}/L))|$
+$=$ $~\sum_{q=0}^{p}{p\choose q}y^{p-q}\cdot O(pR_{w}R^{q}L^{3/2}/\epsilon)$
+$=$ $~O(\epsilon^{-1}pR_{w}L^{3/2}\sum_{q=0}^{p}{p\choose q}y^{p-q}R^{q})$
+$=$ $~O(\epsilon^{-1}pR_{w}L^{3/2}(y+R)^{p})$
+$=$ $~O(\epsilon^{-1}pR_{w}(2R)^{p}\log^{3/2}n),$
+where the third step follows from Eq. (5), and the last step is from $L=\log n$ and $y\in[0,R]$.
+Therefore, by triangle inequality and two parts of error, the total error is
+$O(R^{p}/n^{p-1})+O(\epsilon^{-1}pR_{w}(2R)^{p}\log^{3/2}n)\leq O(\epsilon^{-1}pR_{w}(2R)^{p}\log^{3/2}n),$
+since $p\geq 1$ and $n\in\mathbb{N}_{+}$. This completes the proof. ∎
+Appendix E High-Dimensional Weighted $\ell_{p}^{p}$ Query
+In this section, we show how we can solve the high dimensional weighted $\ell_{p}^{p}$ distance problem, generalizing results from [8] and [45].
+In Section E.1, we give the analysis of Algorithm 6.
+In Section E.2, we give the theorem of our DPTreeHighDim data structure.
+Algorithm 4,5 can be naturally extended to higher dimensions because of the decomposability of the $\ell_{p}^{p}$ distance function.
 We construct $d$ separate one-dimensional distance query data structures, each corresponding to a coordinate projection of the dataset.
-G.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
-We now give the analysis of our Algorithm 7 for high dimensional weighted $\ell_{1}$-distance query.
-Algorithm 7 High-dimensional weighted $\ell_{1}$ distance query
-1:datastrucutre DPTreeHighDim $\triangleright$ Theorem G.3
+E.1 Privacy and Accuracy Analysis for High Dimensional Weighted Distance
+We now give the analysis of our Algorithm 6 for high dimensional weighted $\ell_{p}^{p}$ distance query.
+Algorithm 6 High-dimensional weighted $\ell_{p}^{p}$ distance query
+1:datastrucutre DPTreeHighDim $\triangleright$ Theorem E.3
 2:members
-3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$triangleright Alg. 5
+3:   ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeDistance}$ $\triangleright$ Alg. 4
 4:   $X:[0,R]^{n\times d}$
@@
 6:end members
-7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$) $\triangleright$ Lemma E.2
+7:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$)
 8:  $X\leftarrow X$
@@
 10:  for $i=1\to d$ do
-11:         $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 5
+11:   $\mathcal{D}_{i}.$Init($X_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$, $\delta/d$) $\triangleright$ Alg. 4
 12:  end for
 13:end procedure
-14:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma G.1, Lemma G.2
+14:procedure DistanceQuery($y\in[0,R]^{d}$) $\triangleright$ Lemma E.1, Lemma E.2
 15:  Value $\leftarrow 0$
 16:  for $i=1\to d$ do
-17:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(y_{i},\alpha)$triangleright Alg. 6
+[…]
 18:  end for
@@
 21:end datastrucutre
-Lemma G.1 (Privacy of DistanceQuery, Algorithm 7).
+Lemma E.1 (Privacy of DistanceQuery, Algorithm 6).
 If the following conditions hold
@@
   • Let $c\in(0,0.1)$ be a small constant and $A$ be the output of DistanceQuery in Algorithm 6, where each one-dimensional algorithm is configured to be $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP (see Line 11).
-  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ represent the true distance query value.
+  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{p}^{p}$ represent the true distance query value.
   • Let $\epsilon=O(\log(1/\delta^{\prime}))$.
@@
 Proof.
-The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem B.8.
+The $(\epsilon,\delta+\delta^{\prime})$-DP guarantee follows from the approximate DP advanced composition result Theorem A.10.
 Our algorithm instantiate each one-dimensional data structure with $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP total $d$ times.
-From advanced composition in Theorem B.8, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
+From advanced composition in Theorem A.10, for a sufficient small parameter $\epsilon$ and constant $c$, we have the final privacy loss parameter be:
 $O(c\epsilon\sqrt{2d\log(1/\delta^{\prime})}/\sqrt{d\log(1/\delta^{\prime})})=O(\epsilon)$
@@
 $d\delta/d+\delta^{\prime}=\delta+\delta^{\prime}.$
-Lemma G.2 (Accuracy of DistanceQuery, Algorithm 7).
+Lemma E.2 (Accuracy of DistanceQuery, Algorithm 6).
 If the following conditions hold
@@
   • Let $c\in(0,0.1)$ be a small constant and $A$ be the output of DistanceQuery in Algorithm 6, where each one-dimensional algorithm is configured to be $(c\epsilon/\sqrt{d\log(1/\delta^{\prime})},\delta/d)$-DP (see Line 11).
-  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ represent the true distance query value.
+  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{p}^{p}$ represent the true distance query value.
 With probability $0.99$, we have
-$|A-A_{*}|\leq\alpha A_{*}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}d\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
+$|A-A_{*}|\leq O(\epsilon^{-1}dp(2R)^{p}R_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
 Proof.
@@
 Observe that $A_{*}=\sum_{i=1}^{d}A_{*,i}$ and $A=\sum_{i=1}^{d}A_{i}$.
-We follow the similar idea in the proof of Lemma F.5.
-Let $z_{j,i}$ be the random variables that represent $z_{j}$ (used in the proof of Lemma F.5) for the $i$-th coordinate.
-We can observe that the overall error across $d$ coordinates can be upper bounded by
-$|\sum_{i=1}^{d}\sum_{j=0}^{O(\log(n)/\alpha)}\frac{Rz_{j,i}}{(1+\alpha)^{j}}|$
-where each $z_{j,i}$ is the sum of $O(\log n)$ truncated Laplace random variables independent to others.
-With $\epsilon$ scaled down by $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$ and $\delta$ scaled down by $\delta/d$, the variance of each individual dimension is given by (see Eq.
-(5))
-$O(\alpha^{-1}\epsilon^{-2}dR^{2}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
+We follow the similar idea in the proof of Theorem D.1.
+With $\epsilon$ scaled down by $c\epsilon/\sqrt{d\log(1/\delta^{\prime})}$ and $\delta$ scaled down by $\delta/d$, the variance of each individual dimension is given by (see proof of Theorem D.1)
+$O(\epsilon^{-2}dp^{2}(2R)^{2p}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
 Thus, the total variance for $d$ instantiated data structures is then
-$O(\alpha^{-1}\epsilon^{-2}d^{2}R^{2}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
-Finally, from Lemma B.3, we have the additive error given by
-$O(\alpha^{-1/2}\epsilon^{-1}dRR_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
-G.2 High Dimension Single Data Structure
-We have the data structure that can solve weighted $\ell_{1}$-distance problem in $d$-dimensional data.
-Theorem G.3 (DPTreeHighDim data structure).
-There is a data structure DPTreeHighDim (Algorithm 7) that uses $O(nd)$ spaces to solve weighted $\ell_{1}$-distance query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
-  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1))$.
-(Algorithm 7) It takes $O(nd)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 7) It takes $O(\alpha^{-1}d\log^{2}n)$ time to output a number $z$ such that
-    • the process of output $z$ satisfies is $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$,
-    • $|z-\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}|\leq\alpha\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}+O(\epsilon^{-1}\alpha^{-1/2}RR_{w}d\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
+$O(\epsilon^{-2}d^{2}p^{2}(2R)^{2p}R_{w}^{2}\log(1/\delta^{\prime})\log^{3}n).$
+Finally, from Lemma A.3, we have the additive error given by
+$O(\epsilon^{-1}dp(2R)^{p}R_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n).$
+E.2 High Dimension Single Data Structure
+We have the data structure that can solve weighted $\ell_{p}^{p}$ distance problem in $d$-dimensional data.
+Theorem E.3 (DPTreeHighDim data structure).
+There is a data structure DPTreeHighDim (Algorithm 6) that uses $O(npd)$ spaces to solve weighted $\ell_{p}^{p}$ distance query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
+[…]
+[…]
+    • the process of output $z$ satisfies is $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{p}^{p}$,
+    • $|z-\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}|\leq O(\epsilon^{-1}dp(2R)^{p}R_{w}\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
     • it holds with probability $0.99$.
@@
 For the runtime analysis, since we loop data structure DPTreeDistance $d$ times, an additional $d$ factor will appear for both initialization and query time complexity.
-The DP is proved by Lemma G.1.
-The accuracy is proved by Lemma G.2.
+The DP is proved by Lemma E.1.
+The accuracy is proved by Lemma E.2.
 ∎
-Appendix H Adaptive Query
-In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from QRS+ [86].
+Appendix F Adaptive Query
+In this section, we introduce how we can solve the adaptive query problem by our algorithm, using some tools from [70].
 Our idea is that, if we can prove that our algorithm can solve any query in the query space with certain error.
 Then, since adaptive query must lie in this space, we can handle adaptive query.
-In Section H.1, we show how we can boost the constant probability of our algorithm to high probability.
-In Section H.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net.
-In Section H.3, we show how we can bound all points in the query space by introducing an additive error.
-In Section H.4, we examine the effects of different norms on our adaptive query proof.
-First, from Theorem G.3, given query $y\in[0,R]^{d},\alpha\in(0,1)$ we have DistanceQuery$(y,\alpha)$ that can solve $d$-dimension weighted $\ell_{1}$-distance problem with constant probability $0.99$.
+In Section F.1, we show how we can boost the constant probability of our algorithm to high probability.
+In Section F.2, we show how we can apply the notion of $\epsilon_{0}$-net and bound all query points in net.
+In Section F.3, we show how we can bound all points in the query space by introducing an additive error.
+First, from Theorem E.3, given query $y\in[0,R]^{d}$ we have DistanceQuery(y)(y) that can solve $d$-dimension weighted $\ell_{p}^{p}$ distance problem with constant probability $0.99$.
 Now we show how to improve it to solve adaptive query problem.
-H.1 Boost the Constant Probability to High Probability
-We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma B.2.
-Lemma H.1 (Using Chernoff bound to boost the probability).
+Here, we focus on the case when $p=1$.
+F.1 Boost the Constant Probability to High Probability
+We can repeat the data structure multiple times and take the median to boost the constant probability using Chernoff bound from Lemma A.2.
+Lemma F.1 (Using Chernoff bound to boost the probability).
 If the following conditions hold:
   • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
-  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
+  • Let the failure probability $p_{f}\in(0,0.01)$.
   • We create $l=O(\log(1/p_{f}))$ independent copies of data structure DPTreeHighDim and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
-  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ be the true answer.
-  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
-Then for each fixed query point $y$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}$.
+  • Let $B=O(\epsilon^{-1}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
+Then for each fixed query point $y$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $B$ with probability $1-p_{f}$.
 Proof.
-By basic composition Fact B.6, we prove the DP. Similar to the proof of Theorem D.2, we prove the error by Chernoff bound (Lemma B.2). ∎
-H.2 From Each Fixed Query Point to All On-net Points
+By basic composition Fact A.8, we prove the DP. Similar to the proof of Theorem B.2, we prove the error by Chernoff bound (Lemma A.2). ∎
+F.2 From Each Fixed Query Point to All On-net Points
 In this section, we build $\epsilon_{0}$-net and generalize from each fixed query point to all on-net points.
-Definition H.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in Ver [96]).
-We define $N$ be $\ell_{p}$ $\epsilon_{0}$-net of :=${\mathcal{B}}:=\{q\in[0,R]^{d}\}$ such that, for every point $q$ in ${\mathcal{B}}$, there exists $y\in N$ satisfying $\|y-q\|_{p}\leq\epsilon_{0}$.
-Fact H.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
+Definition F.2 ($\ell_{p}$ $\epsilon_{0}$-net, see Definition 4.2.1 in [77]).
+We define $N$ be $\ell_{p}$ $\epsilon_{0}$-net of ${\mathcal{B}}:=\{q\in[0,R]^{d}\}$ such that, for every point $q$ in $\mathcal{B}$}, there exists $y\in N$ satisfying $\|y-q\|_{p}\leq\epsilon_{0}$.
+Fact F.3 ($\ell_{\infty}$ $\epsilon_{0}$-net).
 Let $N$ be the $\ell_{\infty}$ $\epsilon_{0}$-net of $\mathcal{B}$}, and $|N|$ be the size of net $N$. We have $|N|\leq(5R/\epsilon_{0})^{d}$.
-Fact H.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in Woo [102]).
+Fact F.4 ($\ell_{2}$ $\epsilon_{0}$-net, see Lemma 5 in [83]).
 Let $N$ be the $\ell_{2}$ $\epsilon_{0}$-net of $\mathcal{B}$}, and $|N|$ be the size of net $N$. We have $|N|\leq(5R/\epsilon_{0})^{d}$.
-Fact H.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in GS [41]).
+Fact F.5 ($\ell_{1}$ $\epsilon_{0}$-net, see Theorem 2 in [28]).
 Let $N$ be the $\ell_{1}$ $\epsilon_{0}$-net of $\mathcal{B}$}, and $|N|$ be the size of net $N$. We have $|N|\leq(5R\sqrt{d}/\epsilon_{0})^{d}$.
-Lemma H.6 (From for each query point to for all points in net).
+Lemma F.6 (From for each query point to for all points in net).
 If the following conditions hold:
@@
   • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
-  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
+  • Let the failure probability $p_{f}\in(0,0.01)$.
   • We create $l=O(\log(|N|/p_{f}))$ independent copies of data structure DPTreeHighDim and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
-  • Let $A_{*}=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$ be the true answer.
-  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
-Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $\alpha A_{*}+B$.
+  • Let $B=O(\epsilon^{-1}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
+Then with probability $1-p_{f}$, for all query points $y\in N$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error is upper bounded by $B$.
 Proof.
-By basic composition Fact B.6, we prove the DP. From Lemma H.1, we know for each $y\in N$, the error is upper bounded by $\alpha A_{*}+B$ with probability $1-p_{f}/|N|$.
-Then, by union bound, with probability $1-p_{f}$, the error of all $|N|$ query points in the net $y\in N$ is upper bounded by $\alpha A_{*}+B$. ∎
-H.3 From Net Points to All Points
+By basic composition Fact A.8, we prove the DP. From Lemma F.1, we know for each $y\in N$, the error is upper bounded by $B$ with probability $1-p_{f}/|N|$.
+Then, by union bound, with probability $1-p_{f}$, the error of all $|N|$ query points in the net $y\in N$ is upper bounded by $B$. ∎
+F.3 From Net Points to All Points
 In this section, we show how to generalize points from net to all points in the query space. Since adaptive query must lie in this space, we complete the proof of adaptive query.
-Lemma H.7 (Lipschitz of query function).
+Lemma F.7 (Lipschitz of query function).
 If the following conditions hold:
@@
 We can show
-$|Z(y)-Z(\widetilde{y})|=$ $~{}|\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}-\sum_{i\in[n]}w_{i}\|\widetilde{y}-x_{i}\|_{1}|$
+[…]
 $\leq$ $~\sum_{i\in[n]}|w_{i}|\cdot|\|y-x_{i}\|_{1}-\|\widetilde{y}-x_{i}\|_{1}|$
@@
 ∎
-Lemma H.8 (From points in net to all points in query space).
+Lemma F.8 (From points in net to all points in query space).
 If the following conditions hold:
@@
   • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
-  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
+  • Let the failure probability $p_{f}\in(0,0.01)$.
   • We create $l=O(\log((R/\epsilon_{0})^{d}/p_{f}))$ independent copies of data structure $\{\textsc{DPTreeHighDim}_{j}\}_{j=1}^{l}$ and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
-  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeHighDim}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
+  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeHighDim}_{j}.\textsc{DistanceQuery}(y)\}_{j=1}^{l})$.
   • Let $Z(y):=\sum_{i\in[n]}w_{i}\|y-x_{i}\|_{1}$, where $Z(y)$ is $L$-Lipschitz with $L=nR_{w}$.
-  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
+  • Let $B=O(\epsilon^{-1}lRR_{w}d\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
 Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfy
-$|f(y)-Z(q)|\leq\alpha Z(q)+B+2Ld\epsilon_{0}.$
+$|f(y)-Z(q)|\leq B+Ld\epsilon_{0}.$
 Proof.
-By basic composition Fact B.6, we prove the DP.
+By basic composition Fact A.8, we prove the DP.
 We define an event $E$ such that:
 $~\forall y\in N$
-$~{}|f(y)-Z(y)|\leq\alpha Z(y)+B.$
-From Lemma H.1, with $l=O(\log(|N|/p_{f}))$ we know
+$~|f(y)-Z(y)|\leq B.$
+From Lemma F.1, with $l=O(\log(|N|/p_{f}))$ we know
 $\Pr[\mathrm{event}~E~\mathrm{holds}]\geq 1-p_{f}$
@@
 $=$ $~O(\log((R/\epsilon_{0})^{d}/p_{f})$
-where the first step follows from definition of $l$, the second step follows from Fact H.3.
-We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition H.2), for each $q\notin N$, there exists $y\in N$ such that
+where the first step follows from definition of $l$, the second step follows from Fact F.3.
+We condition on event $E$ to be held. Then, by definition of $\ell_{\infty}$ $\epsilon_{0}$-net (see Definition F.2), for each $q\notin N$, there exists $y\in N$ such that
 $\|y-q\|_{\infty}\leq\epsilon_{0}$ (6)
@@
 $\leq$ $~L\cdot d\epsilon_{0}$ (7)
-where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (6).
+where the first step follows from Lemma F.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from Eq. (6).
 Using the on-net query $y$ to answer the off-net query $q$, for any $q\notin N$, we have
@@
 $\leq$ $~|f(y)-Z(y)|+L\cdot d\cdot\epsilon_{0}$
-$\leq$ $~{}\alpha Z(y)+B+L\cdot d\cdot\epsilon_{0}$
-$\leq$ $~{}\alpha Z(q)+B+2L\cdot d\cdot\epsilon_{0}$ (8)
-where the first step follows from triangular inequality, the second step follows from Eq. (H.3), the third step follows from Lemma H.6, and the last step follows from Eq. (H.3).
+$\leq$ $~B+L\cdot d\cdot\epsilon_{0}$ (8)
+where the first step follows from triangular inequality, the second step follows from Eq. (F.3), the third step follows from Lemma F.6.
 Thus, we complete the proof. ∎
 Therefore, even adaptive queries can be answered accurately, since any adaptive query can be assumed in $\mathcal{B}$}.
-H.4 Effect of Different Norms on the Result
-In the above proof, we have two different measure spaces, i.e. $\ell_{\infty}$ distance of $\epsilon_{0}$-net (Definition H.2) and $\ell_{1}$ Lipschitz (Lemma H.7).
-One might ask, will the norm we choose in two spaces have an impact on the final result? We can show that the norm we choose currently is sufficient to use.
-For different norms, the only differences in the proofs will be Lipschitz smoothness in Eq. (H.3) and the cardinality of $\epsilon_{0}$-net, i.e. $|N|$ in Fact H.3.
-Lemma H.9.
-If we use $\ell_{\infty}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(nR/p_{f}))$.
-Proof.
-If we use $\ell_{\infty}$ to bound the distance to net, Eq. (H.3) is:
-$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
-$\leq$ $~{}nR_{w}\cdot d\|y-q\|_{\infty}$
-$\leq$ $~{}nR_{w}\cdot d\epsilon_{0}$
-where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq d\|x\|_{\infty}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{\infty}$ $\epsilon_{0}$-net.
-Then, Eq. (H.3) is
-$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot d\cdot\epsilon_{0}$
-For $\ell_{\infty}$ distance, we have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact H.3.
-We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot d\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
-$l=$ $~{}O(\log(|N|/p_{f})$
-$=$ $~{}O(\log((R/\epsilon_{0})^{d}/p_{f})$
-$=$ $~{}O(\log((nR)^{d}/p_{f}))$
-$=$ $~{}O(d\log(nR/p_{f}))$
-where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
-Lemma H.10.
-If we use $\ell_{2}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(nR/p_{f}))$.
-Proof.
-If we use $\ell_{2}$ to bound the distance to net, Eq. (H.3) changes to be:
-$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
-$\leq$ $~{}nR_{w}\cdot\sqrt{d}\cdot\|y-q\|_{2}$
-$\leq$ $~{}nR_{w}\cdot\epsilon_{0}\sqrt{d}$
-where the first step follows from Lemma H.7, the second step follows from $\|x\|_{1}\leq\sqrt{d}\cdot\|x\|_{2}$ for $x\in\mathbb{R}^{d}$, and the last step follows from $\ell_{2}$ $\epsilon_{0}$-net.
-Then, Eq. (H.3) changes to be
-$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot\epsilon_{0}\sqrt{d}$
-For $\ell_{2}$ distance, we also have $|N|\leq(5R/\epsilon_{0})^{d}$ in Fact H.4.
-We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\sqrt{d}\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
-$l=$ $~{}O(\log(|N|/p_{f})$
-$=$ $~{}O(\log((R/\epsilon_{0})^{d}/p_{f})$
-$=$ $~{}O(\log((nR)^{d}/p_{f}))$
-$=$ $~{}O(d\log(nR/p_{f}))$
-where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
-Lemma H.11.
-If we use $\ell_{1}$ $\epsilon_{0}$-net and use $\ell_{1}$ Lipschitz in Lemma H.8, we have copies of data structure $l=O(d\log(ndR/p_{f}))$.
-Proof.
-If we use $\ell_{1}$ to bound the distance to net, Eq. (H.3) changes to be:
-$|Z(y)-Z(q)|\leq$ $~{}nR_{w}\cdot\|y-q\|_{1}$
-$\leq$ $~{}nR_{w}\cdot\epsilon_{0}$
-where the first step follows from Lemma H.7, and the last step follows from $\ell_{1}$ $\epsilon_{0}$-net.
-Then, Eq. (H.3) changes to be
-$|f(y)-Z(q)|\leq$ $~{}\alpha Z(q)+B+2nR_{w}\cdot\epsilon_{0}$
-For $\ell_{1}$ distance, we have $|N|\leq(5R\sqrt{d}/\epsilon_{0})^{d}$.
-We can choose $\epsilon_{0}=\Theta(1/n)$ to hide $nR_{w}\cdot\epsilon_{0}$ term in $B$ in Lemma H.8. Thus,
-$l=$ $~{}O(\log(|N|/p_{f})$
-$=$ $~{}O(\log((R\sqrt{d}/\epsilon_{0})^{d}/p_{f})$
-$=$ $~{}O(\log((nR\sqrt{d})^{d}/p_{f}))$
-$=$ $~{}O(d\log(nRd/p_{f}))$
-where the last step follows from $\log(a^{d}/b)=O(d\log(a/b))$ for any $a>1,0<b<1,d>1$. ∎
-From the above analysis, we can show that $\ell_{\infty}$ or $\ell_{2}$ $\epsilon_{0}$-net is slightly better than $\ell_{1}$ $\epsilon_{0}$-net.
-  • $\ell_{\infty}$ $\epsilon_{0}$-net, Lemma H.9: we have $l=O(d\log(nR/p_{f}))$.
-  • $\ell_{2}$ $\epsilon_{0}$-net, Lemma H.10: we have $l=O(d\log(nR/p_{f}))$.
-  • $\ell_{1}$ $\epsilon_{0}$-net, Lemma H.11: we have $l=O(d\log(nRd/p_{f}))$.
-Thus, the norm we choose for $\epsilon_{0}$-net is sufficient good.
-Appendix I Softmax Activation
-In this section, we introduce how we extend previous $\ell_{1}$ distance results to the Softmax activation function, which is the most widely used distance measure in attention mechanism based models.
-In Section I.1, we show how to extend to the Softmax distance function in Lemma I.6.
-In Section I.2, we show how to adjust our algorithms.
-In Section I.3, we extend our algorithm to be robust to adaptive query.
-In Section I.4, we give the proof of our main result Theorem 3.1.
-I.1 Exponential Inner Product
-In this section, we show how we obtain the Softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from AS [6].
-Definition I.1 (Definition 3.1 in AS [6]).
+Appendix G Softmax Activation
+In this section, we introduce how we extend previous $\ell_{p}^{p}$ distance results to the Softmax activation function, which is the most widely used distance measure in attention mechanism based models.
+In Section G.1, we show how to extend to the Softmax distance function in Lemma G.6.
+In Section G.2, we show how to adjust our algorithms.
+In Section G.3, we extend our algorithm to be robust to adaptive query.
+In Section G.4, we give the proof of our main result Theorem 4.1.
+G.1 Exponential Inner Product
+In this section, we show how we obtain the Softmax distance using $\ell_{2}^{2}$ distance query. First, we provide some helpful results from [4].
+Definition G.1 (Definition 3.1 in [4]).
 Let $r\geq 1$ denote a positive integer.
@@
   • $|\widetilde{A}_{i,j}-A_{i,j}|\leq\epsilon\cdot A_{i,j}$ for all $(i,j)\in[n]^{2}$.
-Lemma I.2 (Lemma 3.4 in AS [6]).
+Lemma G.2 (Lemma 3.4 in [4]).
 Suppose $Q,K\in\mathbb{R}^{n\times d}$, with $\|Q\|_{\infty}\leq R$, and $\|K\|_{\infty}\leq R$.
@@
 For accuracy parameter $\epsilon\in(0,0.1)$, there is a positive integer $s$ bounded above by
-$s=O\Big{(}\max\Big{\{}\frac{\log(1/\epsilon)}{\log(\log(1/\epsilon)/R)},R^{2}\Big{\}}\Big{)},$ (9)
+$s=O\Big(\max\Big\{\frac{\log(1/\epsilon)}{\log(\log(1/\epsilon)/R^{2})},R^{2}\Big\}\Big),$ (9)
 and a positive integer $r$ bounded above by
 $r\leq{2s+2d\choose 2s}$ (10)
-such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition I.1) of $A\in\mathbb{R}^{n\times n}$.
+such that: There is a matrix $\widetilde{A}\in\mathbb{R}^{n\times n}$ that is an $(\epsilon,r)$-approximation (Definition G.1) of $A\in\mathbb{R}^{n\times n}$.
 Furthermore, the matrices $U_{1}$ and $U_{2}$ defining $\widetilde{A}$ can be computed in $O(n\cdot r)$ time.
-Here we consider the vector version of Lemma I.2.
-Definition I.3.
+Here we consider the vector version of Lemma G.2.
+Definition G.3.
 We define $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$.
-Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in AS [6].
-Remark I.4.
+Then, we have $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ where $P(\cdot)$ is polynomial kernel function defined in [4].
+Remark G.4.
 We use $\Gamma_{R,s}$ to denote the value range of our polynomial kernel methods function, i.e., $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$.
@@
 We use the polynomial approximation method, which has been applied to accelerate Transformer model extensively [4, 5, 6, 56, 54].
-Lemma I.5 (Polynomial approximation).
-For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in AS [6] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
+Lemma G.5 (Polynomial approximation).
+For any accuracy parameter $\epsilon_{s}\in(0,0.1)$, let $R\geq 1$, and let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in [4] where $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$.
 Then, for any $x,y\in[0,R]^{d}$, we have
@@
 Proof.
-Let $n=1$. The proof follows from directly applying Lemma I.2. ∎
-Using the results from AS [6] above, we can extend our results to Softmax activation.
-Lemma I.6 (Weighted Softmax approximation, formal version of Lemma C.7).
+Let $n=1$. The proof follows from directly applying Lemma G.2. ∎
+Using the results from [4] above, we can extend our results to Softmax activation.
+Lemma G.6 (Weighted Softmax approximation ).
 Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
 Let $R\geq 1$.
-Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma I.5.
+Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$.
+Let $P(x):[0,R]^{d}\to[0,\Gamma_{R,s}]^{r}$ be the $s$-th order polynomial kernel function defined in Lemma G.5.
 Then we can approximate exponential inner product using polynomial kernel function:
-$w^{\top}\exp(Xy/d)=$ $~{}-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$
-$+~{}O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s})$
+$|-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}+\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})-w^{\top}\exp(Xy/d)|$
+$=$ $~O(|w^{\top}\exp(Xy/d)\cdot\epsilon_{s}|)$
 Moreover, the vectors $P(\cdot)$ can be computed in $O(r)$ time.
 Proof.
-From Lemma I.5, we can use polynomial kernel to approximate the Softmax function:
-$w^{\top}\exp(Xy/d)=$ $~{}\sum_{i\in[n]}w_{i}P(x_{i})^{\top}P(y)+O(w^{\top}\exp(Xy/d)\cdot\epsilon_{s}).$
-The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma I.5.
+From Lemma G.5, we can use a polynomial kernel to approximate the Softmax function:
+$|\sum_{i\in[n]}w_{i}P(x_{i})^{\top}P(y)-w^{\top}\exp(Xy/d)|=$ $O(|w^{\top}\exp(Xy/d)\cdot\epsilon_{s}|).$
+The proof of approximation error and time complexity of constructing $P(\cdot)$ follows from Lemma G.5.
 Then, we can show
@@
 where the first step follows from $\|x-y\|_{2}^{2}=\|x\|_{2}^{2}+\|y\|_{2}^{2}-2\langle x,y\rangle$, and the second step follows $\|x\|_{2}^{2}=\sum_{j=1}^{d}|x_{j}|^{2}$ for $x\in\mathbb{R}^{d}$. ∎
-I.2 Algorithm Modifications
-Based on Lemma I.6, we can now extend our DP algorithms to handle Softmax activation.
+G.2 Algorithm Modifications
+Based on Lemma G.6, we can now extend our DP algorithms to handle Softmax activation.
 First, we need to construct $P(y)$ and $P(x_{i})$ for $i\in[n]$, each costing $O(r)$ time.
-Then, for the second term in Lemma I.6, i.e.
-$\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|^{2}_{2}+\|P(y)\|^{2}_{2})$, we don’t need to add DP noises in it; instead, we calculate this term exactly, preprocess it, and store the results in the algorithm.
-For the first term, $-\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}$, we can adjust our high dimensional DP distance query algorithm to solve it.
-Due to the decomposability of $\ell_{p}^{p}$ norm, i.e.
-$\sum_{i\in[n]}w_{i}\|x_{i}-y\|_{p}^{p}=\sum_{j\in[d]}\sum_{i\in[n]}w_{i}|x_{i,j}-y_{j}|^{p},$
-we can compute $\ell_{2}^{2}$ norm easily (see details in Lemma E.2).
-We then show how to extend our one dimensional $\ell_{1}$ distance algorithm (Algorithm 5 and 6) to $\ell_{2}^{2}$ distance with minor modifications.
-Theorem I.7 (DPTreeDistance $\ell_{2}^{2}$ distance).
-With $\alpha$ scaled down by a factor of $2$ and all Query instead multiplied by $R^{2}/(1+\alpha/2)^{2j}$ in Lines 8 and 13 of Algorithm 6, i.e., from
-  • Lines 8 and 13: Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R}{(1+\alpha)^{j}}$
-to
-  • Lines 8 and 13: Value $\leftarrow$ Value + ${\mathcal{D}}$.Query($l_{j},r_{j}$) $\cdot\frac{R^{2}}{(1+\alpha/2)^{2j}}$.
-The data structure DPTreeDistance (Algorithm 5,6) uses $O(n)$ spaces to solve weighted $\ell_{2}^{2}$-distance query problem for dataset $X\subset[0,R]$ and support the following operations:
-  • Init$(X\subset[0,R],n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1))$. (Algorithm 5) It takes $O(n)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R],\alpha\in(0,1))$. (Algorithm 6)
-    It takes $O(\alpha^{-1}\log^{2}n)$ time to output a number $z$ such that
-    • the process of output $z$ satisfies $(\epsilon,\delta)$-DP private, which computes $\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}$,
-    • $|z-\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}|\leq\alpha\sum_{i\in[n]}w_{i}|y-x_{i}|^{2}+O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$,
+For the first term, $\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-P(y)_{j}|^{2}$, we can adjust our high dimensional DP distance query algorithm to solve it.
+For the second term in Lemma G.6, i.e., $\frac{1}{2}\sum_{i\in[n]}w_{i}(\|P(x_{i})\|_{2}^{2}+\|P(y)\|_{2}^{2})$, it can be expressed as $\frac{1}{2}\sum_{j\in[r]}\sum_{i\in[n]}w_{i}|P(x_{i})_{j}-0|^{2}$ and $\frac{1}{2}\sum_{i\in[n]}w_{i}(\sum_{j\in[r]}P(y)_{j}^{2})$.
+The former can be computed using query 0, while the latter can be solved using the precomputed value $\sum_{i\in[n]}w_{i}$, which can be obtained from the data ${\bf 1}_{n}$ and query 0.
+Thus, we only need to consider the case $p=2$ in weighted $\ell_{p}^{p}$ distance algorithms.
+Now we can give our result that can answer Softmax query.
+Theorem G.7 (Softmax query, formal version of Theorem 5.2).
+Let $R\geq 1$.
+Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$.
+Let $\Gamma_{R,s}$ be defined in Definition G.3.
+Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
+There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
+[…]
+[…]
+    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
+    • $|z-w^{\top}\exp(Xy/d)|\leq|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|+O(\epsilon^{-1}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
     • it holds with probability $0.99$.
 Proof.
-The proof is similar to that of Theorem F.6, except that now our additive error includes $R$ increased by a power of $2$, i.e., from $O(\epsilon^{-1}\alpha^{-1/2}RR_{w}\log^{3/2}n)$ to $O(\epsilon^{-1}\alpha^{-1/2}R^{2}R_{w}\log^{3/2}n)$.
-∎
-Now we can give our result that can answer Softmax query.
-Theorem I.8 (Softmax query, formal version of Theorem 4.2).
-Let $R\geq 1$.
-Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}$ be defined in Definition I.3.
-Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
-There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
-  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1))$.
-(Algorithm 3) It takes $O(nr)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 3) It takes $O(\alpha^{-1}r\log^{2}n)$ time to output a number $z$ such that
-    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
-    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
-      ${}~{}~{}~{}~{}~{}~{}~{}+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
-    • it holds with probability $0.99$.
-Proof.
-The DP proof is the same as the proof of Lemma G.1.
+Let $P_{wx}:=\sum_{i\in[n]}w_{i}\|P(x_{i})\|^{2}_{2}$ and $s_{w}:=\sum_{i\in[n]}w_{i}$.
+Observe that $P_{wx}=\sum_{i\in[n]}w_{i}\|P(x_{i})-0\|^{2}_{2}$, meaning we can calculating $P_{wx}$ using query 0.
+Similarly, $s_{w}=\sum_{i\in[n]}w_{i}\|{\bf 1}_{n}-0\|^{2}_{2}$, meaning we can calculating $s_{w}$ using data ${\bf 1}_{n}$ and query 0.
+Thus, we compute $P_{wx},s_{w}$ in Line 16 and 19 in Algorithm 3 in this way.
+From the privacy proof of Lemma E.1 and the way we choose privacy parameters, similarly we get the output process of calculating $P_{wx}$ and Value is $(\epsilon/3,\delta/3+\delta^{\prime}/2)$-DP.
+Also, the output process of calculating $s_{w}$ is $(\epsilon/3,\delta/3)$-DP.
+Then, by Fact A.8, overall process is $(\epsilon,\delta+\delta^{\prime})$-DP in Line 28 of Algorithm 3.
 We then show the time complexity.
-From Lemma I.6, we know that constructing $P(\cdot)$ requires $O(r)$ time.
+From Lemma G.6, we know that constructing $P(\cdot)$ requires $O(r)$ time.
 In the first for loop of Init, the dominating time consumption is $O(nr)$.
@@
 In the DistanceQuery function, constructing $P(y)$ takes $O(r)$ time.
-Within the for loop, it requires $O(\alpha^{-1}r\log^{2}n)$.
-Thus, the total time complexity for DistanceQuery is $O(\alpha^{-1}r\log^{2}n)$.
+Within the for loop, it requires $O(r\log n)$.
+Thus, the total time complexity for DistanceQuery is $O(r\log n)$.
 The space complexity is $O(nr)$, since storing the $n\times r$ matrix $P$ is the dominating factor.
-The proof of the error follows from the triangle inequality by combining the errors in Lemma I.6 and Theorem I.7. ∎
-I.3 Adaptive Softmax
+The proof of the error follows from the triangle inequality by combining the errors in Lemma G.6 and Theorem E.3.
+Here, we omit the constant factors of 2 and 3 used for the privacy guarantee in Algorithm 3, incorporating it into the big-$O$ notation for the error analysis.
+To be more specific, in Line 28 of Algorithm 3, we have 3 terms to bound the error, namely $P_{wx},s_{w}\|P(y)\|_{2}^{2}$ and Value.
+From Lemma G.6, the first source of error comes from the approximation error introduced by polynomial kernel method, i.e.,
+$~|w^{\top}\exp(Xy/d)-\frac{1}{2}(\underbrace{\sum_{i\in[n]}w_{i}\|P(x_{i})\|^{2}_{2}}_{P_{wx}}+\underbrace{\sum_{i\in[n]}w_{i}}_{s_{w}}\|P(y)\|^{2}_{2}-\underbrace{\sum_{i\in[n]}w_{i}\|P(x_{i})-P(y)\|_{2}^{2}}_{\mathrm{Value}})|$
+$=$ $~O(|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|).$
+Then, the second source of error comes from the DP noises in Theorem E.3, where we use Algorithm 4 to compute the three terms.
+The two terms $P_{wx}$ and Value have additive error $O(\epsilon^{-1}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$ (Theorem E.3) due to to the way we choose the DP parameters, the appl==ication of advanced composition (Theorem A.10), and the transformation of the value range from $[0,R]$ to $0,\Gamma_{R,s}]$ by the polynomial kernel.
+See more details in the proof of Lemma E.2.
+As for the term $s_{w}\|P(y)\|_{2}^{2}$, the addtive error of $s_{w}$ is $O(\epsilon^{-1}R_{w}\log^{3/2}n)$.
+But since $\|P(y)\|_{2}^{2}\leq r\Gamma_{R,s}^{2}$, we have the addtive error is $O(\epsilon^{-1}\Gamma_{R,s}^{2}R_{w}r\log^{3/2}n)$ which is smaller than other two terms.
+We ignore the constant 3 introduced by summing three terms by triangle inequality of absolute function, i.e., $|-t_{1}+t_{2}+t_{3}|\leq|t_{1}|+|t_{2}|+|t_{3}|$.
+Finally, summing the two sources of error by triangle inequality, we finish the proof. ∎
+G.3 Adaptive Softmax
 In this section, we show how to make Algorithm 3 robust to adaptive query.
-We follow the same idea from Section H.
-We notice that, in the Softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$-distance in Section H.
+We follow the same idea from Section F.
+We notice that, in the Softmax activation, we have query function $Z(y):=w^{\top}\exp(Xy/d)$ different from the $\ell_{1}$ distance in Section F.
 Therefore, we need to recalculate Lipschitz constant first.
-Lemma I.9 (Lipschitz of weighted Softmax).
+Lemma G.8 (Lipschitz of weighted Softmax).
 If the following conditions hold:
@@
 We can show
-$|Z(y)-Z(\widetilde{y})|=$ $~{}|\sum_{i\in[n]}w_{i}\exp(x_{i}^{\top}y/d)-\sum_{i\in[n]}w_{i}\exp(x_{i}^{\top}\widetilde{y}/d)|$
+[…]
 $\leq$ $~\sum_{i\in[n]}|w_{i}|\cdot|\exp(x_{i}^{\top}y/d)-\exp(x_{i}^{\top}\widetilde{y}/d)|$
@@
 $\leq$ $~nd^{-1/2}RR_{w}\exp(R^{2})\|y-\widetilde{y}\|_{1}$
-where the first step follows from definition of $Z(y),Z(\widetilde{y})$, the second step follows from triangular inequality, the third step follows from Fact B.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$.
+where the first step follows from definition of Z(y),Z(y~)Z(y),Z($\widetilde${y}), the second step follows from triangular inequality, the third step follows from Fact A.4, the fourth step follows from Cauchy–Schwarz inequality $|u^{\top}v|\leq\|u\|_{2}\cdot\|v\|_{2}$ for $u,v\in\mathbb{R}^{d}$, the fifth step follows from $w_{i}\in[-R_{w},R_{w}]$ and $x_{i}\in[0,R]^{d}$, and the last step follows from $\|u\|_{2}\leq\|u\|_{1}$ for $u\in\mathbb{R}^{d}$.
 ∎
 Then we can show how to extend our algorithm to be robust to adaptive query.
-Lemma I.10 (Adaptive Softmax, formal version of Lemma C.8).
+Lemma G.9 (Adaptive Softmax ).
 If the following conditions hold:
@@
   • Let data set $X\in[0,R]^{n\times d}$, weights $w\in[-R_{w},R_{w}]^{n}$, query $y\in[0,R]^{d}$.
-  • Let relative error parameter $\alpha\in(0,1)$, the failure probability $p_{f}\in(0,0.01)$.
+  • Let the failure probability $p_{f}\in(0,0.01)$.
   • We create $l=O(\log((R/\epsilon_{0})^{r}/p_{f}))$ independent copies of data structure $\{\textsc{DPTreeSoftmax}_{j}\}_{j=1}^{l}$ (Algorithm 3) and take the median of the outputs with each data structure instantiated with $(\epsilon/l,(\delta+\delta^{\prime})/l)$-DP.
-  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y,\alpha)\}_{j=1}^{l})$.
+  • Let $f(y):=\mathrm{Median}(\{\textsc{DPTreeSoftmax}_{j}.\textsc{DistanceQuery}(y)\}_{j=1}^{l})$.
   • Let $Z(y):=w^{\top}\exp(Xy/d)$, where $Z(y)$ is $L$-Lipschitz with $L=nd^{-1/2}RR_{w}\exp(R^{2})$.
-  • Let $B=O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
+  • Let $B=O(\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$.
 Then with probability $1-p_{f}$, for all query points $q\in{\mathcal{B}}$, there exists a point $y\in N$ which is the closest to $q$, we can have the process of outputting the median of $l$ responses is $(\epsilon,\delta+\delta^{\prime})$-DP and the error satisfies
-$|f(y)-Z(q)|\leq(\alpha+\epsilon_{s})Z(q)+B+2n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}.$
+$|f(y)-Z(q)|\leq|\epsilon_{s}Z(q)|+B+O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0}).$
 Proof.
-The proof follows from the same idea as the proof of Lemma H.8, except that we use Theorem I.8 and the Lipschitz in Lemma I.9. ∎
-Theorem I.11 (Adaptive query Softmax data structure, formal version of Theorem 4.4).
+The proof follows from the same idea as the proof of Lemma F.8, except that we use Theorem G.7 and the Lipschitz in Lemma G.8. ∎
+Algorithm 7 Adaptive query data structure
+1:datastructure DPTreeSoftmaxAdaptive $\triangleright$ Theorem 5.3
+2:members
+3:   ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$ $\triangleright$ Algorithm 3
+4:end members
+[…]
+6:  $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
+7:  for $i=1\to l$ do
+[…]
+9:  end for
+10:end procedure
+11:procedure DistanceQuery($y\in[0,R]^{d}$)
+12:  $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
+13:  $r\leftarrow 0^{l}$
+14:  for $i=1\to l$ do
+15:   $r_{i}\leftarrow{\mathcal{D}}_{i}.\textsc{DistanceQuery}(y)$
+16:  end for
+17:  return Median of $r$
+18:end procedure
+19:end datastructure
+Theorem G.10 (Adaptive query Softmax data structure, formal version of Theorem 5.3).
 Let $R\geq 1$.
-Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$.
-Let $\Gamma_{R,s}$ be defined in Definition I.3.
+Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R^{2})},R^{2}\})$.
+Let $\Gamma_{R,s}$ be defined in Definition G.3.
 Let accuracy parameter be $\epsilon_{s}\in(0,0.1)$.
-Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter, and $p_{f}$ be the failure probability parameter.
+Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, and $p_{f}$ be the failure probability parameter.
 Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$.
 There is a data structure DPTreeSoftmaxAdaptive (Algorithm 7) that uses $O(lnr)$ spaces to solve weighted Softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
-  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$.
-(Algorithm 1) It takes $O(lnd)$ time to initialize the data structure.
-  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 1) It takes $O(\alpha^{-1}ld\log^{2}n)$ time to output a number $z$ such that
+[…]
+[…]
     • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
-    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
-      ${}~{}~{}~{}~{}~{}~{}~{}+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • $|z-w^{\top}\exp(Xy/d)|\leq|\epsilon_{s}\cdot w^{\top}\exp(Xy/d)|+O(\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
     • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
@@
 Proof.
-We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma I.10.
+We only need to show how to pick $\epsilon_{0}$ in the parameter $l$, because everything else is the same as Lemma G.9.
 We know the additive error introduced by adaptive query is $E_{a}:=O(n\sqrt{d}RR_{w}\exp(R^{2})\epsilon_{0})$ and the relative error introduced by polynomial kernel approximation is $E_{p}:=w^{\top}\exp(Xy/d)\cdot\epsilon_{s}$.
@@
 ∎
-I.4 Proof of Main Result
+G.4 Proof of Main Result
 In this section, we give the proof of our main result of Theorem 4.1.
-Theorem I.12 (Softmax cross-attention, formal version of Theorem 3.1).
+Theorem G.11 (Softmax cross-attention, formal version of Theorem 4.1).
 Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1.
-Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter.
-Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma C.7).
-Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3).
+Assume the input context length $n$ is large enough.
+Let $p_{f}$ be the probability of failure parameter.
+Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma G.6).
+Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3).
 Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$.
-There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
-  • We initialize $d$ data structures using Init$(K,n,V_{*,k},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$ (Algorithm 1), for $k\in[d]$.
-It takes $O(lnr)$ time to initialize one data structure.
-  • At query time, for user input $Q$, we process one token at a time by passing the $i$-th row of $Q$, denoted $Q_{i}\in\mathbb{R}^{d}$, to DistanceQuery$(Q_{i},\alpha\in(0,1))$ (Algorithm 1) for each $i\in[m]$.
-It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
+There is a data structure DPTreeCrossAttention (Algorithm 1) that uses $O(lnrd)$ spaces to ensure cross-attention DP and supports the following operations:
+[…]
+[…]
     • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP,
-    • the process of output $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
-    • the process of output $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
+    • the process of output $z$ has relative error $2\epsilon_{s}/(1-\epsilon_{s})$,
+    • the process of output $z$ has additive error $O((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n),$
     • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
@@
 Proof.
+We first prove the privacy and then prove error for each coordinate of the output $O$ of Algorithm 1.
+Proof of Privacy:
+From Theorem G.10, $\mathcal{D}_{k}$.DistanceQuery for $k\in\{0,1,\dots,d\}$ in Algorithm 1 answer $(\epsilon/2,\delta/2+\delta^{\prime}/2)$-DP queries that are robust to adaptive queries.
+By Fact A.8, the procedure for calculating each coordinate of vector $O$ is $(\epsilon,\delta+\delta^{\prime})$-DP in Line 15 of Algorithm 1.
+Proof of Error:
+We prove the error bound of the cross-attention module.
+We omit the constant factor of 2 used for the privacy guarantee in Algorithm 1, incorporating it into the big-$O$ notation for the error analysis.
+Let $AV$ be the true value and $\widetilde{AV}$ be the noisy value.
+Let $D$ be the true value and $\widetilde{D}$ be the noisy value.
+First, we use triangular inequality to decompose the error:
+$~|(D^{-1}AV)_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$
+$\leq$ $~|(D^{-1}AV)_{i,k}-(D^{-1}\widetilde{AV})_{i,k}|+|(D^{-1}\widetilde{AV})_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$ (11)
+We now prove for each term.
+Part 1: Error bound for AVAV
 From Section 4, we know that we can ensure matrix $AV$ in cross-attention computation satisfies DP.
 Next, from Theorem 5.3, for $i\in[m],j\in[n],k\in[d]$, we have $(AV)_{i,k}$ is $(\epsilon,\delta+\delta^{\prime})$-DP and also robust to adaptive query.
-For $D$, we compute the exact true value.
-By the post-processing property of DP (Fact B.5), our cross-attention process is DP.
-Let $(AV)_{i,k}$ be the true value and $\widetilde{(AV)}_{i,k}$ be the noisy value. Then, from Theorem 4.4, we have
-$|(AV)_{i,k}-\widetilde{(AV)}_{i,k}|\leq(\alpha+\epsilon_{s})\cdot(AV)_{i,k}+O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n).$ (11)
+Let $\zeta:=\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n$ denote the additive error. Then, from Theorem G.10, we have
+$|(AV)_{i,k}-\widetilde{(AV)}_{i,k}|\leq$ $~|\epsilon_{s}\cdot(AV)_{i,k}|+O(\zeta)$ (12)
 For $D_{i,i}$, we can show
@@
 because $\langle Q_{i},K_{j}\rangle\geq 0$ for bounded $Q,K$.
-Finally, we can show the error of one entry is bounded by
+Finally, we can show the error of first term in Eq. (G.4) is bounded by
 $|(D^{-1}AV)_{i,k}-(D^{-1}\widetilde{AV})_{i,k}|=$ $~|D_{i,i}^{-1}((AV)_{i,k}-\widetilde{(AV)}_{i,k})|$
 $=$ $~|D_{i,i}^{-1}|\cdot|((AV)_{i,k}-\widetilde{(AV)}_{i,k})|$
-$\leq$ $~{}n^{-1}(\alpha+\epsilon_{s})\cdot(AV)_{i,k}$
-$~{}+O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}(n))$
-where the first step follows from definition, the second step follows from simple algebra, and the last step follows from Eq.(11) and (12). ∎
+$\leq$ $~|\epsilon_{s}\cdot D_{i,i}^{-1}(AV)_{i,k}|+O(n^{-1}\zeta)$
+where the first step follows from definition, the second step follows from simple algebra, and the last step follows from Eq. (12) and (13).
+Part 2: Error bound for $D$
+[…]
+Recall that
+$D_{i,i}=\sum_{i=1}^{n}\exp(\langle Q_{i},K_{j}\rangle/d)).$
+This can be viewed as the weighted Softmax problem but with weight ${\bf 1}_{n}$.
+To be more clear, let us recall that $R_{w}$ is the upper bound of the entries in $V$, and define $R_{w}^{\prime}$ as the upper bound of the entries in ${\bf 1}_{n}$.
+Observe that we can reuse previous results in Theorem G.10 with adjustment only on the value of $R_{w}^{\prime}$ (which is 1) in $\mathcal{D}_{0}$.
+We wish to bound
+$|D^{-1}_{i,i}-\widetilde{D}^{-1}_{i,i}|=$ $~\frac{|D_{i,i}-\widetilde{D}_{i,i}|}{D_{i,i}\cdot\widetilde{D}_{i,i}}.$
+For the term $|D_{i,i}-\widetilde{D}_{i,i}|$, similar to Eq. (12), from Theorem G.10, we have
+$|D_{i,i}-\widetilde{D}_{i,i}|\leq|\epsilon_{s}\cdot D_{i,i}|+O(\zeta),$ (14)
+where we assume $R_{w}\geq 1=R_{w}^{\prime}$ and loose the $R_{w}^{\prime}$ in additive error parameter in $\mathcal{D}_{0}$ from 1 to $R_{w}$.
+Now we need the lower bound of $\widetilde{D}_{i,i}$. From Eq. (14), we have
+$\widetilde{D}_{i,i}\geq$ $~D_{i,i}-(|\epsilon_{s}\cdot D_{i,i}|+O(\zeta))\geq|(1-\epsilon_{s})\cdot D_{i,i}|-O(\zeta).$
+Then, we have
+$|D^{-1}_{i,i}-\widetilde{D}^{-1}_{i,i}|=$ $~D_{i,i}^{-1}\frac{|D_{i,i}-\widetilde{D}_{i,i}|}{\widetilde{D}_{i,i}}\leq D_{i,i}^{-1}\frac{|\epsilon_{s}\cdot D_{i,i}|+O(\zeta)}{|(1-\epsilon_{s})\cdot D_{i,i}|-O(\zeta)}$
+We assume $n$ is large enough and thus ignore other small factors.
+Observe that $O(\zeta)=O(\log^{3/2}n)$, and $D_{i,i}\geq n=O(n)$ from Part 1.
+Thus, $O(\zeta)$ is a small order term compared to $D_{i,i}$.
+As a consequence, we get
+$|D^{-1}_{i,i}-\widetilde{D}^{-1}_{i,i}|\leq D_{i,i}^{-1}\frac{|\epsilon_{s}\cdot D_{i,i}|}{|(1-\epsilon_{s})\cdot D_{i,i}|}=D_{i,i}^{-1}\frac{\epsilon_{s}}{(1-\epsilon_{s})},$ (15)
+since $\epsilon_{s}\in(0,0.1)$.
+From Eq. (12), we have
+$|\widetilde{(AV)}_{i,k}|\leq$ $~(1+\epsilon_{s})\cdot|(AV)_{i,k}|+O(\zeta)$
+We consider the second term in Eq.(G.4). Then,
+$~|(D^{-1}\widetilde{AV})_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$
+$=$ $~|D^{-1}_{i,i}-\widetilde{D}^{-1}_{i,i}|\cdot|(\widetilde{AV})_{i,k}|$
+$\leq$ $~D_{i,i}^{-1}\frac{\epsilon_{s}}{(1-\epsilon_{s})}((1+\epsilon_{s})\cdot|(AV)_{i,k}|+O(\zeta))$
+$=$ $~\epsilon_{s}\frac{(1+\epsilon_{s})}{(1-\epsilon_{s})}\cdot D_{i,i}^{-1}|(AV)_{i,k}|+O(\frac{\epsilon_{s}}{(1-\epsilon_{s})}D_{i,i}^{-1}\zeta)$
+$\leq$ $~\epsilon_{s}\frac{(1+\epsilon_{s})}{(1-\epsilon_{s})}\cdot|D_{i,i}^{-1}(AV)_{i,k}|+O(\frac{\epsilon_{s}}{(1-\epsilon_{s})}n^{-1}\zeta)$
+where the first step follows from simple algebra, the second step follows from the previous derived upper bounds, the third step follows from simple algebra, and the last step follows from Eq.(13).
+Part 3: Final error bound
+Combining results from Part 1 and 2, the final error bound is
+$~|(D^{-1}AV)_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$
+$\leq$ $~|(D^{-1}AV)_{i,k}-(D^{-1}\widetilde{AV})_{i,k}|+|(D^{-1}\widetilde{AV})_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$
+$=$ $~\epsilon_{s}\cdot|D_{i,i}^{-1}(AV)_{i,k}|+O(n^{-1}\zeta)+~\epsilon_{s}\frac{(1+\epsilon_{s})}{(1-\epsilon_{s})}\cdot|D_{i,i}^{-1}(AV)_{i,k}|+O(\frac{\epsilon_{s}}{(1-\epsilon_{s})}n^{-1}\zeta)$
+$=$ $~\frac{2\epsilon_{s}}{(1-\epsilon_{s})}\cdot|(D^{-1}AV)_{i,k}|+O((1-\epsilon_{s})^{-1}n^{-1}\zeta)$
+Therefore, we prove the error bound.